### v2

# The gauge symmetry - unbroken and broken - in the QCD ground state

Abstract
The main purpose of this work is to report on convincing progress made in the investigation of the one of the long-standing conceptual problems in QCD, namely whether the gauge symmetries of the QCD Lagrangian and its ground state coincide or not? For this purpose, we use only the tensor algebra derivation rigorous rules in dealing with the corresponding skeleton loop integrals, contributing into the full gluon self-energy. It has been exactly (i.e., without making any kind of the simplifications) proven that in the QCD ground state may exist the two principally different types of the gluon field configurations. One of them maintains the QCD Lagrangian’s gauge symmetry, but another one does not.The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared. The exact proof has been given that it cannot be disregarded from the theory and its ground state by any means. The gauge symmetry of the QCD Lagrangian forbids such term, but the perturbative renormalizability of the theory will not be affected. This allows to present the new significant insights into the gauge and dynamical structures of the true QCD ground state, and sheds light on a mass squared dynamical generation in it.
keywords:
theoretical particle physics, Quantum Chromodynamics, ground state, gluon full self-energy, dynamical and gauge structures.
PACS:
11.10.-z, 11.15.-q, 12.38.-t, 12.38.Aw, 12.38.Lg
MSC:
[2020] 30D30, 81Q40, 81T13, 81T16
1 Introduction
The quark model (QM) treats the strongly-interacting hadrons (baryons and mesons) as bound-states of quarks, by emitting and absorbing gluons. The theory, which purpose is to describe the properties of the observed hadrons in terms of the non-observable quarks and gluons from first principles, is Quantum Chromodynamics (QCD) 1 ; 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 . It is widely accepted as the quantum gauge field theory of strong interactions. Being such a theory, it has to undergo the two phase transitions: the first one at the fundamental (microscopic) quark-gluon level is – the confinement phase transition – in order to explain why all the physical states are colour-singlets. The second one at the (macroscopic) hadronic level is – the PCAC (partially conserved axial currents) phase transition – in order to explain the soft pion physics 2 ; 23 .
However, this purpose remains a formidable task yet because of the multiple dynamical and topological complexities of low-energy particles (hadrons) physics, originated from QCD and its ground state (vacuum). This happens because QCD as a gauge theory of the gluon-gluon and quark-gluon strong interactions suffers from a few important conceptual problems. Here we focus on the two of them, being in the close connection with each other.
A. Whether the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is a symmetry of its ground state or not?
B. The dynamical generation of a mass squared in the vacuum of QCD, since the gauge invariance of its Lagrangian forbids such kind of terms, like the gluon mass term $M^{2}_{g}A_{\mu}A_{\mu}$.
The solutions of these fundamental problems is a key to the explanation of the first confinement phase transition in QCD. The properties and symmetries of the QCD Lagrangian, and thus including its Yang – Mills (YM) part, are well-known 1 ; 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 ; 10 . The propagation of gluons is one of the main dynamical effects in the QCD ground state. The importance of the corresponding equation of motion is due to the fact that its investigation is supposed to reflect the dynamical and gauge structures of the QCD ground state. The gluon Schwinger – Dyson (SD) equation 2 ; 10 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 is highly non-linear (NL) because of the self-interaction of massless gluon modes. They are taken into account by the full gluon self-energy, which is the main part of the gluon SD equation itself. We will show that the general properties of the full gluon self-energy point out on some new dynamical and gauge aspects of the true QCD ground state. Due to these inputs, the novel insights into a mass dynamical generation in the QCD vacuum are also present within proposed below formalism, based on the tensor algebra derivation rigorous rules only.
2 The gluon SD equation
The structure of the gluon SD equation is given in this section in some necessary details. Analytically it looks like
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{0}_{\mu\rho}(q)i\Pi_{\rho\sigma}(q;D)D_{\sigma\nu}(q),$ (2.1)
where $D_{\mu\nu}(q)$ and $D^{0}_{\mu\nu}(q)$ denote the full gluon propagator and its free counterpart, respectively. $\Pi_{\rho\sigma}(q;D)$ is the full gluon self-energy which depends on the full gluon propagator due to the non-abeian character of QCD. Here and everywhere below we omit the colour group indices, for simplicity, because of their final factorization, for example $D^{ab}_{\mu\nu}(q)=D_{\mu\nu}(q)\delta^{ab}$. Eq. (2.1) in terms of the corresponding skeleton loop diagrams is shown in Fig. 1.
Figure 1: The SD equation for the full gluon propagator as present in 10 .
Here stringy lines are for the free gluon propagator, while $D$ denotes its full counterpart. $S$ with solid lines denotes the full quark propagator, and $\Gamma$ denotes the full quark-gluon vertex. $G$ with dashed lines denotes the full ghost propagator, and $G_{\mu}$ is the full ghost-gluon vertex. Finally, $T_{3}$ and $T_{4}$ denote the full 3- and 4-gluon vertices, respectively.
The full gluon self-energy is convenient to present as the sum of the three independent terms, namely
$\Pi_{\rho\sigma}(q;D)=\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)+\Pi_{\rho\sigma}^{t}(D),$ (2.2)
where $\Pi^{q}_{\rho\sigma}(q)$ describes the skeleton loop contribution of the quark degrees of freedom as an analogue to the vacuum polarization tensor in Quantum Electrodynamics (QED) 7 ; 21 . Note that here and below the superscript or subscript ’$q$’ means quark (not to be mixed up with the gluon momentum variable $q$). The gluon part of the full gluon self-energy by itself is the sum of a few independent terms as follows:
$\Pi^{g}_{\rho\sigma}(q;D)=\Pi^{gh}_{\rho\sigma}(q)+\Pi^{(1)}_{\rho\sigma}(q;D^{2})+\Pi^{(2)}_{\rho\sigma}(q;D^{4})+\Pi^{(2^{\prime})}_{\rho\sigma}(q;D^{3}),$ (2.3)
and $\Pi^{gh}_{\rho\sigma}(q)$ describes the skeleton loop contribution associated with the ghost degrees of freedom. $\Pi^{(1)}_{\rho\sigma}(q;D^{2})$ represents the skeleton loop contribution, containing the 3-gluon vertices only. Finally, $\Pi^{(2)}_{\rho\sigma}(q;D^{4})$ and $\Pi^{(2^{\prime})}_{\rho\sigma}(q;D^{3})$ describe the skeleton two-loop contributions, which combine the 3- and 4-gluon vertices. All these quantities are given by the corresponding skeleton loop diagrams in Fig. 1, and they are independent from each other. The last four terms explicitly contain the full gluon propagators in the corresponding powers symbolically shown above. They form the NL part of the gluon SD equation. The analytical expressions for the corresponding skeleton loop integrals 22 , in which the symmetry and combinatorial coefficients and signs have been included, are not important here. We are not going to calculate any of them explicitly, and thus to introduce into them any truncations/approximations/assumptions or choose some special gauge. Any skeleton loop integral in Fig. 1 is the sum of the infinite number of terms. Moreover, the full vertices entering these skeleton loop integrals are themselves determined by an infinite series of the corresponding multi-loop skeleton terms 2 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 . In these NL series the dependence on the coupling constant may not be simple, i.e., in fact, it is not known. Such kind of series are the so-called ’cluster’ expansions 23 , indeed.
The constant tadpole term $\Pi_{\rho\sigma}^{t}(D)$ is defined as follows:
$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lD_{\alpha\beta}(l)T^{0}_{\rho\sigma\alpha\beta}=g_{\rho\sigma}\Delta^{2}_{t}(D)=[T_{\rho\sigma}(q)+L_{\rho\sigma}(q)]\Delta^{2}_{t}(D),$ (2.4)
where $L_{\rho\sigma}(q)=q_{\rho}q_{\sigma}/q^{2}$, so that the tadpole term contributes into the both transverse and longitudinal components of the full gluon propagator (2.1). In what follows the constant $\Delta^{2}_{t}(D)$ will be called as the tadpole term itself, since it deviates from the definition (2.4) by the tensor $g_{\rho\sigma}$ only.
It is instructive for further purpose to present the initial gluon SD eq. (2.1), on account of the expressions (2.2) and (2.4) as follows:
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)$ $+$ $D^{0}_{\mu\rho}(q)i\left[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)\right]D_{\sigma\nu}(q)$ (2.5)
$+$ $D^{0}_{\mu\rho}(q)ig_{\rho\sigma}\Delta^{2}_{t}(D)D_{\sigma\nu}(q).$
All the terms which contribute to the full gluon self-energy eq. (2.2), and hence eq. (2.5), are tensors, having the dimensions of a mass squared. All these skeleton loop integrals are therefore quadratically divergent (QD), i.e., ultraviolet (UV) divergent in the perturbation theory (PT) regime, and so they are assumed to be regularized. Contrary to QED, QCD being a non-abelian gauge theory can suffer from the severe infrared (IR) singularities in the $q^{2}\rightarrow 0$ limit (i.e., more singular than the PT one, which is proportional to $\sim 1/q^{2}$). Evidently, they will be generated due to the self-interaction of massless gluon modes. Thus, all the possible subtractions at zero may be dangerous 2 ; 3 ; 4 ; 5 ; 6 ; 7 . That is why in all the quantities below, the dependence on the finite (slightly different from zero) dimensionless subtraction point $\alpha$ has to be understood. In other words, all the subtractions at zero and the Taylor expansions around zero should be understood as the subtractions at $\alpha$ and the structure of the Taylor expansions near $\alpha$, where they are justified to be used. From the technical point of view, however, it is convenient to put formally $\alpha=0$ in all the expressions and derivations below, and to restore the explicit dependence on non-zero $\alpha$ in all the quantities only at the later stage. In all the quantities where the dependence on the dimensionless UV regulating parameter $\tilde{\lambda}$ and $\alpha$ is not shown explicitly, nevertheless, it should be also assumed. For example, $\Pi_{\rho\sigma}(q;D)\equiv\Pi_{\rho\sigma}(q;D,\tilde{\lambda},\alpha)$ and similarly for all other quantities. This means that all the expressions are regularized (i.e., we render infinite skeleton loop integrals to be finite).
The tadpole term $\Delta^{2}_{t}(D)$, defined in eq. (2.4), is nothing else but the QD skeleton loop integral, which is already regularized from below and above as well as all the other such kind of constants which may appear in the theory, as a results of the proper subtraction scheme (see below). Being a mass squared regularized quantity, it is explicitly present in the QCD ground state, see eq. (2.5). However, it violets the gauge symmetry of the QCD Lagrangian. Then a natural question arises why is it present in the vacuum of QCD at all if it makes theory to look un- renormalizable from the very beginning? The standard solution to this dilemma was to remove this and all other QD constants in order to make theory renormalizable in accordance with the QCD Lagrangian’s exact gauge symmetry. However, here we are going to solve this dilemma by asking the question is it possible to retain the tadpole term in the QCD vacuum, but without affecting the PT renormalization properties of the theory? This is important to ask because, unlike to other QD constants, the tadpole term is only one which generates a mass squared scale parameter 6 . Also, one of our primary aims will be to clarify its role in the dynamical and gauge structures of the QCD vacuum in more detail. For this purpose we will develop a suitable formalism, based only on the rigorous rules of the tensor algebra derivations.
Concluding, let us remind that the whole gluon momentum range is $q^{2}\in[0,\infty)$. In what follows we will work in Euclidean metric $q^{2}=q^{2}_{0}+\bf q^{2}$ since it implies $q_{i}\rightarrow 0$ when $q^{2}\rightarrow 0$ and vice versa. This makes it possible to avoid the unphysical IR singularities on the light cone.
3 Slavnov-Taylor identities
The basic relation to which the full gluon propagator should satisfy is the corresponding Slavnov-Taylor (ST) identity 2 ; 24 ; 25 , namely
$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi,$ (3.1)
where $\xi$ is the gauge-fixing parameter. It implies that the general tensor decomposition of the full gluon propagator in the covariant gauge is as follows:
$D_{\mu\nu}(q)=i\left[T_{\mu\nu}(q)d(q^{2})+\xi L_{\mu\nu}(q)\right]{1\over q^{2}},$ (3.2)
where the invariant function $d(q^{2})$ is the corresponding Lorentz structure of the full gluon propagator (or the gluon invariant function). Throughout this paper we use the standard definition of $T_{\mu\nu}(q)=\delta_{\mu\nu}-q_{\mu}q_{\nu}/q^{2}=\delta_{\mu\nu}-L_{\mu\nu}(q)$ in Euclidean metric. Any invariant functions associated with the projective operators $T_{\mu\nu}(q)$ and $\delta_{\mu\nu}$ are the same, and thus $D_{\mu\nu}(q)$ is defined up to its longitudinal part $L_{\mu\nu}(q)$. This should be also valid for its equation of motion.
If one neglects all the contributions to the full gluon self-energy in eq. (2.1), by setting $d(q^{2})=1$ in eq. (3.2), then one obtains the free gluon propagator
$D^{0}_{\mu\nu}(q)=i\left[T_{\mu\nu}(q)+\xi_{0}L_{\mu\nu}(q)\right](1/q^{2}),$ (3.3)
where $\xi_{0}$ is the corresponding gauge-fixing parameter. The general ST identity (3.1) will look like
$q_{\mu}q_{\nu}D^{0}_{\mu\nu}(q)=i\xi_{0}.$ (3.4)
Let us note in advance that from the very beginning the gauge-fixing parameter for the full gluon propagator $\xi$ is not to be identified with its free counterpart $\xi_{0}$. Under which conditions they are the same or different is one of the main subjects of our investigation in this paper. Since the gauge freedom in QCD is determined by $\xi_{0}$, the gauge-fixing parameter $\xi$ has to be a function of $\xi_{0}$ i.e., $\xi=f(q^{2};\xi_{0})$ and called the gauge-fixing function.
Emphasizing, these and all other ST identities which appear in QCD 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 9 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 are important for the general renormalizability of the theory, and thus they ”are exact constraints on any solution to QCD” 2 .
3.1 Transverse relations for the full gluon self-energy
Let us now perform some algebraic derivations which are necessary for the further purposes. Contracting the full gluon SD eq. (2.1) and eq. (2.5) with $q_{\mu}$ and $q_{\nu}$, on account of the relations (3.1)-(3.4), one gets
$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)$ $=$ ${(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2},$
$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]$ $=$ ${(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}-q^{2}\Delta^{2}_{t}(D),$ (3.5)
and they related to each other through the relation (2.2) as well as eq. (2.1) is related to eq. (2.5), which is as it should be. The first line in the transverse relations (3.5) is the tensor relation, while its second one is nothing else but the tensor identity. The second transverse relation explicitly depends on the tadpole term $\Delta^{2}_{t}(D)$. If it is formally put zero, then the both transverse relations (3.5) coincide with each other. This underlines its important role in the QCD ground state, which one precisely is to be clarified in our paper.
[…]
However, by this mechanism the QD constants due to the quark degrees of freedom only can be excluded from the QCD theory. What is about the exclusion of the QD constants from its YM sector? QCD being much more complicated quantum field gauge theory than QED, has the two transverse relations (3.5), second of which shows the gluon degrees of freedom explicitly. By solution of these relations, we mean that at least one of them has to be necessarily satisfied. Finally, this will lead to the removal of the corresponding QD constants from the YM theory as well as to fix the function $\xi=f(q^{2};\xi_{0})$, see exact derivations in sections 4, 5 and 6. This strategy is definitely accessible with conventional means of the gauge theories to make them renormalizable, and QCD is a such theory 2 , indeed. So that whether these transverse relations will be satisfied/saturated (i.e., equal zero) or not and further include them into the proper subtraction scheme requires much more careful investigation in QCD. Obviously, there exist only the two different requirements in order to satisfy at least one of the transverse relations (3.5). However, it is time now to introduce the proper subtraction scheme.
4 The proper subtraction scheme
The first step in the renormalization program of any gauge theory is the removal of the quadratic UV divergences in order to make the corresponding theory renormalizable in the PT sense. It can be achieved by introducing the proper subtraction scheme in order to separate them from the PT logarithmic divergences. The preliminary step in the regularization program has been already done by introducing the corresponding regulating parameters $\tilde{\lambda}$, $\alpha$.
Within our approach nothing will depend on how exactly the regulating parameters have been introduced. Even the hard cut-off procedures can be used, since we are mainly interested in the regulating of leading order divergences 2 . They have to disappear from the theory after the PT and the non-perturbative (NP) renormalization programs will be performed.
Let us now introduce the subtractions for the quark and gluon contributions to the full gluon self-energy as follows:
$\Pi^{s(q)}_{\rho\sigma}(q)$ $=$ $\Pi^{q}_{\rho\sigma}(q)-\Pi^{q}_{\rho\sigma}(0)=\Pi^{q}_{\rho\sigma}(q)-\delta_{\rho\sigma}\Delta^{2}_{q},$
$\Pi^{s(g)}_{\rho\sigma}(q;D)$ $=$ $\Pi^{g}_{\rho\sigma}(q;D)-\Pi^{g}_{\rho\sigma}(0;D)=\Pi^{g}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{g}(D),$ (4.1)
and thus $\Pi^{s(q)}_{\rho\sigma}(0)=\Pi^{s(g)}_{\rho\sigma}(0;D)=0$, by definitions. In these relations $\Delta^{2}_{q}$ is the skeleton loop integral for the quark degrees of freedom at $q=0$, and $\Delta^{2}_{g}(D)$ is the sum of the corresponding skeleton loop integrals at $q=0$ contributing to eq. (2.3), namely
$\Delta^{2}_{g}(D)=\Delta^{2}_{gh}+\Delta^{2}_{1}(D^{2})+\Delta^{2}_{2}(D^{4})+\Delta^{2}_{2^{\prime}}(D^{3}).$ (4.2)
Let us remind that all the QD constants, shown up in eqs. (4.1)-(4.2), as well as the tadpole term $\Delta^{2}_{t}(D)$ itself, are independent from each other and are defined by the skeleton loop integrals, which have been already regularized from above and below. The subtraction at zero is to be understood in a such way that we subtract at the external gluon momentum $q^{2}=-\mu^{2}$ 2 with $\mu^{2}\rightarrow 0$ final limit (its dimensionless counterpart is $\alpha$ mentioned above). However, due to the self-interaction of the massless gluon modes, some of these constants (apart from the tadpole term) may depend on the internal zero gluon momentum (if it is connected to any closed loop, for example see appendix A). Then it has to be replaced by the subtraction at $q^{2}_{i}=-M^{2}_{i}$ with $M^{2}_{i}\rightarrow 0$ final limit as well, and subindex $i$ determines the number of such internal gluons. For the any full gluon propagator the subindex $i$ can be treated as the number of the necessary subtractions made in it in accordance with the rules of the theory of distributions (generalized functions) 26 .
Also it is necessary to emphasize that the subtractions (4.1) have not been introduced by hand. They are equivalent to add zero to the corresponding identity. For example $\Pi^{q}_{\rho\sigma}(q)=\Pi^{q}_{\rho\sigma}(q)-\Pi^{q}_{\rho\sigma}(0)+\Pi^{q}_{\rho\sigma}(0)$, so that the initial quark contribution to the gluon self-energy is not changed. It makes it possible to present the initial quark contribution $\Pi^{q}_{\rho\sigma}(q)=\Pi^{s(q)}_{\rho\sigma}(q)+\Pi^{q}_{\rho\sigma}(0)$ as a sum of the two terms, one of which shows up all the corresponding QD but regularized constants. The first term is a regular function of the external gluon momentum $q$, by definition, and it may be only logarithmically divergent at large $q$. The same is true for the initial gluon contribution as well. Such an exact separation will be very useful for the renormalization programs of any kind to be performed for the single full gluon propagator, and we consider this as an advantage of the formalism developed here. For the interacting full gluon propagators these independent terms will interact with each other, but such a separation will simplify the corresponding renormalization programs, anyway.
The independent tensor decompositions of the quark and gluon degrees of freedom, which appear in the subtraction relations (4.1), are
$\Pi^{s(q)}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(q)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(q)}_{l}(q^{2}),$
$\Pi^{q}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{q}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{q}_{l}(q^{2}),$ (4.3)
and
$\Pi^{s(g)}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(g)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(g)}_{l}(q^{2}),$
$\Pi^{g}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{g}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{g}_{l}(q^{2}),$ (4.4)
respectively. In all the quantities above and below the dependence on $D$ is omitted, for simplicity, and will be restored when necessary. Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary. However, all the invariant functions $\Pi^{s(q)}_{t}(q^{2})$, $\Pi^{s(q)}_{l}(q^{2})$ as well as $\Pi^{s(g)}_{t}(q^{2})$, $\Pi^{s(g)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the subtractions (4.1).
Substituting all these decompositions (4.3)-(4.4) into the subtractions (4.1), and doing some tensor algebra derivations, one finally obtains
$\Pi^{s(q)}_{t}(q^{2})=\Pi^{q}_{t}(q^{2})-{\Delta^{2}_{q}\over q^{2}},\quad\Pi^{s(q)}_{l}(q^{2})=\Pi^{q}_{l}(q^{2})+{\Delta^{2}_{q}\over q^{2}},$ (4.5)
and
$\Pi^{s(g)}_{t}(q^{2})=\Pi^{g}_{t}(q^{2})-{\Delta^{2}_{g}(D)\over q^{2}},\quad\Pi^{s(g)}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})+{\Delta^{2}_{g}(D)\over q^{2}}.$ (4.6)
Using these relations it is easy to show that for the regularized (i.e., finite) quark and gluon contributions to the initial gluon self-energy their values at zero always are $\Pi^{q}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{q}$ and $\Pi^{g}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{g}(D)$, indeed, i.e, they are general ones, and have been already used in the subtraction relations (4.1) above, as it has to be. Also these relations will be true if expressed in terms of the above-mentioned subtraction point $q^{2}=-\mu^{2}$.
Let us now make an important observation in advance. If the initial contributions of the quarks and gluons degrees of freedom to the full gluon self-energy, which are present by the second lines in the relations (4.3) and (4.4), respectively, are not transverse, then one cannot remove the corresponding QD constants $\Delta^{2}_{q}$ and $\Delta^{2}_{g}(D)$ from the theory. On the contrary, if their are transverse, i.e., $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$, then from the relations (4.5) and (4.6) it follows that one has to put the corresponding QD constants $\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0$, and thus removing them from the theory, since the corresponding subtracted invariant functions cannot have the pole-type singularities, by definition. This observation explicitly shows the necessity to satisfy one of the transverse relations (3.5) from the general point of view.
An interesting observation follows from the relations (4.5) and (4.6) as well, namely $\Pi^{q}_{t}(q^{2})+\Pi^{q}_{l}(q^{2})=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(q)}_{l}(q^{2})$ and $\Pi^{g}_{t}(q^{2})+\Pi^{g}_{l}(q^{2})=\Pi^{s(g)}_{t}(q^{2})+\Pi^{s(g)}_{l}(q^{2})$. So that the corresponding sums do not depend on the corresponding QD but regularized constants, which once more underlines the general character of our subtraction scheme. Also, the sum of the subtracted invariant functions is never zero, though the subtracted contributions to the full gluon self-energy are zero at $q^{2}=0$, by definitions, see relations (4.1).
5 Gauge symmetry is unbroken
In order to satisfy the first transverse relation (3.5) one needs to put $\xi=f(\xi_{0})=\xi_{0}$, and thus it becomes transverse $q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)=0$, indeed. Then from the second of the relations (3.5) it follows that
$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]=-q^{2}\Delta^{2}_{t}(D),$ (5.1)
and it is not transverse, as expected. Substituting further the subtractions (4.1) into the relation (5.1), and doing some algebra, one obtains
$q_{\rho}q_{\sigma}[\Pi^{s(q)}_{\rho\sigma}(q)+\Pi^{s(g)}_{\rho\sigma}(q;D)]=-q^{2}\Delta^{2}(D),$ (5.2)
where
$\Delta^{2}(D)=\Delta^{2}_{q}+\Delta^{2}_{g}(D)+\Delta^{2}_{t}(D),$ (5.3)
is the sum of all the QD but regularized constants, which may appear in the theory.The QD but regularized constant $\Delta^{2}_{g}(D)$ itself is defined in eq. (4.2).
Substituting further the first tensor decompositions of the relations (4.3) and (4.4) into the transverse relation (5.2), and doing some algebra, one obtains for their longitudinal invariant functions the following relation
$\Pi^{s(q)}_{l}(q^{2})+\Pi^{s(g)}_{l}(q^{2})={\Delta^{2}(D)\over q^{2}},$ (5.4)
which is impossible, since these invariant functions cannot have the pole-type singularities, by definition, as underlined just above. So that the QD but regularized constant $\Delta^{2}(D)$ should be discarded from the theory, i.e., put formally zero $\Delta^{2}(D)=0$ on this mathematical basis. Then $\Pi^{s(q)}_{l}(q^{2})=\Pi^{s(g)}_{l}(q^{2})=0$ as well, since they are independent from each other.
From now on in this section, we replace $D$ by $D^{PT}$ (for the explanation see the text below). The definitions (4.2) and (5.3) require that this system of the relations for the QD but regularized constants now looks like
$\Delta^{2}_{q}=\Delta^{2}_{g}(D^{PT})=\Delta^{2}_{t}(D^{PT})=0,$ (5.5)
and
$\Delta^{2}_{gh}=\Delta^{2}_{1}(D^{2}_{PT})=\Delta^{2}_{2}(D^{4}_{PT})=\Delta^{2}_{2^{\prime}}(D^{3}_{PT})=0,$ (5.6)
because all of them are independent from each other. The dimensional regularization method (DRM) 4 ; 6 ; 27 ; 28 ; 29 provides a gauge-invariant scheme to correctly calculate the finite parts of the generally QD loop integrals, while omitting their QD parts, i.e., simply to ignore them and to deal further only with the logarithmic divergences of the PT. This was a prescription rather than an exact result, as pointed out in 4 . But the relations (5.5)-(5.6) are not prescriptions, they are mathematically exact results. How to understand such equalities is convenient to postpone until next section.
Collecting now all our results obtained above, one arrives at the system of the relations as follows:
$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D^{PT})=0,$ (5.7)
and
$\xi=f(\xi_{0})=\xi_{0},\quad\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D^{PT})=T_{\rho\sigma}(q)q^{2}\Pi^{s}(q^{2};D^{PT}),$ (5.8)
where $\Pi^{s}(q^{2};D^{PT})=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(g)}_{t}(q^{2};D^{PT})$, since the quark and gluon contributions coincide with their subtracted counterparts and are transverse in this case as well, i.e., $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$, due to the second equations in the relations (4.5)-(4.6). The invariant function $\Pi^{s}(q^{2};D^{PT})$ is a regular function at small $q^{2}$ and may be only logarithmically divergent at large $q^{2}$.
The gluon SD eq. (2.5) becomes
$D^{PT}_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)q^{2}\Pi^{s}(q^{2};D^{PT})D^{PT}_{\sigma\nu}(q),$ (5.9)
and from it one arrives at $q_{\mu}q_{\nu}(q)D^{PT}_{\mu\nu}(q)=q_{\mu}q_{\nu}D^{0}_{\mu\nu}(q)=i\xi_{0}$, and thus the gauge-fixing parameter of the PT full gluon propagator is fixed as follows: $\xi=\xi_{0}$, in complete agreement with eqs. (5.8). Substituting decompositions (3.2) and (3.3) into the eq. (5.9), one finally obtains
$D^{PT}_{\mu\nu}(q)={iT_{\mu\nu}(q)\over q^{2}[1+\Pi^{s}(q^{2};D^{PT})]}+i\xi_{0}L_{\mu\nu}(q){1\over q^{2}},$ (5.10)
so that the gluon invariant function is $d_{PT}(q^{2})=[1+\Pi^{s}(q^{2};D^{PT})]^{-1}$. Obviously, this equation describes the propagation of the PT massless gluons, since it has the PT singularity on the mass-shell $q^{2}=0$, i.e., the singularity of the free gluon propagator $\sim 1/q^{2}$. Let us note in advance that the equality $\xi=\xi_{0}$ takes place only for the regularized massless gluon fields. For their renormalized counterparts these gauge-fixing parameters differ by the corresponding renormalization constant 2 ; 3 ; 4 ; 5 ; 6 ; 7 .
The absence of a mass scale parameters in eq. (5.10) can be now attributed to the satisfied transverse relations (5.7)-(5.8). It is possible to say that just these ones decrease the quadratic UV divergences of the corresponding skeleton loop integrals to a logarithmic ones at large $q^{2}$. Due to the transverse relations (5.7)-(5.8) all the scale parameters, having the dimensions of a mass squared, shown up in eqs. (5.5)-(5.6), should be removed/disappeared from the theory, i.e., put formally zero.
The system of the regularized eqs.(5.9)-(5.10), which is uniquely defined, is free of all the types of the scale parameters having the dimensions of a mass squared, forbidden by the exact gauge symmetry of the QCD Lagrangian. Therefore, it constitutes that the gauge symmetry of the QCD ground state, reflected by this system, coincides with the symmetry of its Lagrangian. Such coincidence is analogous to QED where the abelian $U(1)$ gauge symmetry of the Lagrangian is the same as of its ground state (omitting formally the YM degrees of freedom here, this coincidence has been proven just in this way in 10 ). However, it is necessary to remind that the coupling constant remains strong, so to call it as $D^{PT}$ is only a mere convention.
The essential feature of the above-mentioned coincidence is the equality $\xi=\xi_{0}$. In this connection, let us note that one can start from the expression (3.2). It is the general decomposition of the full gluon propagator, depending on the one variable, into the independent tensor structures. Then it will automatically satisfy to the relation (3.1), called the ST identity, determining thus the gauge-fixing parameter. This has nothing to do with the statement that the ST identity is a consequence of the exact gauge symmetry. It becomes its consequence, indeed, when $\xi$ is fixed by the relation (5.8). The well-known PT linear relation $\xi=\xi_{0}$ is the consequence of the exact gauge invariance both for the QCD Lagrangian and its ground state if it is treated in the PT framework for the regularized gluon fields. In principle, the PT QCD is a well-elaborated quantum gauge field theory 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 , and it is not our concern. We present it here within our formalism for the readers convenience to directly compare it with our approach to the QCD below.
6 Gauge symmetry is broken!
QCD, being a non-abelian gauge theory with the strong coupling constant, is to be treated beyond the PT, especially its ground state. It contains much more information on the dynamical and gauge structures of the theory than its Lagrangian can provide at all 2 . So that the general case when $\xi\neq\xi_{0}$ should be investigated in detail, while keeping the ST identities (3.1) and (3.4) valid, anyway. The another independent requirement to satisfy the second one of the transverse relations (3.5) is defined by the condition
${(\xi_{0}-\xi)\over\xi\xi_{0}}q^{2}-\Delta^{2}_{t}(D)=0.$ (6.1)
Then for the initial gluon self-energy its transverse relation looks like
$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)=q^{2}\Delta^{2}_{t}(D),$ (6.2)
and it is not transverse, while the second line of the system (3.5) becomes transverse,
$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]=0,$ (6.3)
as expected and compare it with eq. (5.1). In the previous section we do not use that the contribution of the quark skeleton loop to the full gluon self-energy can be made always transverse, i.e., put $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=0$, independently from its gluon counterpart (this is an analogue to the satisfied transverse relation for the vacuum polarization tensor in QED, mentioned above). So that the satisfied transverse relation (6.3) can be effectively splintered into the two independent transverse relations, namely
$q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=0,\quad q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0.$ (6.4)
However, from the second lines of the tensor decompositions (4.3)-(4.4) and the satisfied transverse relations (6.4), one arrives to the conditions $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$, like in the previous case. Then the second relations in eqs (4.5) and (4.6) are to be reduced to
$\Pi^{s(q)}_{l}(q^{2})={\Delta^{2}_{q}\over q^{2}},\quad\Pi^{s(g)}_{l}(q^{2};D)={\Delta^{2}_{g}(D)\over q^{2}},$ (6.5)
but that is impossible, since the corresponding invariant functions cannot have pole-type singularities, by definitions, as explained in the previous section. One has to put these constants to zero on this general mathematical basis, i.e., $\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0$, and $\Pi^{s(q)}_{l}(q^{2})=\Pi^{s(g)}_{l}(q^{2};D)=0$ as well.
All our results for the QD constants can be now summarized as follows:
$\Delta_{q}^{2}=\Delta_{g}^{2}(D)=0,\quad\Delta^{2}(D)\rightarrow\Delta_{t}^{2}(D)\neq 0,$ (6.6)
and hence
$\Delta^{2}_{gh}=\Delta^{2}_{1}(D^{2})=\Delta^{2}_{2}(D^{4})=\Delta^{2}_{2^{\prime}}(D^{3})=0,$ (6.7)
as well because of the relation (4.2). It is instructive to compare this system of the relations with the relations (5.5)-(5.6). The common feature of these relations is that the quark and gluon constants should be removed from the theory not only in the PT QCD but in the QCD as well. This means that the correspondence with the PT system (5.5)-(5.6) is maintained by the system (6.6)-(6.7). It is important for the renormalizability of the theory in the PT sense. The difference is that in the QCD the tadpole term remains intact. The essential dynamical source of this difference is that only the tadpole term can generate a mass squared scale parameter. All the other terms present in the skeleton loops decomposition for the full gluon self-energy, seen in Fig. 1, cannot do this 6 , and thus none of their subtracted counterparts with their quark and gluon constants.The system of the relations (6.6)-(6.7) is not only a general but it is a unique one as well. They are not prescriptions, since based on a rigorous tensor algebra derivations, i.e., they are exact mathematical results. The relations (6.6) hold because these constants are related to the corresponding invariant functions for the subtracted counterparts, as described above and shown in the relations (6.5). Evidently, such kind of the relation does not exist for the constant tadpole term $\Delta^{2}_{t}(D)$, that is why it cannot be disregarded at all on this mathematical basis.
All these constants, defined by the corresponding skeleton loop integrals, are QD at the upper limit. They have to be removed/disregarded from the theory on the general basis, i.e., formally put zero, as described above, apart from the tadpole term (2.4). It should remain intact in the QCD full gluon propagator, while in the PT QCD full gluon propagator it has to be removed as well. In what follows all these constants will be called the tadpole-like/type terms, for simplicity. The general question arises now, namely how to understand these exact equalities to zero in the relations (6.6)-(6.7), and hence in the relations (5.5)-(5.6) as well? Each tadpole-like term in these relations is the skeleton loop integral, which thus is a sum of the infinite number of terms. But an equality does not mean that the sum itself is zero. These equalities mean that any tadpole-like term which may appear in the theory by any possible way has to be discarded/disregarded in the theory, i.e, put formally zero, independently from any other tadpole-like terms. In this way such sums will be always zero, indeed. Such constants may appear even not as the result of the subtractions, but, for example as the result of the NL iteration procedure for the full gluon propagator shown in Fig. 1. In this case the tadpole-like terms may be even multiplied by some regularized functions, see appendix A below. So that, whatever their origins would have been in the full gluon propagator, all of them belong to the infinite manifold of the relations (6.6)-(6.7).
Collecting all our results obtained above for the quark and gluon contributions to the gluon self-energy, one obtains
$\Pi^{q}_{\rho\sigma}(q)=T_{\rho\sigma}(q)q^{2}\Pi^{s(q)}_{t}(q^{2}),\quad\Pi^{g}_{\rho\sigma}(q;D)=T_{\rho\sigma}(q)q^{2}\Pi^{s(g)}_{t}(q^{2};D),$ (6.8)
since they become transverse and coincide with their subtracted counterparts. Let us also remind that the invariant functions present in the previous relations are regular functions at small $q^{2}$, i.e., they have no pole-type singularities, and may be only logarithmically divergent at large $q^{2}$. Substituting their sum into the eq. (2.5), one gets
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)$ $+$ $D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)q^{2}\Pi^{s}_{t}(q^{2};D)D_{\sigma\nu}(q)$ (6.9)
$+$ $D^{0}_{\mu\rho}(q)i\delta_{\rho\sigma}\Delta^{2}_{t}(D)D_{\sigma\nu}(q),$
where $\Pi^{s}_{t}(q^{2};D)=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(g)}_{t}(q^{2};D)$.
The corresponding ST identities therefore are not equal to each other, namely $q_{\mu}q_{\nu}(q)D_{\mu\nu}(q)\neq q_{\mu}q_{\nu}D^{0}_{\mu\nu}(q)$, which implies $\xi\neq\xi_{0}$. The gauge-fixing parameter $\xi$ as a function of $\xi_{0}$ is given now by the condition (6.1). Contrary to the gluon SD eq. (5.9), the gluon SD eq. (6.9) is now equivalent to the initial gluon SD eq. (2.5), only being re-written in the different form. That is the reason why we left it denoted as $D$ here and everywhere below. Evidently, the PT gluon SD eq. (5.9) is not equivalent to eq. (2.5). The both full gluon SD eqs. (2.5) and (6.9) by keeping the mass squared parameter explicitly - the tadpole term $\Delta^{2}_{t}(D)$ constitutes the fact that the gauge symmetry of the QCD ground state is not a symmetry of its Lagrangian. The PT renormalizability of the theory seems to be compromised, however see arguments below.
It is important to emphasize once more that we are under an obligation to satisfy at least one of the transverse relations (3.5). Reminding that this is necessary to do in order to get finally to eq. (5.9) in the PT QCD and eq. (6.9) in the QCD. The last one allows to distinguish between the quark $\Delta^{2}_{q}$ and the gluon $\Delta^{2}_{g}(D)$ constants, which are not present in the initial gluon SD eq. (2.5) and the tadpole constant $\Delta^{2}_{t}(D)$, which is explicitly present in it. The satisfied transverse relations have been implemented into the our formalism in a self-consistent way. The fact of the matter is that the system of the relations (6.2)-(6.3) and (6.8)-(6.9) have been uniquely and exactly derived. Then the condition (6.1) can be recovered from the gluon SD eq. (6.9).
7 The QCD full gluon propagator in the generalized gauge
The basic relations to which the full gluon propagator and its free counterpart should satisfy are the corresponding ST identities (3.1) and (3.4), respectively. As emphasized above in section 3, they are important for the general renormalizability of the theory, and thus provide the exact constraints on any solution to QCD 2 . But by themselves they cannot remove the quadratic UV divergences from the theory, as it has been shown above in the relations (6.6)-(6.7), apart from the tadpole term. We have achieved this by formulating the proper subtraction scheme in which the corresponding transverse conditions have been implemented. If some equations, relations or the regularization schemes, etc. do not satisfy the ST identities automatically, i.e., without any additional conditions, then they should be modified and not the identity (3.1). In other words, all the relations, equations, regularization schemes, etc. should be adjusted to it and not vice versa. For example, the transverse condition for the full gluon self-energy (6.2) is not satisfied, but, nevertheless, the general form of the ST identity (3.1) is to be maintained in any case, see eq. (7.3) below. Let us begin with the gluon SD eq. (6.9), which can be equivalently re-written as follows:
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)$ $+$ $D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)\left[q^{2}\Pi(q^{2};D)+\Delta^{2}_{t}(D)\right]D_{\sigma\nu}(q)$ (7.1)
$+$ $D^{0}_{\mu\rho}(q)iL_{\rho\sigma}(q)\Delta^{2}_{t}(D)D_{\sigma\nu}(q),$
saving the tadpole term in the longitudinal component of eq. (6.9), it makes it possible for the tadpole term to appear in the transverse component of the full gluon propagator (7.1) as well. The invariant function $\Pi(q^{2};D)$ is a replacement (for simplicity) of the initial one $\Pi^{s}_{t}(q^{2};D)$ in eq. (6.9). It is worth remanding that it is regular at zero and may have only logarithmic divergences in the PT $q^{2}\rightarrow\infty$ limit. Combining this equation with the decompositions (3.2) and (3.3), one obtains
$d(q^{2})={1\over 1+\Pi(q^{2};D)+(\Delta^{2}_{t}(D)/q^{2})}.$ (7.2)
This relation is not a solution for the full gluon invariant function. It is the NL transcendental equation for the different invariant functions $d(q^{2}),\ \Pi(q^{2};D)$ and the constant $\Delta^{2}_{t}(D)$, i.e., $d=f(D(d))$. Nevertheless, from this expression is clearly seen that in the PT $q^{2}\rightarrow\infty$ regime, the contribution $(\Delta^{2}_{t}(D)/q^{2})$ can be neglected, but the invariant function $\Pi(q^{2};D)$ may still depend on this ratio under the PT logarithms. At the same time, in the NP region of finite and small gluon momenta this term is dominant, and the dependence of $d(q^{2})$ on $\Delta^{2}_{t}(D)$ may be much more complicated due to the transcendental character of eq. (7.2).
Contracting the full gluon SD eq. (7.1) with $q_{\mu}$ and $q_{\nu}$, and substituting its result into the general ST identity (3.1), one arrives at
$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi_{0}\left(1-\xi{\Delta^{2}_{t}(D)\over q^{2}}\right)=i\xi,$ (7.3)
which solution is
$\xi\equiv\xi(q^{2};\xi_{0})={\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)},$ (7.4)
i.e., in this case the gauge-fixing parameter becomes the function $\xi\equiv\xi(q^{2};\xi_{0})$ and thus it is not a constant equal to $\xi_{0}$. Let us point out that the expression (7.4) satisfies condition (6.1), which is as it should be. It is worth emphasizing that the solution (7.4) shows the NL dependence on $\xi_{0}$, which is only possible if gauge symmetry is broken in the ground state. From eq. (7.4) it is clearly seen that in the PT $q^{2}\rightarrow\infty$ regime, which effectively equivalent to the formal $\Delta^{2}_{t}(D)=0$ limit and vice versa, the gauge-fixing parameter $\xi$ goes to $\xi_{0}$, as expected. The generalized gauge (7.4) is the known function of its arguments, while the relation (7.2) is the NL transcendental one. Thus, behind the general inequality $\xi\neq\xi_{0}$ is the constant $\Delta^{2}_{t}(D)$ as its dynamical source, when its contribution can be neglected only then $\xi=\xi_{0}$. The equality holds only for the regularized gluon fields, as underlined above in section 5.
Substituting equations (7.2) and (7.4) into the general decomposition (3.2) for the full gluon propagator, one finally obtains
$D_{\mu\nu}(q)={iT_{\mu\nu}(q)\over q^{2}+q^{2}\Pi(q^{2};D)+\Delta^{2}_{t}(D)}+iL_{\mu\nu}(q){\lambda^{-1}\over q^{2}+\lambda^{-1}\Delta^{2}_{t}(D)},$ (7.5)
[…]
The corresponding ST identity now becomes
$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi(q^{2};\lambda^{-1})=i{\lambda^{-1}q^{2}\over(q^{2}+\lambda^{-1}\Delta^{2}_{t}(D))}.$ (7.6)
This is the generalized gauge since it depends on the constant $\Delta^{2}_{t}(D)$, and when it is zero, one recovers the gauge-fixing parameter for the free gluon propagator. To our best knowledge the generalized gauge (7.6) has been derived for the first time in a such new manner. In the generalized gauge the gauge-fixing parameter $\lambda^{-1}$ is convenient to vary continuously from zero to infinity. The functional dependence of the generalized gauge-fixing parameter $\xi$ (7.4) is fixed up to an arbitrary gauge-fixing parameter $\xi_{0}=\lambda^{-1}$. Unless we fix it, and thus $\xi$ itself, we will call such situation as the general gauge dependence (GGD), see equations (3.1), (7.5), and (7.6). Choosing $\xi_{0}=\lambda^{-1}$ explicitly, we will call such situation as the explicit gauge dependence (EGD). For example, $\xi_{0}=\lambda^{-1}=0$ is called the unitary (Landau) gauge, $\xi_{0}=\lambda^{-1}=1$ is called the t’ Hooft–Feynman gauge, etc. 30 ; 31 ; 32 ; 33 . The formal $\xi_{0}=\lambda^{-1}=\infty$ limit is called as the canonical gauge in 32 . This distinction seems a mere convention, but, nevertheless, it is useful one in QCD because of the presence of a mass scale parameter in its ground state. The generalized gauge requires that there is no other functional expression for $\xi$, apart from given by the relation (7.4) at finite $\xi_{0}=\lambda^{-1}$, in the full gluon propagator (7.5) and the ST identity (7.6) for the regularized gluon fields.
The system of the regularized equations, consisting of the gluon SD eq. (7.1), the expression (7.5) and the ST identity (7.6), constitutes that the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is not a symmetry of its ground state, and it has been exactly and uniquely defined. The tadpole term enters the full gluon self-energy linearly, see the expression (2.2). However, in the full gluon propagator (7.5) it appears in the NL way, because its contribution has been summed up with the help of the gluon SD eq. (7.1). In the PT $q^{2}\rightarrow\infty$ limit, the term $(\Delta^{2}_{t}(D)/q^{2})$ in the expression (7.5) is to be suppressed, and it will be reduced to the PT QCD expression (5.10), reminding that $\xi_{0}=\lambda^{-1}$. The full gluon propagator will behave like the PT QCD gluon propagator at short distance ($q^{2}\rightarrow\infty$), namely $\sim 1/q^{2}$. This is the one of the necessary constraints that a theory is perturbative finite. Other ones such as the corresponding behaviour of the spinor Green’s function, a unitary of $S$-matrix and analyticity (causality) 2 , evidently are beyond the scope of the present paper. Here we have proven that the true dynamical and gauge structures of the QCD/YM ground state are much more complicated than it is required by its Lagrangian’s gauge symmetry. Despite the gauge symmetry is broken in the QCD ground state, its PT renormalizability is not affected within the QCD full gluon propagator in the generalized gauge (7.5), i.e., it has the PT renormalizable behaviour at large $q^{2}$.
8 Discussion
Let us now discuss some interesting features of the explicit presence of the tadpole term in the QCD ground state, i.e., at the fundamental quark-gluon level. First of all, let us underline once more that the tadpole term, existing in the QCD ground state, cannot be discarded in the QCD/YM theory by any means. This was exactly proven in section 6 and used in section 7. Then it will drastically change the structure of the full gluon propagator in the deep IR region $q^{2}\rightarrow 0$. For the sake of transparency and clarity let us use the free gluon propagator in the t’ Hooft-Feynman gauge $\lambda^{-1}=1$, i.e., $D^{0}_{\mu\nu}(q)=i{\delta_{\mu\nu}/q^{2}}$. Then the full gluon propagator (7.1) up to the fist iteration term looks like
$D_{\mu\nu}(q)\sim i{\delta_{\mu\nu}\over q^{2}}-i{T_{\mu\nu}(q)\over q^{2}}\Bigl{[}\Pi(q^{2};D)+{\Delta^{2}_{t}(D)\over q^{2}}\Bigr{]}-iL_{\mu\nu}(q){\Delta^{2}_{t}(D)\over(q^{2})^{2}}+...\ ,$ (8.1)
and it is worth reminding that the invariant function $\Pi(q^{2};D)$ is a regular one of its argument, while all the tadpole-like terms to one loop approximation have been omitted due to the relations (6.6)-(6.7). See appendix A as well, where this iteration is discussed in more detail. Let us also remind that the gluon eq. (7.1) is equivalent to the initial gluon SD eq. (2.1), shown in Fig. 1, and hence to eq. (2.5), only being re-written down in the form suitable for the NP renormalization program to be performed. The higher-order iteration terms will be much more singular with respect to the general ratio $(\Delta^{2}_{t}(D)/q^{2})$, since each iteration invokes additional powers of this ratio. Because of the NL character of the iteration procedure any next iteration will contribute into the previous iteration terms, and thus they become the sums of the infinite number of terms 34 , namely
$D_{\mu\nu}(q)\sim iT_{\mu\nu}{\Delta^{2}_{t}(D)\over(q^{2})^{2}}\sum_{k=0}^{\infty}\Bigl{(}{\Delta^{2}_{t}(D)\over q^{2}}\Bigr{)}^{k}\Phi_{k}(g^{2},\tilde{\lambda}...)+O_{\mu\nu}(1/q^{2}).$ (8.2)
For simplicity, in this iteration expansion is explicitly shown only singular part of the corresponding Laurent series, and thus it represents the summation of all the severe IR singularities (more singular than the PT IR singularity $\sim 1/q^{2}$), which can be accumulated by the full gluon propagator. Evidently, such a summation is only possible due to the existence of the tadpole term. We have shown the singular part, associated with the transverse component of the full gluon propagator, since its longitudinal part is the regular function of the tadpole term, see expression (7.5). The dependence of the coefficients $\Phi_{k}(g^{2},\tilde{\lambda}...)$ on the coupling constant is not known, i.e., such kind of series are the cluster expansions, mentioned above 23 . These coefficients, being the sum of the infinite number of terms by themselves, are resulted from the skeleton loop integrals, contributing into full gluon self-energy through the invariant function $\Pi(q^{2};D)$. Evidently, the iteration ab  infinity leads to the essential singularity at $q^{2}\rightarrow 0$! in the full gluon propagator (8.2), which will remain massless, but depending on the tadpole term. Let us note that the same series can be obtained by expanding the initial NL transcendental relation (7.2) in powers of the ratio $(\Delta^{2}_{t}(D)/q^{2})$.
From the NL iteration expression (8.2) it follows that the full gluon propagator in the IR region ($q^{2}\rightarrow 0$) is dominated by the tadpole contribution, while in the UV region ($q^{2}\rightarrow\infty$) it will be dominated by the free gluon one. This is completely different to QED, where the dressed photon propagator behaves at small photon momentum as its free counterpart ($\sim 1/q^{2}$). This is also true for the full gluon propagator in the PT QCD with unbroken gauge symmetry, investigated in section 5. The principal difference between these theories is due to the existence of the tadpole term in the QCD ground state, which precisely controls the structure of a single full gluon propagator at small $q^{2}$. That is why the dressed PT gluon (5.10) and free gluon states can not appear in the physical spectrum at large distances ($q^{2}\rightarrow 0$), i.e., they will be suppressed in this region by the tadpole term. If this statement will survive the corresponding renormalization program beyond the PT then confinement of the PT and free gluons will be explained. This is important since there is no any other dynamical mechanism to confine these gluon states to the vacuum of QCD or to keep them inside hadrons. To prove the confinement of the singular full gluon propagator itself due to the above-mentioned essential singularity, though possible, but requires completely different investigation within the theory of distributions 26 , and thus is beyond the scope of the present work. Let us note that by the NP program we mean that the full gluon propagator will become free from all the non-physical parameters of the theory in any of its possible solutions (not only in the iteration one).
Eq. (7.5) shows that gluon may acquire mass dynamically if the denominator of this equation becomes zero at some finite point $q^{2}=M^{2}_{g}$. The expression for the free massive vector particle propagator follows in a rather simple way, indeed. For this goal it is necessary to omit the corresponding invariant function and identify the tadpole term with the gluon pole mass $M^{2}_{g}$. In Minkowski metric (by replacing $q^{2}\rightarrow-q^{2}$ and $\delta_{\mu\nu}\rightarrow g_{\mu\nu}$), one gets
$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2}_{g})}\left[g_{\mu\nu}-(1-\lambda^{-1}){q_{\mu}q_{\nu}\over(q^{2}-\lambda^{-1}M^{2}_{g})}\right],$ (8.3)
i.e., it will be expressed in the Stueckelberg gauge 4 . Contracting this expression with respect to $q_{\mu}$ and $q_{\nu}$, one obtains the ST identity (7.6) by replacing there $q^{2}\rightarrow-q^{2}$. In other words, the Stueckelberg gauge (8.3) is a particular case of the generalized gauge (7.6). The PT QCD full gluon propagator (5.10) cannot generate the gluon pole mass at some finite point, since it remains always massless. The full gluon propagator (7.5) provides a rather convenient framework in order to formulate and further develop the NP renormalization program for the massive gluon fields. This program will suppose to show that the full gluon propagator (7.5), and hence its free massive counterpart (8.3), can not be put on the mass-shell $q^{2}=M^{2}_{g}$. Acquiring mass dynamically, the gluon, nevertheless, will remain the off-mass-shell object. So that it will not appear in the physical spectrum (confinement of the massive gluon fields within our approach).
In the presence of the mass squared scale parameter - the tadpole term - the role of the QCD coupling constant $g^{2}$ becomes unimportant. This is also evidence of the ’dimensional transmutation’, $g^{2}\rightarrow\Delta^{2}_{t}(D)$ 2 ; 35 ; 36 , which occurs whenever a massless theory acquires mass dynamically. It is a general feature of spontaneously symmetry breaking in field theories. We distinguish between the PT and full QCD by the explicit presence of the tadpole term in the latter one, and not by the magnitude of the coupling constant even at the regularized gluon fields level yet. In the both cases the gluon fields remain strongly interacted, apart from the asymptotic freedom (AF) regime 37 ; 38 .
Any deviation of the full gluon propagator from the free one in QCD requires the presence of a mass squared scale parameter. Even in the AF regime there is a scale breaking,
$D_{\mu\nu}(q)\sim\delta_{\mu\nu}\Bigl{[}{ig^{2}\over 1+g^{2}b_{0}\ln(q^{2}/\Lambda^{2}_{QCD})}\Bigr{]}(1/q^{2}),\quad q^{2}\rightarrow\infty,$ (8.4)
which has been discovered many years ago 37 ; 38 . When $q^{2}\rightarrow\infty$ the dependence on the coupling constant $g^{2}$ vanishes, indeed. The expression (8.4) is written down in the t’ Hooft-Feynman gauge and in the approximation of the main PT logarithms with the colour group factor $b_{0}>0$. $\Lambda^{2}_{QCD}$ determines the scale of the non-trivial PT dynamics in the QCD. However, in section 5, we have already shown that the PT cannot generate the scale parameter, having the dimension of a mass squared. From the renormalization group equations analysis, it follows that any mass to which can be assigned any physical meaning is exponentially suppressed in the weak coupling limit to every order of the PT 2 ; 35 , namely
$M\sim\mu\exp(-1/b_{0}g^{2}),\quad\quad g^{2}\rightarrow 0,$ (8.5)
where $\mu$ is the arbitrary renormalization point. None a finite mass can survive in the PT weak coupling limit or, equivalently, in the PT $q^{2}\rightarrow\infty$ regime, indeed. So the question where the finite mass comes from? cannot be answered by the PT! It is entirely of the NP origin! It has to come from the IR region which is controlled by the tadpole term, as discusses just above in this section. Its finite version should be somehow related to $\Lambda^{2}_{QCD}$ (reminding that the ratio $(\Delta^{2}_{t}(D)/q^{2})$ may appear under the PT logarithms).
9 Conclusions
Our formalism makes it possible to remove from the theory the QD constants of the gluon degrees of freedom in the same way as the QD constants of the quark degrees of freedom. This has lead to the discovery that in addition to the PT gluon field configurations in the QCD ground state (more precisely in its YM sector), preserving the exact gauge symmetry of its Lagrangian, there also exists such a type of the gluon field configurations which do not maintain this symmetry. But the PT renormalizable properties of the theory will not be affected. The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared. It is dynamically generated by the self-interaction of the multiple massless gluon modes involving only the point-like 4-gluon vertex, see eq. (2.4) and Fig. 1. The full 4-gluon vertex survives when all the gluon momenta involved go to zero, becoming thus its point-like counterpart, while the full 3-gluon vertex does not survive in this limit 10 ; 23 ; 39 . By the design the tadpole term is present in the QCD ground state from the very beginning, i.e., not introducing by hand. It is only one which explicitly contributes into the skeleton loops decomposition of the full gluon propagator (shown in Fig. 1), while all the other tadpole-like terms don’t. Unlike them, being the QD but already regularized constant, the tadpole term is not connected to any external gluon momentum, and thus may generate a mass squared scale parameter 6 , indeed. Thus, there is no need to extract such a scale parameter in the QCD vacuum in a some other way. In 10 the importance of the tadpole term has been recognized and exploited on an intuitive basis rather than a logical one. However, from now on its existence cannot be undermined. Besides its combinatorial (mathematical) meaning it has a deep dynamical one as a source of the gauge symmetry violation in the QCD vacuum. Its renormalized (i.e., finite) version will determine a scale of the NP dynamics in QCD and in its true ground state.
For the both types of the gluon field configurations have been derived the corresponding expressions for the full gluon propagators describing their propagation in the vacuum of QCD. For the second type the corresponding NP QCD full gluon propagator (7.5) is written down in the generalized gauge (7.6). From these expressions it is clearly seen that in the PT $q^{2}\rightarrow\infty$ limit eq. (7.5) coincides with eq. (5.10). This is an evidence that the PT renormalizable properties of the theory will not be affected by the new type of the gluon field configurations. In other words, the renormalizability of the theory will be maintained within our approach to the full gluon propagator. It is worth emphasizing once more that the massive and severely singular gluon field configurations can be described by eq. (7.5) only, while eq. (5.10) describes the propagation of the massless gluon modes with the PT IR singularities $\sim 1/q^{2}$ only. Our ground-breaking findings are as follows:
a. One of the transverse relations (3.5) is to be necessarily satisfied, in order to make the theory renormalizable in the PT sense.
b. The subtraction scheme has been formulated, which makes it possible to remove from theory all the quadratically UV divergent constants on the general basis.
c. The exact gauge symmetry of the YM Lagrangian is not a symmetry of its ground state. The dynamical source of this effect is the tadpole term. If it formally put zero then these gauge symmetries coincide.
d. The system of the relations (6.6)-(6.7) are exact mathematical results, i.e., they are not prescriptions.
e. The new NP approach to the YM theory and its ground state has been formulated, by keeping the corresponding ST identities always valid. It is based on the system of eqs.  (7.5)-(7.6), which is defined up to the corresponding invariant function with the known asymptotic properties. Eqs. (7.5)-(7.6) are suitable for the NP renormalization program to be performed.
f. The NL dependence $\xi$ on $\xi_{0}$ in the generalized gauge (7.4) in the PT $q^{2}\rightarrow\infty$ or, equivalently, formal $\Delta^{2}_{t}(D)=0$ limits becomes linear one $\xi=\xi_{0}$.
g. The DRM 6 ; 25 ; 26 ; 27 prescription to simply ignore all the QD constants (including the tadpole term) in the PT QCD has been put on a firm mathematical ground from now on, see eqs. (5.4)-(5.6).
In obtaining all these results no specific regularization schemes have been used. Neither special gauge choice nor any truncations, approximations and assumptions have been made either in dealing with the regularized skeleton loop integrals, contributing to the gluon self-energy. The exact analytic derivations have been done, based on a rigorous tensor algebra rules, such as the decomposition into the independent structures, the subtractions, etc.
10 Summary
It is shown in the most general and unique way that the gauge symmetry of the QCD Lagrangian is broken in its ground state, indeed. This has been achieved without solving the full gluon SD eq. (2.5), but recovering only its true dynamical and gauge structures in more detail in the framework of our formalism. So that, the significant peace of the information on these important issues was not missed within our general approach. Its breakthrough results can be summarized in terms of the following theorem:
Tadpole term existence and the true QCD ground state: Prove that in the QCD ground state exist such gluon field configurations which break/violet the gauge symmetry of its Lagrangian, but the PT renormalizability of the theory is maintained. Existence of this effect includes establishing the true role of the tadpole term in the QCD ground state.
Just this theorem has been proven above in this paper, which resolves the dilemma formulated at the end of section 2. In the proof of this theorem we have shown how to save the tadpole term in QCD and its ground state, without affecting the PT renormalizable properties of the theory, while removing all the other QD constants from it. The new derived solution with the gauge symmetry breaking will apparently satisfy the bounder conditions in QCD at the fundamental quark-gluon level: confinement at large distances ($q^{2}\rightarrow 0$) and AF at short distances ($q^{2}\rightarrow\infty$), as discussed in section 8. The solution with the gauge symmetry preservation can not satisfy them. All this underlines once more the importance of the theorem above.
The perspectives for further work are as follows: to formulate the NP multiplicative (MP) renormalization programs for the full gluon propagator (7.5), which is only regularized one at this stage. Unfortunately, this is not a simple task, since such kind of a programs have to be different for the massive and severely singular gluon field configurations, briefly discussed in section 7. For this purpose it is necessary to take into account the transcendental and NL structure of eq. (7.2). They are also needed to be formulated in the most general and unique ways. To ensure a unitary of $S$-matrix elements by these NP MP renormalization programs should be requested as well.
To understand a mass dynamical generation at the quark-gluon level is the first necessary step to understand the existence of the physical mass spectrum at the hadronic level in QCD as a theory of strong interactions. The confinement first and only after the PCAC phase transitions in QCD have to be understood 2 ; 23 . QCD is a self-consistent quantum gauge field theory. It does not need any extra degrees of freedom (such as Higgs fields 40 ) in order to generate a mass at the fundamental quark-gluon level.
Concluding, a few remarks are in order. The wide-spread opinion that the gauge symmetry cannot be broken/violated is misleading. The Elitzur’s theorem 41 has been proven in lattice field theory. It is supposed to be valid for the continuous theories as well, but if yes than only for the abelian systems, like QED. To the non-abelian systems, such as the YM fields, this theorem cannot be applied. In this paper we have just proven that the gauge symmetry of the YM theory is broken in its vacuum. The relevant existing knowledge on a mass dynamical generation and spontaneously symmetry breaking subjects has been taken into account in the list of publications, presented in our paper (and see references in each of the cited items there).
Acknowledgments
The authors are grateful to P. Levai, P. Forgács, J. Nyiri, T.S. Biró, Gy. Kluge, M. Vasúth for useful suggestions, remarks, discussions and help. The work was supported by the Hungarian National Research, Development and Innovation Office (NKFIH) under the contract numbers OTKA K135515 and NKFIH 2019-2.1.11-TÉT-2019-00078, 2019-2.1.11-TÉT-2019-00050, the Wigner Scientific Computing Laboratory.
Appendix A The full gluon propagator up to one loop
Figure 2: The SD equation for the full gluon propagator up to one loop as present in 6 . It is useful to compare it with Fig. 1 and its description in the text after it.
In order to explicitly compare the tadpole term (2.4) with the other tadpole-type terms let us consider the iteration of the gluon SD eq. (2.1), shown in Fig. 1, up to one loop, shown in Fig. A.2. The last two terms in this figure appear because of the NL nature of the iteration series, when the next iterations contribute into the previous ones. That is the reason why such multiplied tadpole-type terms are not present in the skeleton loop decomposition for the full gluon propagator in Fig. 1, but have to be shown up in the iteration terms, seen in Fig. A.2. This iteration has been exactly calculated in 6 by using the DRM, while rightfully now ignoring the tadpole term itself and all the other tadpole-likes ones, in accordance with the relations (5.5)-(5.6). It is worth emphasizing that in 6 this iteration has not been called the PT expansion up to $O(g^{2})$, though looks like it, but expansion up to $O(\hbar)$. It is well-known that in QCD the expansion in powers of the coupling constant does not make any sense because it is strong. The PT expansion in QCD makes sense only in the AF 37 ; 38 regime when it becomes weak. In order to perform analytical derivations below, we will take the interaction vertices as the point-like ones. The external gluon propagators will be taken as the free ones. All the skeleton loop propagators will remain as the full ones. Then the corresponding iteration will look like Fig. A.2, and hence can be called as quasi-one loop expansion, discussed in some details in section 8, when it was called as the cluster one. Note that below we do not show explicitly the dependence on the coupling constant, since our skeleton loop quark, gluon and ghost propagators are the full ones, so that the exact dependence on the coupling constant of the expressions below is not known. Also we will omit below all the colour group factors and other overall numerical numbers, since all of them are unimportant here.
However, it is instructive first to show explicitly how works our subtraction scheme, using only the third diagram in the second line of Fig. A.2, for simplicity. From 6 by replacement $\mu\rightarrow\rho$ and $\nu\rightarrow\sigma$ and fixing $\omega=2$ and $\delta_{\rho\rho}=4$ in eq. (6.22) there, one finally obtains an analytical expression for this term as follows:
$\Pi^{(3)}_{\rho\sigma}(q)\sim\int d^{4}l\ {N_{\rho\sigma}(l,q)\over l^{2}(l+q)^{2}},$ (A.1)
where, apart from below, we put the gluon loop propagators as free ones in the t’ Hooft-Feynman gauge as well, for the sake of transparency of our derivations for this diagram. The tensor function $N_{\rho\sigma}(l,q)$ is
$N_{\rho\sigma}(l,q)=10l_{\rho}l_{\sigma}+5(l_{\rho}q_{\sigma}+q_{\rho}l_{\sigma})-2q_{\rho}q_{\sigma}+(5q^{2}+2(lq)+2l^{2})\delta_{\rho\sigma}.$ (A.2)
Putting in eqs. (A.1) and (A.2) $q=0$, one gets
$\Pi^{(3)}_{\rho\sigma}(0)\sim\int d^{4}l\ {N_{\rho\sigma}(l,0)\over(l^{2})^{2}}=\int d^{4}l\ {(10l_{\rho}l_{\sigma}+2l^{2}\delta_{\rho\sigma})\over(l^{2})^{2}},$ (A.3)
and, obviously, this integral is quadratically divergent at the upper limit, but is assumed to be regularized from above and below from the very beginning, as emphasized throughout this work.
Let us now compose the subtracted counterpart of the initial integral (A.1), in accordance with the subtractions (4.1), as follows:
$\Pi^{(3s)}_{\rho\sigma}(q)=\Pi^{(3)}_{\rho\sigma}(q)-\Pi^{(3)}_{\rho\sigma}(0)\sim\int d^{4}l\ {N^{s}_{\rho\sigma}(l,q)\over(l^{2})^{2}(l+q)^{2}},$ (A.4)
where
$N^{s}_{\rho\sigma}(l,q)$ $=$ $5l^{2}(l_{\rho}q_{\sigma}+q_{\rho}l_{\sigma})-20(ql)l_{\rho}l_{\sigma}+l^{2}(q^{2}-2(lq))\delta_{\rho\sigma}$ (A.5)
-- $2l^{2}q_{\rho}q_{\sigma}-10q^{2}l_{\rho}l_{\sigma}.$
Because of the two last terms and term $l^{2}q^{2}\delta_{\rho\sigma}$ in this equation, the subtracted integral (A.4) is only logarithmically divergent at the upper limit. So that the integral (A.3) is one of the tadpole-like terms which should be disregarded on the general basis due to the relations (6.6)-(6.7) within our approach. The subtracted integral (A.4) finally will contribute to the invariant function which appear in the full gluon propagator (7.5).
The tadpole term up to one skeleton loop, shown first in the third line of Fig. A.2, analytically can be written down as follows:
$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lT^{0}_{\rho\sigma\alpha\beta}D_{\alpha\beta}(l)=\delta_{\rho\sigma}\Delta^{2}_{t}(D).$ (A.6)
Being the QD but already regularized constant, it is not connected to any external gluon momentum. In this respect it is different from all other tadpole-like terms which appear as a result of the subtraction scheme or in the NL iteration procedure. Also it directly generate the mass squared parameter, and thus explicitly violets the gauge symmetry of the QCD Lagrangian in its ground state 6 . We have shown that in the PT QCD it should be neglected on the general basis as well as all other tadpole-type terms. But in the QCD it should remain intact, while all other tadpole-types terms have to be removed from the theory, as it has been proven in section 6.
Let us now investigate the tadpole-like term with the gluon skeleton loop, shown second in the third line of Fig. A.2. Its analytical expression is
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l\ T^{0}_{\nu\zeta\mu}(l,p)D_{\nu\zeta}(l),$ (A.7)
where the gluon momentum $p$ is, in fact, zero, i.e., $p=0$ but it is convenient to go to this limit at the final step only. Both gluon propagators in the t’ Hooft-Feynman gauge are $D^{0}_{\mu\mu^{\prime}}(p)=\delta_{\mu\mu^{\prime}}/p^{2}$ and $D_{\nu\zeta}(l)=\delta_{\nu\zeta}d(l^{2})/l^{2}$, where $d(l^{2})$ is the corresponding invariant function of the full gluon propagator. Let us remind that such kind of the skeleton loop integrals are assumed to be regularized from above and below. Using now the Euclidean space Feynman rules for the corresponding vertices present in 6 , one arrives at
$T^{0}_{\rho\sigma\mu^{\prime}}(q,p)$ $=$ $-(2q+p)_{\mu^{\prime}}\delta_{\rho\sigma}+(q-p)_{\sigma}\delta_{\rho\mu^{\prime}}+(2p+q)_{\rho}\delta_{\sigma\mu^{\prime}},$
$T^{0}_{\nu\zeta\mu}(l,p)$ $=$ $2(l+p)_{\mu}\delta_{\nu\zeta}-l_{\zeta}\delta_{\nu\mu}-(2p+l)_{\nu}\delta_{\zeta\mu}.$ (A.8)
Substituting these expressions into the eq. (A.7), one finally obtains
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p){2\over p^{2}}\int d^{4}l{d(l^{2})\over l^{2}}[3l_{\mu}+p_{\mu}].$ (A.9)
By the symmetry integration ($d^{4}l=l^{3}dl=(1/2)l^{2}dl^{2}$ and all the overall numerical numbers due to the integration over angular variables in four-dimensional Euclidean space 6 omitted) the first loop integral is zero, while the second one leads to
$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p)p_{\mu}{2\over p^{2}}\times\tilde{\Delta}^{2}_{g}(d),$ (A.10)
where $\tilde{\Delta}^{2}_{g}(d)=\int d^{4}l(d(l^{2})/l^{2})$ is the QD but regularized skeleton loop integral, having the dimension of a mass squared. It is one of the constants present in the relations (6.7)-(6.8). It should be discarded on the general basis within our approach, i.e., put formally zero, so that $\tilde{\Delta}^{2}_{g}(d)=0$, even before the final $p^{2}=-M^{2}\rightarrow 0$ limit. Let us note that the cancellation of the pole $1/p^{2}=-1/M^{2}$ in eq. (B.10) can be always achieved by going to the dimensionless variable $x=l/M$ in the skeleton loop integral (A.9), i.e., this pole is not a problem here. Thus, $\Pi^{(4)}_{\rho\sigma}(q,p)=0$.
Let us finally investigate the tadpole-like term with the ghost skeleton loop, shown third in the third line of Fig. A.2. In analogy with eq. (A.7) its analytical expression is
$\Pi^{(5)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l{l_{\mu}\over l^{2}}d_{gh}(l^{2}),$ (A.11)
where $d_{gh}(l^{2})$ is the invariant function for the full ghost propagator $G(l)$. The vertex $T^{0}_{\rho\sigma\mu^{\prime}}(q,p)$ is explicitly shown in the relations (A.8), and the remaining free gluon propagator is also the same. The altered rule 6 for the gluon-ghost vertex gives $(1/2)[(l_{\mu}+l{\mu})=l_{\mu}$ and the gluon momentum $p$ is zero again at the final stage. Since the nominator of this skeleton loop integral contains the loop variable linearly, then by the symmetry integration it is simply zero from the very beginning, i.e., $\Pi^{(5)}_{\rho\sigma}(q,p)=0$.
Concluding, neither the tadpole-like term with the gluon skeleton loop nor its ghost counterpart contribute into the full gluon propagator. In these cases the subtractions with respect to the external gluon momentum $q$ (though possible), but were not even necessary to perform, since the final results would have been the same zeros. The primary goal of this work is to prove that the tadpole term cannot be disregarded by any means. All the other tadpole-like terms are to be always discarded on the general basis, i.e., put zeros, as explicitly demonstrated here as well. This was not a prescription, but being an exact mathematical result within our formalism.

### v3

@@
-The gauge symmetry - unbroken and broken - in the QCD ground state
+The gauge symmetry violation in the QCD ground state
 Abstract
-The main purpose of this work is to report on convincing progress made in the investigation of the one of the long-standing conceptual problems in QCD, namely whether the gauge symmetries of the QCD Lagrangian and its ground state coincide or not?
+The purpose of this work is to report on convincing progress made in the investigation of the true dynamical and gauge structures of the QCD ground state (vacuum).
+It also sheds new significant insights into a mass squared dynamical generation in it.
 For this purpose, we use only the tensor algebra derivation rigorous rules in dealing with the corresponding skeleton loop integrals, contributing into the full gluon self-energy.
-It has been exactly (i.e., without making any kind of the simplifications) proven that in the QCD ground state may exist the two principally different types of the gluon field configurations.
-One of them maintains the QCD Lagrangian’s gauge symmetry, but another one does not.The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared.
+The novel constraint on any solution to QCD has been exactly and uniquely derived in the framework of the the Slavnov-Taylor identities for the gauge particles propagators.
+Its general solution makes it possible to prove that the gauge symmetry of the QCD Lagrangian is broken in its ground state.
+The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared.
 The exact proof has been given that it cannot be disregarded from the theory and its ground state by any means.
-The gauge symmetry of the QCD Lagrangian forbids such term, but the perturbative renormalizability of the theory will not be affected.
-This allows to present the new significant insights into the gauge and dynamical structures of the true QCD ground state, and sheds light on a mass squared dynamical generation in it.
+The gauge symmetry of the QCD Lagrangian forbids such kind of terms, but the perturbative renormalizability of the theory will not be affected.
+The particular solution of the novel constraint describes the gluon field configurations with unbroken gauge symmetry in the QCD vacuum.
+The solution with broken gauge symmetry dominates over the solution with unbroken one at large distances, while being in agreement with it at short distances.
 keywords:
-theoretical particle physics, Quantum Chromodynamics, ground state, gluon full self-energy, dynamical and gauge structures.
+Theoretical particle physics, Quantum Chromodynamics, ground state, full gluon self-energy, gauge
 PACS:
-11.10.-z, 11.15.-q, 12.38.-t, 12.38.Aw, 12.38.Lg
-MSC:
-[2020] 30D30, 81Q40, 81T13, 81T16
+11.10.-z , 11.15.-q , 12.38.-t , 12.38.Lg , 12.38.Aw
+[…]
+[table]capposition=top
 1 Introduction
 The quark model (QM) treats the strongly-interacting hadrons (baryons and mesons) as bound-states of quarks, by emitting and absorbing gluons.
-The theory, which purpose is to describe the properties of the observed hadrons in terms of the non-observable quarks and gluons from first principles, is Quantum Chromodynamics (QCD) 1 ; 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 .
+The theory, which purpose is to describe the properties of the observed hadrons in terms of the non-observable quarks and gluons from first principles, is Quantum Chromodynamics (QCD) [1, 2, 3, 4, 5, 6, 7, 8, 9].
 It is widely accepted as the quantum gauge field theory of strong interactions.
 Being such a theory, it has to undergo the two phase transitions: the first one at the fundamental (microscopic) quark-gluon level is – the confinement phase transition – in order to explain why all the physical states are colour-singlets.
-The second one at the (macroscopic) hadronic level is – the PCAC (partially conserved axial currents) phase transition – in order to explain the soft pion physics 2 ; 23 .
+The second one at the (macroscopic) hadronic level is – the PCAC (Partially Conserved Axial Currents) phase transition – in order to explain the soft pion physics [2, 23].
 However, this purpose remains a formidable task yet because of the multiple dynamical and topological complexities of low-energy particles (hadrons) physics, originated from QCD and its ground state (vacuum).
@@
 A. Whether the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is a symmetry of its ground state or not?
-B. The dynamical generation of a mass squared in the vacuum of QCD, since the gauge invariance of its Lagrangian forbids such kind of terms, like the gluon mass term $M^{2}_{g}A_{\mu}A_{\mu}$.
+B.
+The dynamical generation of a mass squared in the vacuum of QCD, since the gauge invariance of its Lagrangian forbids such kind of terms, for example like the gluon mass term $M^{2}_{g}A_{\mu}A_{\mu}$.
 The solutions of these fundamental problems is a key to the explanation of the first confinement phase transition in QCD.
-The properties and symmetries of the QCD Lagrangian, and thus including its Yang – Mills (YM) part, are well-known 1 ; 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 ; 10 .
+The properties and symmetries of the QCD Lagrangian, and thus including its Yang – Mills (YM) part, are well-known [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
 The propagation of gluons is one of the main dynamical effects in the QCD ground state.
 The importance of the corresponding equation of motion is due to the fact that its investigation is supposed to reflect the dynamical and gauge structures of the QCD ground state.
-The gluon Schwinger – Dyson (SD) equation 2 ; 10 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 is highly non-linear (NL) because of the self-interaction of massless gluon modes.
+The gluon Schwinger – Dyson (SD) equation [2, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] is highly non-linear (NL) because of the self-interaction of massless gluon modes.
 They are taken into account by the full gluon self-energy, which is the main part of the gluon SD equation itself.
 We will show that the general properties of the full gluon self-energy point out on some new dynamical and gauge aspects of the true QCD ground state.
-Due to these inputs, the novel insights into a mass dynamical generation in the QCD vacuum are also present within proposed below formalism, based on the tensor algebra derivation rigorous rules only.
-2 The gluon SD equation
-The structure of the gluon SD equation is given in this section in some necessary details. Analytically it looks like
+Due to these impacts, the novel insights into a mass dynamical generation in the QCD vacuum are also present within proposed below formalism.
+It is based on the tensor algebra derivation rigorous rules only, which is widely using in the theoretical and mathematical physics.
+2 Gluon SD equation
+The system of the SD equations of motions in QCD, including the above-mentioned gluon SD equation, in general, contains much more informations on the properties of the theory than its Lagrangian can provide at all [2].
+So it might be the place where a mas squared scale parameter, forbidden by the QCD Lagrangian gauge symmetry, could be generated.
+Among them, the special role belongs to the equations, describing the propagation of the fundamental constituents in the QCD vacuum – gluons and quarks.
+The gluon SD equation has a rather complicated tensor structure because of its above-mentioned NL character.
+Let us emphasize that all the known interactions in nature (electroweak, gravitational and strong) are described by the gauge theories.
+The common mathematical language for them is the tensor algebra, which is only one used throughout this paper.
+That is why our works will be easily understood by a broad spectrum of researches from the different areas of the theoretical physics.
+The gluon SD equation analytically looks like
 $D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{0}_{\mu\rho}(q)i\Pi_{\rho\sigma}(q;D)D_{\sigma\nu}(q),$ (2.1)
@@
 1.
-Figure 1: The SD equation for the full gluon propagator as present in 10 .
+Figure 1: The SDE for the full gluon propagator, as present in [10].
 Here stringy lines are for the free gluon propagator, while $D$ denotes its full counterpart.
@@
 $\Pi_{\rho\sigma}(q;D)=\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)+\Pi_{\rho\sigma}^{t}(D),$ (2.2)
-where $\Pi^{q}_{\rho\sigma}(q)$ describes the skeleton loop contribution of the quark degrees of freedom as an analogue to the vacuum polarization tensor in Quantum Electrodynamics (QED) 7 ; 21 .
-Note that here and below the superscript or subscript ’$q$’ means quark (not to be mixed up with the gluon momentum variable $q$).
+where $\Pi^{q}_{\rho\sigma}(q)$ describes the skeleton loop contribution of the quark degrees of freedom as an analogue to the vacuum polarization tensor in Quantum Electrodynamics (QED) [7, 21].
+Here and below the superscript or subscript ’$q$’ means quark (not to be mixed up with the gluon momentum variable $q$).
 The gluon part of the full gluon self-energy by itself is the sum of a few independent terms as follows:
-$\Pi^{g}_{\rho\sigma}(q;D)=\Pi^{gh}_{\rho\sigma}(q)+\Pi^{(1)}_{\rho\sigma}(q;D^{2})+\Pi^{(2)}_{\rho\sigma}(q;D^{4})+\Pi^{(2^{\prime})}_{\rho\sigma}(q;D^{3}),$ (2.3)
+$\Pi^{g}_{\rho\sigma}(q;D)$ $=$ $\Pi^{gh}_{\rho\sigma}(q)+\Pi^{(1)}_{\rho\sigma}(q;D^{2})$ (2.3)
+$+$ $\Pi^{(2)}_{\rho\sigma}(q;D^{4})+\Pi^{(2^{\prime})}_{\rho\sigma}(q;D^{3}),$
 and $\Pi^{gh}_{\rho\sigma}(q)$ describes the skeleton loop contribution associated with the ghost degrees of freedom.
@@
 The last four terms explicitly contain the full gluon propagators in the corresponding powers symbolically shown above.
-They form the NL part of the gluon SD equation.
-The analytical expressions for the corresponding skeleton loop integrals 22 , in which the symmetry and combinatorial coefficients and signs have been included, are not important here.
+The analytical expressions for the corresponding skeleton loop integrals [22], in which the symmetry and combinatorial coefficients and signs have been included, are not important here.
 We are not going to calculate any of them explicitly, and thus to introduce into them any truncations/approximations/assumptions or choose some special gauge.
@@
 1 is the sum of the infinite number of terms.
-Moreover, the full vertices entering these skeleton loop integrals are themselves determined by an infinite series of the corresponding multi-loop skeleton terms 2 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 .
+Moreover, the full vertices entering these skeleton loop integrals are themselves determined by an infinite series of the corresponding multi-loop skeleton terms [2, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20].
 In these NL series the dependence on the coupling constant may not be simple, i.e., in fact, it is not known.
-Such kind of series are the so-called ’cluster’ expansions 23 , indeed.
-The constant tadpole term $\Pi_{\rho\sigma}^{t}(D)$ is defined as follows:
-$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lD_{\alpha\beta}(l)T^{0}_{\rho\sigma\alpha\beta}=g_{\rho\sigma}\Delta^{2}_{t}(D)=[T_{\rho\sigma}(q)+L_{\rho\sigma}(q)]\Delta^{2}_{t}(D),$ (2.4)
-where $L_{\rho\sigma}(q)=q_{\rho}q_{\sigma}/q^{2}$, so that the tadpole term contributes into the both transverse and longitudinal components of the full gluon propagator (2.1).
+Such kind of series are the so-called ’cluster’ expansions [23], indeed.
+The constant tadpole term $\Pi_{\rho\sigma}^{t}(D)$ is defined as,
+$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lD_{\alpha\beta}(l)T^{0}_{\rho\sigma\alpha\beta}=g_{\rho\sigma}\Delta^{2}_{t}(D),$ (2.4)
+where $g_{\rho\sigma}=T_{\rho\sigma}(q)+L_{\rho\sigma}(q)=T_{\rho\sigma}(q)+q_{\rho}q_{\sigma}/q^{2}$.
 In what follows the constant $\Delta^{2}_{t}(D)$ will be called as the tadpole term itself, since it deviates from the definition (2.4) by the tensor $g_{\rho\sigma}$ only.
 It is instructive for further purpose to present the initial gluon SD eq. (2.1), on account of the expressions (2.2) and (2.4) as follows:
-$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)$ $+$ $D^{0}_{\mu\rho}(q)i\left[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)\right]D_{\sigma\nu}(q)$ (2.5)
-$+$ $D^{0}_{\mu\rho}(q)ig_{\rho\sigma}\Delta^{2}_{t}(D)D_{\sigma\nu}(q).$
+$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{T}_{\mu\rho}(q)+D^{L}_{\mu\rho}(q)$ (2.5)
+with
+$D^{T}_{\mu\nu}(q)$ $=$ $D^{0}_{\mu\rho}(q)i\left[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)\right]D_{\sigma\nu}(q),$
+$D^{L}_{\mu\nu}(q)$ $=$ $D^{0}_{\mu\rho}(q)ig_{\rho\sigma}\Delta^{2}_{t}(D)D_{\sigma\nu}(q).$ (2.6)
 All the terms which contribute to the full gluon self-energy eq.
@@
 (2.5), are tensors, having the dimensions of a mass squared.
-All these skeleton loop integrals are therefore quadratically divergent (QD), i.e., ultraviolet (UV) divergent in the perturbation theory (PT) regime, and so they are assumed to be regularized.
+All these skeleton loop integrals are therefore quadratically divergent (QD), i.e., ultraviolet (UV) divergent in the perturbation theory (PT) regime, and so they are assumed to be regularized from above.
 Contrary to QED, QCD being a non-abelian gauge theory can suffer from the severe infrared (IR) singularities in the $q^{2}\rightarrow 0$ limit (i.e., more singular than the PT one, which is proportional to $\sim 1/q^{2}$).
 Evidently, they will be generated due to the self-interaction of massless gluon modes.
-Thus, all the possible subtractions at zero may be dangerous 2 ; 3 ; 4 ; 5 ; 6 ; 7 .
-That is why in all the quantities below, the dependence on the finite (slightly different from zero) dimensionless subtraction point $\alpha$ has to be understood.
-In other words, all the subtractions at zero and the Taylor expansions around zero should be understood as the subtractions at $\alpha$ and the structure of the Taylor expansions near $\alpha$, where they are justified to be used.
-From the technical point of view, however, it is convenient to put formally $\alpha=0$ in all the expressions and derivations below, and to restore the explicit dependence on non-zero $\alpha$ in all the quantities only at the later stage.
-In all the quantities where the dependence on the dimensionless UV regulating parameter $\tilde{\lambda}$ and $\alpha$ is not shown explicitly, nevertheless, it should be also assumed.
-For example, $\Pi_{\rho\sigma}(q;D)\equiv\Pi_{\rho\sigma}(q;D,\tilde{\lambda},\alpha)$ and similarly for all other quantities.
+Thus, all the skeleton loop integrals, contributing to the full gluon self-energy, are assumed to be regularized from below as well.
 This means that all the expressions are regularized (i.e., we render infinite skeleton loop integrals to be finite).
+How precisely the regulating parameters have to be introduced is not important here, since we are not going to calculate any of these skeleton loop integrals explicitly, as emphasized above.
+They will be assumed but not shown in all regularized expressions, for simplicity.
 The tadpole term $\Delta^{2}_{t}(D)$, defined in eq.
-(2.4), is nothing else but the QD skeleton loop integral, which is already regularized from below and above as well as all the other such kind of constants which may appear in the theory, as a results of the proper subtraction scheme (see below).
+(2.4), is nothing else but the QD skeleton loop integral, which is already regularized from below and above as well as all the other such kind of constants which may appear in the theory by any possible ways (the subtraction schemes, the iteration solutions, etc., see below).
 Being a mass squared regularized quantity, it is explicitly present in the QCD ground state, see eq.
@@
 However, it violets the gauge symmetry of the QCD Lagrangian.
-Then a natural question arises why is it present in the vacuum of QCD at all if it makes theory to look un- renormalizable from the very beginning?
+Then a natural question arises why is it present in the vacuum of QCD at all if it makes the theory to look non-renormalizable from the very beginning?
 The standard solution to this dilemma was to remove this and all other QD constants in order to make theory renormalizable in accordance with the QCD Lagrangian’s exact gauge symmetry.
 However, here we are going to solve this dilemma by asking the question is it possible to retain the tadpole term in the QCD vacuum, but without affecting the PT renormalization properties of the theory?
-This is important to ask because, unlike to other QD constants, the tadpole term is only one which generates a mass squared scale parameter 6 .
+This is important to ask because, unlike to other QD constants, the tadpole term is only one which dynamically generates a mass squared scale parameter, dominated by the point-like four-gluon vertex only, and thus does not depending on the external gluon momentum [6].
 Also, one of our primary aims will be to clarify its role in the dynamical and gauge structures of the QCD vacuum in more detail.
-For this purpose we will develop a suitable formalism, based only on the rigorous rules of the tensor algebra derivations.
 Concluding, let us remind that the whole gluon momentum range is $q^{2}\in[0,\infty)$.
@@
 This makes it possible to avoid the unphysical IR singularities on the light cone.
-3 Slavnov-Taylor identities
-The basic relation to which the full gluon propagator should satisfy is the corresponding Slavnov-Taylor (ST) identity 2 ; 24 ; 25 , namely
+3 Exact constraint on any solution to QCD
+It is well known that all the Slavnov-Taylor (ST) identities which appear in QCD [1] are important for the renormalizability of the theory, and thus they ”are exact constraints on any solution to QCD” [2].
+Among them the ST identities [2, 24, 35] for the gauge particle propagators play a dominant role, since they determine the gauge structure of the QCD ground state (vacuum).
+The ST identity for the full gluon propagator looks like
 $q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi,$ (3.7)
@@
 Throughout this paper we use the standard definition of $T_{\mu\nu}(q)=\delta_{\mu\nu}-q_{\mu}q_{\nu}/q^{2}=\delta_{\mu\nu}-L_{\mu\nu}(q)$ in Euclidean metric.
-Any invariant functions associated with the projective operators $T_{\mu\nu}(q)$ and $\delta_{\mu\nu}$ are the same, and thus $D_{\mu\nu}(q)$ is defined up to its longitudinal part $L_{\mu\nu}(q)$.
+Any invariant functions associated with the projection operators $T_{\mu\nu}(q)$ and $\delta_{\mu\nu}$ are the same, and thus $D_{\mu\nu}(q)$ is defined up to its longitudinal part $L_{\mu\nu}(q)$.
 This should be also valid for its equation of motion.
-If one neglects all the contributions to the full gluon self-energy in eq. (2.1), by setting $d(q^{2})=1$ in eq. (3.2), then one obtains the free gluon propagator
+By formally setting $d(q^{2})=1$ in eq.
+(3.8), then one obtains the free gluon propagator
 $D^{0}_{\mu\nu}(q)=i\left[T_{\mu\nu}(q)+\xi_{0}L_{\mu\nu}(q)\right](1/q^{2}),$ (3.9)
@@
 Let us note in advance that from the very beginning the gauge-fixing parameter for the full gluon propagator $\xi$ is not to be identified with its free counterpart $\xi_{0}$.
-Under which conditions they are the same or different is one of the main subjects of our investigation in this paper.
-Since the gauge freedom in QCD is determined by $\xi_{0}$, the gauge-fixing parameter $\xi$ has to be a function of $\xi_{0}$ i.e., $\xi=f(q^{2};\xi_{0})$ and called the gauge-fixing function.
-Emphasizing, these and all other ST identities which appear in QCD 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 9 ; 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 are important for the general renormalizability of the theory, and thus they ”are exact constraints on any solution to QCD” 2 .
-3.1 Transverse relations for the full gluon self-energy
-Let us now perform some algebraic derivations which are necessary for the further purposes.
+Since the gauge freedom in QCD is determined by $\xi_{0}$, there should exist a relation between them.
+Let us now perform some exact algebraic derivations which are necessary for the further purposes.
 Contracting the full gluon SD eq.
-(2.1) and eq.
 (2.1) with $q_{\mu}$ and $q_{\nu}$, on account of the relations (3.7)-(3.10), one gets
-$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)$ $=$ ${(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2},$
-$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]$ $=$ ${(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}-q^{2}\Delta^{2}_{t}(D),$ (3.5)
-and they related to each other through the relation (2.2) as well as eq.
-(2.1) is related to eq.
-(2.5), which is as it should be.
-The first line in the transverse relations (3.5) is the tensor relation, while its second one is nothing else but the tensor identity.
-The second transverse relation explicitly depends on the tadpole term $\Delta^{2}_{t}(D)$.
-If it is formally put zero, then the both transverse relations (3.5) coincide with each other.
+$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2},$ (3.11)
+where the full gluon self-energy, due to the eqs. (2.2) and (2.4), can be present as follows:
+$\Pi_{\rho\sigma}(q;D)=\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)+\delta_{\rho\sigma}\Delta^{2}_{t}(D),$ (3.12)
+and contracting it with with $q_{\rho}$ and $q_{\sigma}$, and because of the previous relation (3.11), one obtains
+$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}-q^{2}\Delta^{2}_{t}(D).$ (3.13)
+Obviously, this transverse relation can be obtained by the corresponding contraction of eq.
+(2.5).
+It explicitly depends on the tadpole term $\Delta^{2}_{t}(D)$.
+If it is formally put zero, i.e., by hand, then the both transverse relations (3.11) and (3.13) coincide with each other.
 This underlines its important role in the QCD ground state, which one precisely is to be clarified in our paper.
+The transverse relation (3.11) and the transverse relation obtained from eq.
+(3.12) are independent from each other, since the first one has been derived by using the gluon SD eq.
+(2.2), while the second one by using the initial definition (2.6).
+From all these transverse relations one concludes that by themselves they cannot remove the QD constants which may appear in the theory (in order to make it in the PT sense renormalizable), and to fix the relation between $\xi$ and $\xi_{0}$.
+Let us now remind that in QCD the quark skeleton loop contribution into the full gluon self-energy can be always made transverse because of the electric charge conservation which flows around the closed loop (in complete analogy with QED, see [10] and derivations below).
+Therefore, $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}=0$ and this relation holds independently from the gluon contribution, which has been defined in eq.
+(2.3).
+At the same time, it is well-known that in QCD that just the Faddeev – Popov (FP) ghost contribution [43], $\Pi^{gh}_{\rho\sigma}(q)$ makes the transverse relation for this term satisfied (independently from the tadpole term (2.4), which itself has the transverse projection operator structure).
+It is instructive to emphasize that each term in the sum (2.3) cannot be made transverse, only the whole sum.
+How explicitly this works can be found in any text book on QCD.
+For the most transparent calculations up to one loop contribution to the transverse gluon self-energy (2.3) see for example [3, 4, 5, 6].
+Of course, such a transversality will be achieved in the higher loop iteration terms as well, i.,e, it is a general fact, as it is requested by the satisfied transverse relation for the (2.3) term.
+It is a sum of the skeleton loop contributions, as underlined from the very beginning in this work.
+This means that $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ and this relation holds independently from the quark contribution.
+So that, in general, in QCD one has
+$q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0,$ (3.14)
+and then from the relation (3.13), one immediately arrives at the exact condition, namely
+${(\xi_{0}-\xi)\over\xi\xi_{0}}q^{2}=\Delta^{2}_{t}(D),$ (3.15)
+which can be treated as the exact constraint on any solution to QCD, since coming out from the corresponding ST identities, and deriving before going to the formulation of any proper subtraction scheme.This underlines its intrinsic, independent and important status in the theory.
+3.1 General solution
+The general solution of the exact constraint (3.15) determines the function $\xi=f(q^{2};\xi_{0})$ as follows:
+$\xi=f(q^{2};\xi_{0})={\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)}.$ (3.16)
+We will call it as the generalized gauge-fixing function approach to QCD as well.
+It has been exactly (i.e., without making any kind of the simplifications) and uniquely defined in a such new manner for the first time.This expression demonstrates the NL dependence $\xi=f(q^{2};\xi_{0})$ on $\xi_{0}$.
+The linear relation $\xi\rightarrow\xi_{0}$ will be recovered in the PT $q^{2}\rightarrow\infty$ limit, when the ratio $(\Delta^{2}_{t}(D)/q^{2})$ is to be suppressed in this regime at finite $\xi_{0}$.
+We have established the general boundary condition in QCD for the behaviour of the gauge-fixing function in the PT limit for the regularized full gluon propagator.
+Since the general solution (3.16) has been obtained within the ST identities formalism, we will call it as the general solution to QCD.
+Let us underline that to know it means to know the true gauge structure of the QCD vacuum and its theory at the fundamental quark-gluon level (for more detailed description see section 5 as well).
+3.2 Particular solution
+The exact constraint (3.15) has also a particular solution, namely
 […]
-However, by this mechanism the QD constants due to the quark degrees of freedom only can be excluded from the QCD theory.
-What is about the exclusion of the QD constants from its YM sector?
-QCD being much more complicated quantum field gauge theory than QED, has the two transverse relations (3.5), second of which shows the gluon degrees of freedom explicitly.
-By solution of these relations, we mean that at least one of them has to be necessarily satisfied.
-Finally, this will lead to the removal of the corresponding QD constants from the YM theory as well as to fix the function $\xi=f(q^{2};\xi_{0})$, see exact derivations in sections 4, 5 and 6.
-This strategy is definitely accessible with conventional means of the gauge theories to make them renormalizable, and QCD is a such theory 2 , indeed.
-So that whether these transverse relations will be satisfied/saturated (i.e., equal zero) or not and further include them into the proper subtraction scheme requires much more careful investigation in QCD.
-Obviously, there exist only the two different requirements in order to satisfy at least one of the transverse relations (3.5).
-However, it is time now to introduce the proper subtraction scheme.
-4 The proper subtraction scheme
+but these equalities should be put by hand in this case, i.e., they are the prescriptions rather than the exact mathematical results.
+We will call it as the particular solution to QCD.
+Let us note in advance that these conditions determine the so-called conventional or, equivalently, conventional PT QCD (though the coupling constant remains always strong apart from the asymptotic free (AF) regime [2, 38, 39]).
+Precisely, this solution to QCD has been studied the last five decades.
+In this solution the gauge symmetries of the QCD Lagrangian and its ground state coincide, like in QED (see section 6 as well).
+4 Proper subtraction scheme
+QCD being much more complicated quantum field gauge theory than QED, has the two satisfied transverse relations (3.14), and so requires much more careful investigation in QCD.
+They have to be included in a self-consistent way into the proper subtraction scheme.
+Finally this will lead to the removal of the corresponding QD constants from the theory.
+How they fixed the gauge-fixing function approach to QCD in the most general and unique ways has been just described here.
 The first step in the renormalization program of any gauge theory is the removal of the quadratic UV divergences in order to make the corresponding theory renormalizable in the PT sense.
 It can be achieved by introducing the proper subtraction scheme in order to separate them from the PT logarithmic divergences.
-The preliminary step in the regularization program has been already done by introducing the corresponding regulating parameters $\tilde{\lambda}$, $\alpha$.
+The preliminary step in the regularization program has been already done by introducing the corresponding regulating parameters, mentioned above.
 Within our approach nothing will depend on how exactly the regulating parameters have been introduced.
-Even the hard cut-off procedures can be used, since we are mainly interested in the regulating of leading order divergences 2 .
+Even the hard cut-off procedures can be used, since we are mainly interested in the regulating of leading order UV divergences [2].
 They have to disappear from the theory after the PT and the non-perturbative (NP) renormalization programs will be performed.
 Let us now introduce the subtractions for the quark and gluon contributions to the full gluon self-energy as follows:
-$\Pi^{s(q)}_{\rho\sigma}(q)$ $=$ $\Pi^{q}_{\rho\sigma}(q)-\Pi^{q}_{\rho\sigma}(0)=\Pi^{q}_{\rho\sigma}(q)-\delta_{\rho\sigma}\Delta^{2}_{q},$
-$\Pi^{s(g)}_{\rho\sigma}(q;D)$ $=$ $\Pi^{g}_{\rho\sigma}(q;D)-\Pi^{g}_{\rho\sigma}(0;D)=\Pi^{g}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{g}(D),$ (4.1)
+$\Pi^{s(q)}_{\rho\sigma}(q)=\Pi^{q}_{\rho\sigma}(q)-\Pi^{q}_{\rho\sigma}(0)=\Pi^{q}_{\rho\sigma}(q)-\delta_{\rho\sigma}\Delta^{2}_{q},$ (4.18)
+$\Pi^{s(g)}_{\rho\sigma}(q;D)$ $=$ $\Pi^{g}_{\rho\sigma}(q;D)-\Pi^{g}_{\rho\sigma}(0;D)$ (4.19)
+$=$ $\Pi^{g}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{g}(D),$
 and thus $\Pi^{s(q)}_{\rho\sigma}(0)=\Pi^{s(g)}_{\rho\sigma}(0;D)=0$, by definitions.
@@
 (4.18)-(4.19), as well as the tadpole term $\Delta^{2}_{t}(D)$ itself, are independent from each other and are defined by the skeleton loop integrals, which have been already regularized from above and below.
-The subtraction at zero is to be understood in a such way that we subtract at the external gluon momentum $q^{2}=-\mu^{2}$ 2 with $\mu^{2}\rightarrow 0$ final limit (its dimensionless counterpart is $\alpha$ mentioned above).
-However, due to the self-interaction of the massless gluon modes, some of these constants (apart from the tadpole term) may depend on the internal zero gluon momentum (if it is connected to any closed loop, for example see appendix A).
+The subtraction at zero is to be understood in a such way that we subtract at the external gluon momentum $q^{2}=-\mu^{2}$ [2] with $\mu^{2}\rightarrow 0$ final limit.
+However, due to the self-interaction of the massless gluon modes, some of these constants (apart from the tadpole term) may depend on the internal zero gluon momentum (if it is connected to any closed loop, for example such as which can appear in the iteration solution of the gluon SD equation [6]).
 Then it has to be replaced by the subtraction at $q^{2}_{i}=-M^{2}_{i}$ with $M^{2}_{i}\rightarrow 0$ final limit as well, and subindex $i$ determines the number of such internal gluons.
-For the any full gluon propagator the subindex $i$ can be treated as the number of the necessary subtractions made in it in accordance with the rules of the theory of distributions (generalized functions) 26 .
-Also it is necessary to emphasize that the subtractions (4.1) have not been introduced by hand.
+For the any single full gluon propagator the subindex $i$ can be treated as the number of the necessary subtractions made in it in accordance with the rules of the theory of distributions (generalized functions) [26].
+Also it is necessary to emphasize that the subtractions (4.17)-(4.18) have not been introduced by hand.
 They are equivalent to add zero to the corresponding identity.
@@
 The first term is a regular function of the external gluon momentum $q$, by definition, and it may be only logarithmically divergent at large $q$.
+In other words, we separate the NP QD constants from the PT logarithmically divergent terms in a gauge-invariant way, that is why our subtraction scheme is called as a proper one.
 The same is true for the initial gluon contribution as well.
@@
 For the interacting full gluon propagators these independent terms will interact with each other, but such a separation will simplify the corresponding renormalization programs, anyway.
-The independent tensor decompositions of the quark and gluon degrees of freedom, which appear in the subtraction relations (4.1), are
+The independent tensor decompositions of the quark and gluon degrees of freedom, which appear in the subtraction relations (4.18)-(4.19), are
 $\Pi^{s(q)}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(q)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(q)}_{l}(q^{2}),$
@@
 Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary.
-However, all the invariant functions $\Pi^{s(q)}_{t}(q^{2})$, $\Pi^{s(q)}_{l}(q^{2})$ as well as $\Pi^{s(g)}_{t}(q^{2})$, $\Pi^{s(g)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the subtractions (4.1).
-Substituting all these decompositions (4.3)-(4.4) into the subtractions (4.1), and doing some tensor algebra derivations, one finally obtains
-$\Pi^{s(q)}_{t}(q^{2})=\Pi^{q}_{t}(q^{2})-{\Delta^{2}_{q}\over q^{2}},\quad\Pi^{s(q)}_{l}(q^{2})=\Pi^{q}_{l}(q^{2})+{\Delta^{2}_{q}\over q^{2}},$ (4.5)
+However, all the invariant functions $\Pi^{s(q)}_{t}(q^{2})$, $\Pi^{s(q)}_{l}(q^{2})$ as well as $\Pi^{s(g)}_{t}(q^{2})$, $\Pi^{s(g)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the initial subtractions (4.18)-(4.19).
+Substituting all these decompositions (4.21)-(4.22) into the subtractions (4.18)-(4.19), and doing some tensor algebra derivations, one finally obtains
+$\Pi^{s(q)}_{t}(q^{2})$ $=$ $\Pi^{q}_{t}(q^{2})-{\Delta^{2}_{q}\over q^{2}},$
+$\Pi^{s(q)}_{l}(q^{2})$ $=$ $\Pi^{q}_{l}(q^{2})+{\Delta^{2}_{q}\over q^{2}},$ (4.23)
 and
-$\Pi^{s(g)}_{t}(q^{2})=\Pi^{g}_{t}(q^{2})-{\Delta^{2}_{g}(D)\over q^{2}},\quad\Pi^{s(g)}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})+{\Delta^{2}_{g}(D)\over q^{2}}.$ (4.6)
-Using these relations it is easy to show that for the regularized (i.e., finite) quark and gluon contributions to the initial gluon self-energy their values at zero always are $\Pi^{q}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{q}$ and $\Pi^{g}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{g}(D)$, indeed, i.e, they are general ones, and have been already used in the subtraction relations (4.1) above, as it has to be.
+$\Pi^{s(g)}_{t}(q^{2})$ $=$ $\Pi^{g}_{t}(q^{2})-{\Delta^{2}_{g}(D)\over q^{2}},$
+$\Pi^{s(g)}_{l}(q^{2})$ $=$ $\Pi^{g}_{l}(q^{2})+{\Delta^{2}_{g}(D)\over q^{2}}.$ (4.24)
+Using these relations it is easy to show that for the regularized (i.e., finite) quark and gluon contributions to the initial gluon self-energy their values at zero always are $\Pi^{q}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{q}$ and $\Pi^{g}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{g}(D)$, indeed, i.e, they are general ones, and have been already used in the subtraction relations (4.18)-(4.19) above, as it has to be.
 Also these relations will be true if expressed in terms of the above-mentioned subtraction point $q^{2}=-\mu^{2}$.
-Let us now make an important observation in advance.
-If the initial contributions of the quarks and gluons degrees of freedom to the full gluon self-energy, which are present by the second lines in the relations (4.3) and (4.4), respectively, are not transverse, then one cannot remove the corresponding QD constants $\Delta^{2}_{q}$ and $\Delta^{2}_{g}(D)$ from the theory.
-On the contrary, if their are transverse, i.e., $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$, then from the relations (4.5) and (4.6) it follows that one has to put the corresponding QD constants $\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0$, and thus removing them from the theory, since the corresponding subtracted invariant functions cannot have the pole-type singularities, by definition.
-This observation explicitly shows the necessity to satisfy one of the transverse relations (3.5) from the general point of view.
-An interesting observation follows from the relations (4.5) and (4.6) as well, namely $\Pi^{q}_{t}(q^{2})+\Pi^{q}_{l}(q^{2})=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(q)}_{l}(q^{2})$ and $\Pi^{g}_{t}(q^{2})+\Pi^{g}_{l}(q^{2})=\Pi^{s(g)}_{t}(q^{2})+\Pi^{s(g)}_{l}(q^{2})$.
+An interesting observation follows from the relations (4.23) and (4.24), namely $\Pi^{q}_{t}(q^{2})+\Pi^{q}_{l}(q^{2})=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(q)}_{l}(q^{2})$ and $\Pi^{g}_{t}(q^{2})+\Pi^{g}_{l}(q^{2})=\Pi^{s(g)}_{t}(q^{2})+\Pi^{s(g)}_{l}(q^{2})$.
 So that the corresponding sums do not depend on the corresponding QD but regularized constants, which once more underlines the general character of our subtraction scheme.
-Also, the sum of the subtracted invariant functions is never zero, though the subtracted contributions to the full gluon self-energy are zero at $q^{2}=0$, by definitions, see relations (4.1).
-5 Gauge symmetry is unbroken
-In order to satisfy the first transverse relation (3.5) one needs to put $\xi=f(\xi_{0})=\xi_{0}$, and thus it becomes transverse $q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)=0$, indeed.
-Then from the second of the relations (3.5) it follows that
-$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]=-q^{2}\Delta^{2}_{t}(D),$ (5.1)
-and it is not transverse, as expected. Substituting further the subtractions (4.1) into the relation (5.1), and doing some algebra, one obtains
-$q_{\rho}q_{\sigma}[\Pi^{s(q)}_{\rho\sigma}(q)+\Pi^{s(g)}_{\rho\sigma}(q;D)]=-q^{2}\Delta^{2}(D),$ (5.2)
+Also, the sum of the subtracted invariant functions is never zero, though the subtracted contributions to the full gluon self-energy are zero at $q^{2}=0$, by definitions, see relations (4.18)-(4.19).
+As we already know, the quark and gluon contributions to the full gluon self-energy are transverse $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$, see the relations (3.14).
+Then from the second of the relations (4.21) and (4.22) it follows that $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$ as well.
+So that the second relations in (4.23) and (4.24) will be reduced to
+$\Pi^{s(q)}_{l}(q^{2})={\Delta^{2}_{q}\over q^{2}},\quad\Pi^{s(g)}_{l}(q^{2};D)={\Delta^{2}_{g}(D)\over q^{2}},$ (4.25)
+where we restored the dependence on $D$ in the second relation.
+However, these relations are impossible, since the corresponding invariant functions cannot have the pole-type singularities, by definitions, as explained above.
+One has to put these constants to zero on this general mathematical basis, i.e., $\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0$, and thus $\Pi^{s(q)}_{l}(q^{2})=\Pi^{s(g)}_{l}(q^{2};D)=0$ as well.
+All our results for the QD constants can be now summarized as follows:
+$\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0,\quad\Delta^{2}_{t}(D)\neq 0,$ (4.26)
+and hence $\Delta^{2}_{gh}=\Delta^{2}_{1}(D^{2})=\Delta^{2}_{2}(D^{4})=\Delta^{2}_{2^{\prime}}(D^{3})=0$, as well because of the relation (4.20).
+The system of the relations (4.26) is not only a general but it is a unique one as well.
+They are not prescriptions, since based on a rigorous tensor algebra derivation rules, i.e., they are exact mathematical results.
+The characteristic feature of this system is that the tadpole term $\Delta_{t}^{2}(D)$ remains intact.
+The mathematical reason is that for the constant tadpole term such kind of the non-trivial relations (4.18)-(4.19) does not exist (constant minus the same constant is always zero).
+The essential dynamical source of its survival in the theory is that only the tadpole term can generate a mass squared scale parameter [6].
+All the other quark and gluon terms, contributing to the full gluon self-energy, cannot do this, and thus none of their subtracted counterparts with their quark and gluon constants.
+All these constants, defined by the corresponding skeleton loop integrals, are QD at the upper limit.
+They have to be removed/disregarded from the theory on the general mathematical basis, i.e., put zero, as described above, apart from the tadpole term (2.4).
+In comparison with it, all the other QD quark and gluon constants will be called the tadpole-like/type terms.
+The general question arises now, namely how to understand these exact equalities to zero in the relations (4.26)?
+These equalities mean that any tadpole-like term which may appear in the theory by any possible way has to be discarded/disregarded in the theory, i.e, put zero, independently from any other tadpole-like terms.
+In this way such sums will be always zero, indeed.
+Such constants may appear even not as the result of the subtractions, but, for example as the result of the NL iteration procedure for the full gluon propagator [6] (see appendix A as well).
+In this case the tadpole-like terms may be even multiplied by some regularized functions.
+Whatever their origins would have been in the full gluon propagator, all of them belong to the infinite manifold of the relations (4.26), and thus should be always removed from the theory.
+Collecting our results obtained above for the quark and gluon contributions to the gluon self-energy, one obtains
+$\Pi^{q}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(q)}_{t}(q^{2}),$
+$\Pi^{g}_{\rho\sigma}(q;D)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(g)}_{t}(q^{2};D),$ (4.27)
+since they become transverse and coincide with their subtracted counterparts.
+Reminding also that the invariant functions present in the previous relations are regular functions at small $q^{2}$, i.e., they have no pole-type singularities, and may be only logarithmically divergent at large $q^{2}$.
+This becomes possible only due to the satisfied transverse relations (3.14) for the quark and gluon degrees of freedom.
+Just this decreases the quadratic UV divergences of the corresponding skeleton loop integrals to a logarithmic ones, as it has been described in detail in this section.
+5 General solution to QCD
+Substituting the sum of the terms (4.27) into the eq.
+(3.12), on account of the relation $\delta_{\rho\sigma}=T_{\rho\sigma}(q)+L_{\rho\sigma}(q)$ and further into the gluon SD eq.
+(2.5), one finally gets
+$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{T}_{\mu\nu}(q;D)+D^{L}_{\mu\nu}(q;D),$ (5.28)
 where
-$\Delta^{2}(D)=\Delta^{2}_{q}+\Delta^{2}_{g}(D)+\Delta^{2}_{t}(D),$ (5.3)
-is the sum of all the QD but regularized constants, which may appear in the theory.The QD but regularized constant $\Delta^{2}_{g}(D)$ itself is defined in eq. (4.2).
-Substituting further the first tensor decompositions of the relations (4.3) and (4.4) into the transverse relation (5.2), and doing some algebra, one obtains for their longitudinal invariant functions the following relation
-$\Pi^{s(q)}_{l}(q^{2})+\Pi^{s(g)}_{l}(q^{2})={\Delta^{2}(D)\over q^{2}},$ (5.4)
-which is impossible, since these invariant functions cannot have the pole-type singularities, by definition, as underlined just above.
-So that the QD but regularized constant $\Delta^{2}(D)$ should be discarded from the theory, i.e., put formally zero $\Delta^{2}(D)=0$ on this mathematical basis.
-Then $\Pi^{s(q)}_{l}(q^{2})=\Pi^{s(g)}_{l}(q^{2})=0$ as well, since they are independent from each other.
-From now on in this section, we replace $D$ by $D^{PT}$ (for the explanation see the text below).
-The definitions (4.2) and (5.3) require that this system of the relations for the QD but regularized constants now looks like
-$\Delta^{2}_{q}=\Delta^{2}_{g}(D^{PT})=\Delta^{2}_{t}(D^{PT})=0,$ (5.5)
+$D^{T}_{\mu\nu}(q;D)=D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)[q^{2}\Pi(q^{2};D)+\Delta^{2}_{t}(D)]D_{\sigma\nu}(q),$ (5.29)
 and
-$\Delta^{2}_{gh}=\Delta^{2}_{1}(D^{2}_{PT})=\Delta^{2}_{2}(D^{4}_{PT})=\Delta^{2}_{2^{\prime}}(D^{3}_{PT})=0,$ (5.6)
-because all of them are independent from each other.
-The dimensional regularization method (DRM) 4 ; 6 ; 27 ; 28 ; 29 provides a gauge-invariant scheme to correctly calculate the finite parts of the generally QD loop integrals, while omitting their QD parts, i.e., simply to ignore them and to deal further only with the logarithmic divergences of the PT.
-This was a prescription rather than an exact result, as pointed out in 4 .
-But the relations (5.5)-(5.6) are not prescriptions, they are mathematically exact results.
-How to understand such equalities is convenient to postpone until next section.
-Collecting now all our results obtained above, one arrives at the system of the relations as follows:
-$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D^{PT})=0,$ (5.7)
+$D^{L}_{\mu\nu}(q;D)=D^{0}_{\mu\rho}(q)iL_{\rho\sigma}(q)\Delta^{2}_{t}(D)D_{\sigma\nu}(q),$ (5.30)
+respectively.
+$\Pi(q^{2};D)=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(g)}_{t}(q^{2};D)$, and this invariant function is regular at zero and may have only the logarithmic divergences in the PT $q^{2}\rightarrow\infty$ limit.
+The gluon SD eq.
+(5.28) is now equivalent to the initial gluon SD eq.
+(2.5), only being re-written in the different form.
+Combining the general gluon SD eq. (5.28) with the decompositions (1.2) and (1.3), and on account of the relations (5.29) and (5.30), one obtains
+$d(q^{2})={1\over 1+\Pi(q^{2};D)+(\Delta^{2}_{t}(D)/q^{2})}.$ (5.31)
+This relation is the NL transcendental equation for the different invariant functions $d(q^{2}),\ \Pi(q^{2};D)$ and the constant $\Delta^{2}_{t}(D)$, i.e., $d=f(D(d))$.
+Nevertheless, from this expression is clearly seen that in the PT $q^{2}\rightarrow\infty$ regime, the contribution $(\Delta^{2}_{t}(D)/q^{2})$ can be neglected, but the invariant function $\Pi(q^{2};D)$ may still depend on this ratio under the PT logarithms.
+In the NP region of finite and small gluon momenta this term is dominant, and the dependence of $d(q^{2})$ on $\Delta^{2}_{t}(D)$ may be much more complicated due to the transcendental character of eq (5.31).
+Contracting the full gluon SD eq. (5.28), on account of the relation (5.30), with $q_{\mu}$ and $q_{\nu}$, and substituting its result into the general ST identity (3.7), one arrives at
+$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi_{0}\left(1-\xi{\Delta^{2}_{t}(D)\over q^{2}}\right)=i\xi,$ (5.32)
+which solution is
+$\xi\equiv\xi(q^{2};\xi_{0})={\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)},$ (5.33)
+i.e., in this case the gauge-fixing parameter becomes the function $\xi\equiv\xi(q^{2};\xi_{0})$ and thus it is not a constant equal to $\xi_{0}$.
+Let us point out that the expression (5.33) coincides with the general solution (3.16), which is as it should be.
+So that, its asymptotic properties have been already discussed in the text after eq.
+(3.16).
+The generalized gauge-fixing function (5.33) is the known function of its arguments, while the relation (5.31) is the NL transcendental one.
+Behind the general inequality $\xi\neq\xi_{0}$ is the regularized constant $\Delta^{2}_{t}(D)$ as its dynamical source.
+Substituting equations (5.28) and (5.29)-(5.30) into the general decomposition (3.8) for the full gluon propagator, one finally obtains
+$D_{\mu\nu}(q)$ $=$ $iT_{\mu\nu}(q){1\over q^{2}+q^{2}\Pi(q^{2};D)+\Delta^{2}_{t}(D)}$ (5.34)
+$+$ $iL_{\mu\nu}(q){\lambda^{-1}\over q^{2}+\lambda^{-1}\Delta^{2}_{t}(D)},$
+where we have introduced the useful notation $\xi_{0}=\lambda^{-1}$ [4]. The corresponding ST identity now becomes
+$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi(q^{2};\lambda^{-1})=i{\lambda^{-1}q^{2}\over q^{2}+\lambda^{-1}\Delta^{2}_{t}(D)}.$ (5.35)
+This is the generalized gauge-fixing function since it depends on the constant $\Delta^{2}_{t}(D)$, and when it is zero, one recovers the gauge-fixing parameter for the free gluon propagator.
+In this ST identity the gauge-fixing parameter $\lambda^{-1}$ is convenient to vary continuously from zero to infinity.
+The functional dependence of the generalized gauge-fixing function $\xi$ (5.33) is fixed up to an arbitrary gauge-fixing parameter $\xi_{0}=\lambda^{-1}$.
+Unless we fix it, and thus $\xi$ itself, we will call such situation as the general gauge dependence (GGD), see equations (3.7)-(3.8), (5.32)-(5.34).
+Choosing $\xi_{0}=\lambda^{-1}$ explicitly, we will call such situation as the explicit gauge dependence (EGD).
+For example, $\xi_{0}=\lambda^{-1}=0$ is called the unitary (Landau) gauge, $\xi_{0}=\lambda^{-1}=1$ is called the t’ Hooft–Feynman gauge, etc.
+[30, 31, 32, 33].
+The formal $\xi_{0}=\lambda^{-1}=\infty$ limit is called as the canonical gauge in [32].
+This distinction seems a mere convention, but, nevertheless, it is useful one in QCD because of the presence of a mass scale parameter in its ground state.
+The generalized gauge requires that there is no other functional expression for $\xi$, apart from given by the relation (5.32) at finite $\xi_{0}=\lambda^{-1}$, in the full gluon propagator (5.34) and the ST identity (5.35) for the regularized gluon fields.
+The system of the regularized equations (5.34)-(5.35) constitutes that the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is not a symmetry of its ground state, and it has been exactly and uniquely defined.
+They are present in the form suitable for the NP renormalization program to be performed.
+The tadpole term enters the full gluon self-energy linearly, see the expression (3.12).
+However, in the full gluon propagator (5.34) it appears in the NL way, because its contribution has been summed up with the help of the gluon SD eq.
+(5.28).
+The term $(\Delta^{2}_{t}(D)/q^{2})$ in the expressions (5.34)-(5.34) is to be suppressed in the PT $q^{2}\rightarrow\infty$ at the finite that $\xi_{0}=\lambda^{-1}$ (the above-mentioned canonical gauge will be investigated in detail in the forthcoming paper).
+The full gluon propagator will behave like the free gluon propagator $\sim 1/q^{2}$.
+This is the one of the necessary constraints that a theory is perturbative renormalizable.
+Other ones such as the corresponding behaviour of the spinor Green’s function, a unitary of $S$-matrix and analyticity (causality) [2], evidently are beyond the scope of the present work.
+Here we have proven that the true dynamical and gauge structures of the QCD/YM ground state are much more complicated than it is required by its Lagrangian’s gauge symmetry.
+But despite the gauge symmetry is broken in the QCD ground state, its PT renormalizability is not affected within the QCD full gluon propagator in the generalized gauge (5.33), i.e., it has the PT renormalizable behaviour at large $q^{2}$.
+Concluding, it is worth noting that in terms of the transverse relations and the corresponding QD regularized constants as well as the corresponding relations between $\xi$ and $\xi_{0}$, the general solution can be shown explicitly as the following system of the relations, namely
+$q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)$ $=$ $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0,$
+$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)$ $=$ $q^{2}\Delta^{2}_{t}(D),$ (5.36)
 and
-$\xi=f(\xi_{0})=\xi_{0},\quad\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D^{PT})=T_{\rho\sigma}(q)q^{2}\Pi^{s}(q^{2};D^{PT}),$ (5.8)
-where $\Pi^{s}(q^{2};D^{PT})=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(g)}_{t}(q^{2};D^{PT})$, since the quark and gluon contributions coincide with their subtracted counterparts and are transverse in this case as well, i.e., $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$, due to the second equations in the relations (4.5)-(4.6).
-The invariant function $\Pi^{s}(q^{2};D^{PT})$ is a regular function at small $q^{2}$ and may be only logarithmically divergent at large $q^{2}$.
-The gluon SD eq. (2.5) becomes
-$D^{PT}_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)q^{2}\Pi^{s}(q^{2};D^{PT})D^{PT}_{\sigma\nu}(q),$ (5.9)
-and from it one arrives at $q_{\mu}q_{\nu}(q)D^{PT}_{\mu\nu}(q)=q_{\mu}q_{\nu}D^{0}_{\mu\nu}(q)=i\xi_{0}$, and thus the gauge-fixing parameter of the PT full gluon propagator is fixed as follows: $\xi=\xi_{0}$, in complete agreement with eqs.
-(5.8).
-Substituting decompositions (3.2) and (3.3) into the eq.
-(5.9), one finally obtains
-$D^{PT}_{\mu\nu}(q)={iT_{\mu\nu}(q)\over q^{2}[1+\Pi^{s}(q^{2};D^{PT})]}+i\xi_{0}L_{\mu\nu}(q){1\over q^{2}},$ (5.10)
-so that the gluon invariant function is $d_{PT}(q^{2})=[1+\Pi^{s}(q^{2};D^{PT})]^{-1}$.
-Obviously, this equation describes the propagation of the PT massless gluons, since it has the PT singularity on the mass-shell $q^{2}=0$, i.e., the singularity of the free gluon propagator $\sim 1/q^{2}$.
-Let us note in advance that the equality $\xi=\xi_{0}$ takes place only for the regularized massless gluon fields.
-For their renormalized counterparts these gauge-fixing parameters differ by the corresponding renormalization constant 2 ; 3 ; 4 ; 5 ; 6 ; 7 .
-The absence of a mass scale parameters in eq.
-(5.10) can be now attributed to the satisfied transverse relations (5.7)-(5.8).
-It is possible to say that just these ones decrease the quadratic UV divergences of the corresponding skeleton loop integrals to a logarithmic ones at large $q^{2}$.
-Due to the transverse relations (5.7)-(5.8) all the scale parameters, having the dimensions of a mass squared, shown up in eqs.
-(5.5)-(5.6), should be removed/disappeared from the theory, i.e., put formally zero.
+$\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0,\ \Delta^{2}_{t}(D)\neq 0,\ \xi\neq\xi_{0}=\lambda^{-1}.$ (5.37)
+It is necessary to underline that none of the prescriptions (i.e., putting some of them zero by hand) have been introduced and none of the truncations, approximations and assumptions have been done as well in the obtaining of these relations.
+In other words, all of them are exact mathematical results in the general solution to QCD.
+6 Particular solution to QCD
+The widely used up to present days the linear relation $\xi=\xi_{0}$ is only a particular solution of the derived here the exact constraint (3.15).
+As mentioned above in section 3, it is defined when the tadpole term has to be put zero by hand, i.e., $\Delta^{2}_{t}(D)=0$, then $\xi=\xi_{0}=\lambda^{-1}$ as well and vice versa.
+The corresponding expressions for the full gluon propagator and the ST identity in the particular solution to QCD can be independently derived absolutely in the same way as has been derived the corresponding expressions for the general solution to QCD in the previous section.
+However, the simplest way to show the particular solution is to put formally $\Delta^{2}_{t}(D)=0$ in the system of eqs.
+(5.34)-(5.35), then it looks like
+$D^{PT}_{\mu\nu}(q)={iT_{\mu\nu}(q)\over q^{2}[1+\Pi(q^{2};D^{PT})]}+i\lambda^{-1}L_{\mu\nu}(q){1\over q^{2}},$ (6.38)
+so that the gluon invariant function is $d_{PT}(q^{2})=[1+\Pi(q^{2};D^{PT})]^{-1}$ and formally neglecting the invariant function $\Pi(q^{2};D^{PT})$, i.e., putting it zero by hand $\Pi(q^{2};D^{PT})=0$, one obtains the free gluon propagator (3.9).
+The ST identities in this case are
+$q_{\mu}q_{\nu}D^{PT}_{\mu\nu}(q)=q_{\mu}q_{\nu}D^{0}_{\mu\nu}(q)=i\lambda^{-1}.$ (6.39)
+The corresponding gluon SD eq. (5.28) now becomes
+$D^{PT}_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{T}_{\mu\rho}(q),$ (6.40)
+where
+$D^{T}_{\mu\nu}(q)=D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)q^{2}\Pi(q^{2};D^{PT})D^{PT}_{\sigma\nu}(q).$ (6.41)
+From above and now on in this section we denote the corresponding full gluon propagator $D$ by $D^{PT}$ (for the explanation see the text below).
+Obviously, the expression for the PT full gluon propagator (6.38) describes the propagation of the PT massless gluons, since it has the PT singularity on the mass-shell $q^{2}=0$, i.e., the singularity of the free gluon propagator $\sim 1/q^{2}$.
+Let us note in advance that the equality $\xi=\xi_{0}=\lambda^{-1}$ takes place only for the regularized massless gluon fields.
+For their renormalized counterparts these gauge-fixing parameters differ by the corresponding renormalization constant.
 The system of the regularized eqs.(6.38)-(6.41), which is uniquely defined, is free of all the types of the scale parameters having the dimensions of a mass squared, forbidden by the exact gauge symmetry of the QCD Lagrangian.
 Therefore, it constitutes that the gauge symmetry of the QCD ground state, reflected by this system, coincides with the symmetry of its Lagrangian.
-Such coincidence is analogous to QED where the abelian $U(1)$ gauge symmetry of the Lagrangian is the same as of its ground state (omitting formally the YM degrees of freedom here, this coincidence has been proven just in this way in 10 ).
-However, it is necessary to remind that the coupling constant remains strong, so to call it as $D^{PT}$ is only a mere convention.
+Such coincidence is analogous to QED where the abelian $U(1)$ gauge symmetry of the Lagrangian is the same as of its ground state (omitting formally the YM degrees of freedom here, this coincidence has been proven just in this way in [10]).
+However, it is necessary to remind that the coupling constant remains strong, so to call it as $D^{PT}$ is only a mere convention, that is why the particular solution to QCD can be called as conventional QCD as well.
+Just this conventional solution to QCD has been investigated in full details the last five decades since its formulation in [1].
 The essential feature of the above-mentioned coincidence is the equality $\xi=\xi_{0}$.
@@
 This has nothing to do with the statement that the ST identity is a consequence of the exact gauge symmetry.
-It becomes its consequence, indeed, when $\xi$ is fixed by the relation (5.8).
-The well-known PT linear relation $\xi=\xi_{0}$ is the consequence of the exact gauge invariance both for the QCD Lagrangian and its ground state if it is treated in the PT framework for the regularized gluon fields.
-In principle, the PT QCD is a well-elaborated quantum gauge field theory 2 ; 3 ; 4 ; 5 ; 6 ; 7 ; 8 ; 9 , and it is not our concern.
-We present it here within our formalism for the readers convenience to directly compare it with our approach to the QCD below.
-6 Gauge symmetry is broken!
-QCD, being a non-abelian gauge theory with the strong coupling constant, is to be treated beyond the PT, especially its ground state.
-It contains much more information on the dynamical and gauge structures of the theory than its Lagrangian can provide at all 2 .
-So that the general case when $\xi\neq\xi_{0}$ should be investigated in detail, while keeping the ST identities (3.1) and (3.4) valid, anyway.
-The another independent requirement to satisfy the second one of the transverse relations (3.5) is defined by the condition
-${(\xi_{0}-\xi)\over\xi\xi_{0}}q^{2}-\Delta^{2}_{t}(D)=0.$ (6.1)
-Then for the initial gluon self-energy its transverse relation looks like
-$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)=q^{2}\Delta^{2}_{t}(D),$ (6.2)
-and it is not transverse, while the second line of the system (3.5) becomes transverse,
-$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]=0,$ (6.3)
-as expected and compare it with eq.
-(5.1).
-In the previous section we do not use that the contribution of the quark skeleton loop to the full gluon self-energy can be made always transverse, i.e., put $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=0$, independently from its gluon counterpart (this is an analogue to the satisfied transverse relation for the vacuum polarization tensor in QED, mentioned above).
-So that the satisfied transverse relation (6.3) can be effectively splintered into the two independent transverse relations, namely
-$q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=0,\quad q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0.$ (6.4)
-However, from the second lines of the tensor decompositions (4.3)-(4.4) and the satisfied transverse relations (6.4), one arrives to the conditions $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$, like in the previous case.
-Then the second relations in eqs (4.5) and (4.6) are to be reduced to
-$\Pi^{s(q)}_{l}(q^{2})={\Delta^{2}_{q}\over q^{2}},\quad\Pi^{s(g)}_{l}(q^{2};D)={\Delta^{2}_{g}(D)\over q^{2}},$ (6.5)
-but that is impossible, since the corresponding invariant functions cannot have pole-type singularities, by definitions, as explained in the previous section.
-One has to put these constants to zero on this general mathematical basis, i.e., $\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0$, and $\Pi^{s(q)}_{l}(q^{2})=\Pi^{s(g)}_{l}(q^{2};D)=0$ as well.
-All our results for the QD constants can be now summarized as follows:
-$\Delta_{q}^{2}=\Delta_{g}^{2}(D)=0,\quad\Delta^{2}(D)\rightarrow\Delta_{t}^{2}(D)\neq 0,$ (6.6)
-and hence
-$\Delta^{2}_{gh}=\Delta^{2}_{1}(D^{2})=\Delta^{2}_{2}(D^{4})=\Delta^{2}_{2^{\prime}}(D^{3})=0,$ (6.7)
-as well because of the relation (4.2).
-It is instructive to compare this system of the relations with the relations (5.5)-(5.6).
-The common feature of these relations is that the quark and gluon constants should be removed from the theory not only in the PT QCD but in the QCD as well.
-This means that the correspondence with the PT system (5.5)-(5.6) is maintained by the system (6.6)-(6.7).
-It is important for the renormalizability of the theory in the PT sense.
-The difference is that in the QCD the tadpole term remains intact.
-The essential dynamical source of this difference is that only the tadpole term can generate a mass squared scale parameter.
-All the other terms present in the skeleton loops decomposition for the full gluon self-energy, seen in Fig.
-1, cannot do this 6 , and thus none of their subtracted counterparts with their quark and gluon constants.The system of the relations (6.6)-(6.7) is not only a general but it is a unique one as well.
-They are not prescriptions, since based on a rigorous tensor algebra derivations, i.e., they are exact mathematical results.
-The relations (6.6) hold because these constants are related to the corresponding invariant functions for the subtracted counterparts, as described above and shown in the relations (6.5).
-Evidently, such kind of the relation does not exist for the constant tadpole term $\Delta^{2}_{t}(D)$, that is why it cannot be disregarded at all on this mathematical basis.
-All these constants, defined by the corresponding skeleton loop integrals, are QD at the upper limit.
-They have to be removed/disregarded from the theory on the general basis, i.e., formally put zero, as described above, apart from the tadpole term (2.4).
-It should remain intact in the QCD full gluon propagator, while in the PT QCD full gluon propagator it has to be removed as well.
-In what follows all these constants will be called the tadpole-like/type terms, for simplicity.
-The general question arises now, namely how to understand these exact equalities to zero in the relations (6.6)-(6.7), and hence in the relations (5.5)-(5.6) as well?
-Each tadpole-like term in these relations is the skeleton loop integral, which thus is a sum of the infinite number of terms.
-But an equality does not mean that the sum itself is zero.
-These equalities mean that any tadpole-like term which may appear in the theory by any possible way has to be discarded/disregarded in the theory, i.e, put formally zero, independently from any other tadpole-like terms.
-In this way such sums will be always zero, indeed.
-Such constants may appear even not as the result of the subtractions, but, for example as the result of the NL iteration procedure for the full gluon propagator shown in Fig.
-1.
-In this case the tadpole-like terms may be even multiplied by some regularized functions, see appendix A below.
-So that, whatever their origins would have been in the full gluon propagator, all of them belong to the infinite manifold of the relations (6.6)-(6.7).
-Collecting all our results obtained above for the quark and gluon contributions to the gluon self-energy, one obtains
-$\Pi^{q}_{\rho\sigma}(q)=T_{\rho\sigma}(q)q^{2}\Pi^{s(q)}_{t}(q^{2}),\quad\Pi^{g}_{\rho\sigma}(q;D)=T_{\rho\sigma}(q)q^{2}\Pi^{s(g)}_{t}(q^{2};D),$ (6.8)
-since they become transverse and coincide with their subtracted counterparts.
-Let us also remind that the invariant functions present in the previous relations are regular functions at small $q^{2}$, i.e., they have no pole-type singularities, and may be only logarithmically divergent at large $q^{2}$.
-Substituting their sum into the eq.
-(2.5), one gets
-$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)$ $+$ $D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)q^{2}\Pi^{s}_{t}(q^{2};D)D_{\sigma\nu}(q)$ (6.9)
-$+$ $D^{0}_{\mu\rho}(q)i\delta_{\rho\sigma}\Delta^{2}_{t}(D)D_{\sigma\nu}(q),$
-where $\Pi^{s}_{t}(q^{2};D)=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(g)}_{t}(q^{2};D)$.
-The corresponding ST identities therefore are not equal to each other, namely $q_{\mu}q_{\nu}(q)D_{\mu\nu}(q)\neq q_{\mu}q_{\nu}D^{0}_{\mu\nu}(q)$, which implies $\xi\neq\xi_{0}$.
-The gauge-fixing parameter $\xi$ as a function of $\xi_{0}$ is given now by the condition (6.1).
-Contrary to the gluon SD eq.
-(5.9), the gluon SD eq.
-(6.9) is now equivalent to the initial gluon SD eq.
-(2.5), only being re-written in the different form.
-That is the reason why we left it denoted as $D$ here and everywhere below.
-Evidently, the PT gluon SD eq.
-(5.9) is not equivalent to eq.
-(2.5).
-The both full gluon SD eqs.
-(2.5) and (6.9) by keeping the mass squared parameter explicitly - the tadpole term $\Delta^{2}_{t}(D)$ constitutes the fact that the gauge symmetry of the QCD ground state is not a symmetry of its Lagrangian.
-The PT renormalizability of the theory seems to be compromised, however see arguments below.
-It is important to emphasize once more that we are under an obligation to satisfy at least one of the transverse relations (3.5).
-Reminding that this is necessary to do in order to get finally to eq.
-(5.9) in the PT QCD and eq.
-(6.9) in the QCD.
-The last one allows to distinguish between the quark $\Delta^{2}_{q}$ and the gluon $\Delta^{2}_{g}(D)$ constants, which are not present in the initial gluon SD eq.
-(2.5) and the tadpole constant $\Delta^{2}_{t}(D)$, which is explicitly present in it.
-The satisfied transverse relations have been implemented into the our formalism in a self-consistent way.
-The fact of the matter is that the system of the relations (6.2)-(6.3) and (6.8)-(6.9) have been uniquely and exactly derived.
-Then the condition (6.1) can be recovered from the gluon SD eq.
-(6.9).
-7 The QCD full gluon propagator in the generalized gauge
-The basic relations to which the full gluon propagator and its free counterpart should satisfy are the corresponding ST identities (3.1) and (3.4), respectively.
-As emphasized above in section 3, they are important for the general renormalizability of the theory, and thus provide the exact constraints on any solution to QCD 2 .
-But by themselves they cannot remove the quadratic UV divergences from the theory, as it has been shown above in the relations (6.6)-(6.7), apart from the tadpole term.
-We have achieved this by formulating the proper subtraction scheme in which the corresponding transverse conditions have been implemented.
-If some equations, relations or the regularization schemes, etc.
-do not satisfy the ST identities automatically, i.e., without any additional conditions, then they should be modified and not the identity (3.1).
-In other words, all the relations, equations, regularization schemes, etc.
-should be adjusted to it and not vice versa.
-For example, the transverse condition for the full gluon self-energy (6.2) is not satisfied, but, nevertheless, the general form of the ST identity (3.1) is to be maintained in any case, see eq.
-(7.3) below.
-Let us begin with the gluon SD eq.
-(6.9), which can be equivalently re-written as follows:
-$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)$ $+$ $D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)\left[q^{2}\Pi(q^{2};D)+\Delta^{2}_{t}(D)\right]D_{\sigma\nu}(q)$ (7.1)
-$+$ $D^{0}_{\mu\rho}(q)iL_{\rho\sigma}(q)\Delta^{2}_{t}(D)D_{\sigma\nu}(q),$
-saving the tadpole term in the longitudinal component of eq.
-(6.9), it makes it possible for the tadpole term to appear in the transverse component of the full gluon propagator (7.1) as well.
-The invariant function $\Pi(q^{2};D)$ is a replacement (for simplicity) of the initial one $\Pi^{s}_{t}(q^{2};D)$ in eq.
-(6.9).
-It is worth remanding that it is regular at zero and may have only logarithmic divergences in the PT $q^{2}\rightarrow\infty$ limit.
-Combining this equation with the decompositions (3.2) and (3.3), one obtains
-$d(q^{2})={1\over 1+\Pi(q^{2};D)+(\Delta^{2}_{t}(D)/q^{2})}.$ (7.2)
-This relation is not a solution for the full gluon invariant function.
-It is the NL transcendental equation for the different invariant functions $d(q^{2}),\ \Pi(q^{2};D)$ and the constant $\Delta^{2}_{t}(D)$, i.e., $d=f(D(d))$.
-Nevertheless, from this expression is clearly seen that in the PT $q^{2}\rightarrow\infty$ regime, the contribution $(\Delta^{2}_{t}(D)/q^{2})$ can be neglected, but the invariant function $\Pi(q^{2};D)$ may still depend on this ratio under the PT logarithms.
-At the same time, in the NP region of finite and small gluon momenta this term is dominant, and the dependence of $d(q^{2})$ on $\Delta^{2}_{t}(D)$ may be much more complicated due to the transcendental character of eq.
-(7.2).
-Contracting the full gluon SD eq. (7.1) with $q_{\mu}$ and $q_{\nu}$, and substituting its result into the general ST identity (3.1), one arrives at
-$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi_{0}\left(1-\xi{\Delta^{2}_{t}(D)\over q^{2}}\right)=i\xi,$ (7.3)
-which solution is
-$\xi\equiv\xi(q^{2};\xi_{0})={\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)},$ (7.4)
-i.e., in this case the gauge-fixing parameter becomes the function $\xi\equiv\xi(q^{2};\xi_{0})$ and thus it is not a constant equal to $\xi_{0}$.
-Let us point out that the expression (7.4) satisfies condition (6.1), which is as it should be.
-It is worth emphasizing that the solution (7.4) shows the NL dependence on $\xi_{0}$, which is only possible if gauge symmetry is broken in the ground state.
-From eq.
-(7.4) it is clearly seen that in the PT $q^{2}\rightarrow\infty$ regime, which effectively equivalent to the formal $\Delta^{2}_{t}(D)=0$ limit and vice versa, the gauge-fixing parameter $\xi$ goes to $\xi_{0}$, as expected.
-The generalized gauge (7.4) is the known function of its arguments, while the relation (7.2) is the NL transcendental one.
-Thus, behind the general inequality $\xi\neq\xi_{0}$ is the constant $\Delta^{2}_{t}(D)$ as its dynamical source, when its contribution can be neglected only then $\xi=\xi_{0}$.
-The equality holds only for the regularized gluon fields, as underlined above in section 5.
-Substituting equations (7.2) and (7.4) into the general decomposition (3.2) for the full gluon propagator, one finally obtains
-$D_{\mu\nu}(q)={iT_{\mu\nu}(q)\over q^{2}+q^{2}\Pi(q^{2};D)+\Delta^{2}_{t}(D)}+iL_{\mu\nu}(q){\lambda^{-1}\over q^{2}+\lambda^{-1}\Delta^{2}_{t}(D)},$ (7.5)
-[…]
-The corresponding ST identity now becomes
-$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi(q^{2};\lambda^{-1})=i{\lambda^{-1}q^{2}\over(q^{2}+\lambda^{-1}\Delta^{2}_{t}(D))}.$ (7.6)
-This is the generalized gauge since it depends on the constant $\Delta^{2}_{t}(D)$, and when it is zero, one recovers the gauge-fixing parameter for the free gluon propagator.
-To our best knowledge the generalized gauge (7.6) has been derived for the first time in a such new manner.
-In the generalized gauge the gauge-fixing parameter $\lambda^{-1}$ is convenient to vary continuously from zero to infinity.
-The functional dependence of the generalized gauge-fixing parameter $\xi$ (7.4) is fixed up to an arbitrary gauge-fixing parameter $\xi_{0}=\lambda^{-1}$.
-Unless we fix it, and thus $\xi$ itself, we will call such situation as the general gauge dependence (GGD), see equations (3.1), (7.5), and (7.6).
-Choosing $\xi_{0}=\lambda^{-1}$ explicitly, we will call such situation as the explicit gauge dependence (EGD).
-For example, $\xi_{0}=\lambda^{-1}=0$ is called the unitary (Landau) gauge, $\xi_{0}=\lambda^{-1}=1$ is called the t’ Hooft–Feynman gauge, etc.
-30 ; 31 ; 32 ; 33 .
-The formal $\xi_{0}=\lambda^{-1}=\infty$ limit is called as the canonical gauge in 32 .
-This distinction seems a mere convention, but, nevertheless, it is useful one in QCD because of the presence of a mass scale parameter in its ground state.
-The generalized gauge requires that there is no other functional expression for $\xi$, apart from given by the relation (7.4) at finite $\xi_{0}=\lambda^{-1}$, in the full gluon propagator (7.5) and the ST identity (7.6) for the regularized gluon fields.
-The system of the regularized equations, consisting of the gluon SD eq.
-(7.1), the expression (7.5) and the ST identity (7.6), constitutes that the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is not a symmetry of its ground state, and it has been exactly and uniquely defined.
-The tadpole term enters the full gluon self-energy linearly, see the expression (2.2).
-However, in the full gluon propagator (7.5) it appears in the NL way, because its contribution has been summed up with the help of the gluon SD eq.
-(7.1).
-In the PT $q^{2}\rightarrow\infty$ limit, the term $(\Delta^{2}_{t}(D)/q^{2})$ in the expression (7.5) is to be suppressed, and it will be reduced to the PT QCD expression (5.10), reminding that $\xi_{0}=\lambda^{-1}$.
-The full gluon propagator will behave like the PT QCD gluon propagator at short distance ($q^{2}\rightarrow\infty$), namely $\sim 1/q^{2}$.
-This is the one of the necessary constraints that a theory is perturbative finite.
-Other ones such as the corresponding behaviour of the spinor Green’s function, a unitary of $S$-matrix and analyticity (causality) 2 , evidently are beyond the scope of the present paper.
-Here we have proven that the true dynamical and gauge structures of the QCD/YM ground state are much more complicated than it is required by its Lagrangian’s gauge symmetry.
-Despite the gauge symmetry is broken in the QCD ground state, its PT renormalizability is not affected within the QCD full gluon propagator in the generalized gauge (7.5), i.e., it has the PT renormalizable behaviour at large $q^{2}$.
+It becomes its consequence, indeed, when $\xi$ is fixed by the relation $\xi=\xi_{0}=\lambda^{-1}$.
+The well-known PT linear relation $\xi=\xi_{0}$ is the consequence of the exact gauge invariance both for the QCD Lagrangian and its ground state.
+In principle, the PT QCD or, equivalently, the conventional QCD is a well-elaborated quantum gauge field theory [2, 3, 4, 5, 6, 7, 8, 9], and for its framework of the corresponding dynamical equations of motion see recent review [34] (and references therein).
+However, this theory is not our concern (but for some clarifying discussion see appendix B).
+We present it here within our formalism for the readers convenience to directly compare it with the general solution to QCD.
+For this aim it is useful to reproduce the system of the relations (5.36)-(5.37) for the PT QCD. It will look like
+$q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)$ $=$ $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D^{PT})=0,$
+$q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D^{PT})$ $=$ $0,$ (6.42)
+and
+$\Delta^{2}_{q}=\Delta^{2}_{g}(D^{PT})=0,\ \Delta^{2}_{t}(D^{PT})=0,\ \xi=\xi_{0}=\lambda^{-1}.$ (6.43)
+The first equalities in these relations coincide with the corresponding equalities in the relations (5.36)-(5.37), i.e., in both theories the general and the PT QCD or, equivalently, the particular QCD, they are exact mathematical results.
+At the same time, the second equalities in these relations remain prescriptions, since they have to be put zero by hand, as emphasized above.
+The dimensional regularization method (DRM) [4, 6, 27, 28, 29], formulated within the framework of the conventional QCD, provides a gauge-invariant scheme to correctly calculate the finite parts of the generally QD loop integrals, while omitting their QD constants, i.e., simply to ignore them and to deal further only with the logarithmic divergences of the PT.
+This was a prescription rather than an exact result, as pointed out in [4].
+However, from now on this prescription has been put on a firm mathematical ground.
+How to understand such kind of equalities has been explained in section 5.
 7 Discussion
-Let us now discuss some interesting features of the explicit presence of the tadpole term in the QCD ground state, i.e., at the fundamental quark-gluon level.
-First of all, let us underline once more that the tadpole term, existing in the QCD ground state, cannot be discarded in the QCD/YM theory by any means.
+Let us now discuss some interesting physical consequences of the general solution to QCD.
+It requires the explicit presence of the tadpole term in the QCD ground state at the fundamental quark-gluon level, i.e., it cannot be discarded in the QCD/YM theory by any means.
 This was exactly proven in section 4 and used in section 5.
@@
 For the sake of transparency and clarity let us use the free gluon propagator in the t’ Hooft-Feynman gauge $\lambda^{-1}=1$, i.e., $D^{0}_{\mu\nu}(q)=i{\delta_{\mu\nu}/q^{2}}$.
-Then the full gluon propagator (7.1) up to the fist iteration term looks like
-$D_{\mu\nu}(q)\sim i{\delta_{\mu\nu}\over q^{2}}-i{T_{\mu\nu}(q)\over q^{2}}\Bigl{[}\Pi(q^{2};D)+{\Delta^{2}_{t}(D)\over q^{2}}\Bigr{]}-iL_{\mu\nu}(q){\Delta^{2}_{t}(D)\over(q^{2})^{2}}+...\ ,$ (8.1)
-and it is worth reminding that the invariant function $\Pi(q^{2};D)$ is a regular one of its argument, while all the tadpole-like terms to one loop approximation have been omitted due to the relations (6.6)-(6.7).
+Then the full gluon propagator (5.28), on account of the definitions (5.29) and (5.30), up to the first iteration term looks like
+$D_{\mu\nu}(q)\sim i{\delta_{\mu\nu}\over q^{2}}-i{T_{\mu\nu}(q)\over q^{2}}\Pi(q^{2};D)-iT_{\mu\nu}(q){\Delta^{2}_{t}(D)\over(q^{2})^{2}}+...\ ,$ (7.44)
+where the invariant function $\Pi(q^{2};D)$ is a regular one of its argument, while all the tadpole-like terms up to one loop approximation have been omitted due to the relations (4.26).
 See appendix A as well, where this iteration is discussed in more detail.
-Let us also remind that the gluon eq.
-(7.1) is equivalent to the initial gluon SD eq.
-(2.1), shown in Fig.
-1, and hence to eq.
-(2.5), only being re-written down in the form suitable for the NP renormalization program to be performed.
+Also here and below not shown the terms proportional to the longitudinal component $L_{\mu\nu}(q)$ since finally they will be summed up to the known function, see expression (5.33).
 The higher-order iteration terms will be much more singular with respect to the general ratio $(\Delta^{2}_{t}(D)/q^{2})$, since each iteration invokes additional powers of this ratio.
-Because of the NL character of the iteration procedure any next iteration will contribute into the previous iteration terms, and thus they become the sums of the infinite number of terms 34 , namely
-$D_{\mu\nu}(q)\sim iT_{\mu\nu}{\Delta^{2}_{t}(D)\over(q^{2})^{2}}\sum_{k=0}^{\infty}\Bigl{(}{\Delta^{2}_{t}(D)\over q^{2}}\Bigr{)}^{k}\Phi_{k}(g^{2},\tilde{\lambda}...)+O_{\mu\nu}(1/q^{2}).$ (8.2)
-For simplicity, in this iteration expansion is explicitly shown only singular part of the corresponding Laurent series, and thus it represents the summation of all the severe IR singularities (more singular than the PT IR singularity $\sim 1/q^{2}$), which can be accumulated by the full gluon propagator.
+Because of the NL character of the iteration procedure any next iteration will contribute into the previous iteration terms, and thus they become the sums of the infinite number of terms [35], namely
+$D_{\mu\nu}(q)\sim iT_{\mu\nu}{\Delta^{2}_{t}(D)\over(q^{2})^{2}}\sum_{k=0}^{\infty}\Bigl{(}{\Delta^{2}_{t}(D)\over q^{2}}\Bigr{)}^{k}\Phi_{k}(g^{2},...)+...,$ (7.45)
+where the terms of order $O_{\mu\nu}(1/q^{2})$ are not shown, for simplicity.
+In this iteration expansion is explicitly shown only singular part of the corresponding Laurent series, and thus it represents the summation of all the severe IR singularities (more singular than the PT IR singularity $\sim 1/q^{2}$), which can be accumulated by the full gluon propagator.
 Evidently, such a summation is only possible due to the existence of the tadpole term.
-We have shown the singular part, associated with the transverse component of the full gluon propagator, since its longitudinal part is the regular function of the tadpole term, see expression (7.5).
-The dependence of the coefficients $\Phi_{k}(g^{2},\tilde{\lambda}...)$ on the coupling constant is not known, i.e., such kind of series are the cluster expansions, mentioned above 23 .
-These coefficients, being the sum of the infinite number of terms by themselves, are resulted from the skeleton loop integrals, contributing into full gluon self-energy through the invariant function $\Pi(q^{2};D)$.
+The dependence of the coefficients $\Phi_{k}(g^{2},...)$ on the coupling constant is not known, i.e., such kind of series are the cluster expansions, mentioned above [23].
+These coefficients, being the sum of the infinite number of terms by themselves, are resulted from the skeleton loop integrals, contributing into full gluon self-energy through the invariant function $\Pi(q^{2};D)$ because of the NL character of the iteration procedure.
 Evidently, the iteration ab  infinity leads to the essential singularity at $q^{2}\rightarrow 0$!
@@
 Let us note that the same series can be obtained by expanding the initial NL transcendental relation (5.30) in powers of the ratio $(\Delta^{2}_{t}(D)/q^{2})$.
-From the NL iteration expression (8.2) it follows that the full gluon propagator in the IR region ($q^{2}\rightarrow 0$) is dominated by the tadpole contribution, while in the UV region ($q^{2}\rightarrow\infty$) it will be dominated by the free gluon one.
+From the NL iteration expressions (7.44)-(7.45) it follows that the full gluon propagator in the IR region ($q^{2}\rightarrow 0$) is dominated by the tadpole contribution, while in the UV region ($q^{2}\rightarrow\infty$) it will be dominated by the free gluon one.
 This is completely different to QED, where the dressed photon propagator behaves at small photon momentum as its free counterpart ($\sim 1/q^{2}$).
@@
 That is why the dressed PT gluon (6.38) and free gluon states can not appear in the physical spectrum at large distances ($q^{2}\rightarrow 0$), i.e., they will be suppressed in this region by the tadpole term.
-If this statement will survive the corresponding renormalization program beyond the PT then confinement of the PT and free gluons will be explained.
+If this statement will survive the corresponding renormalization program beyond the PT then confinement of the PT and free gluon states will be explained.
 This is important since there is no any other dynamical mechanism to confine these gluon states to the vacuum of QCD or to keep them inside hadrons.
-To prove the confinement of the singular full gluon propagator itself due to the above-mentioned essential singularity, though possible, but requires completely different investigation within the theory of distributions 26 , and thus is beyond the scope of the present work.
+To prove the confinement of the singular full gluon propagator itself due to the above-mentioned essential singularity, though possible, but requires completely different investigation within the theory of distributions [26], and thus is the subject of the future work.
 Let us note that by the NP program we mean that the full gluon propagator will become free from all the non-physical parameters of the theory in any of its possible solutions (not only in the iteration one).
 Eq.
-(7.5) shows that gluon may acquire mass dynamically if the denominator of this equation becomes zero at some finite point $q^{2}=M^{2}_{g}$.
+(5.34) shows that gluon may acquire mass dynamically if the denominator of this equation becomes zero at some point $q^{2}=M^{2}$.
 The expression for the free massive vector particle propagator follows in a rather simple way, indeed.
-For this goal it is necessary to omit the corresponding invariant function and identify the tadpole term with the gluon pole mass $M^{2}_{g}$.
+For this goal it is necessary to omit the corresponding invariant function and identify the tadpole term with the gluon mass $M^{2}$.
 In Minkowski metric (by replacing $q^{2}\rightarrow-q^{2}$ and $\delta_{\mu\nu}\rightarrow g_{\mu\nu}$), one gets
-$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2}_{g})}\left[g_{\mu\nu}-(1-\lambda^{-1}){q_{\mu}q_{\nu}\over(q^{2}-\lambda^{-1}M^{2}_{g})}\right],$ (8.3)
-i.e., it will be expressed in the Stueckelberg gauge 4 .
+$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2})}\left[g_{\mu\nu}-{(1-\lambda^{-1})q_{\mu}q_{\nu}\over(q^{2}-\lambda^{-1}M^{2})}\right],$ (7.46)
+i.e., it will be expressed in the Stueckelberg gauge [4].
 Contracting this expression with respect to $q_{\mu}$ and $q_{\nu}$, one obtains the ST identity (5.35) by replacing there $q^{2}\rightarrow-q^{2}$.
-In other words, the Stueckelberg gauge (8.3) is a particular case of the generalized gauge (7.6).
-The PT QCD full gluon propagator (5.10) cannot generate the gluon pole mass at some finite point, since it remains always massless.
+In other words, the Stueckelberg gauge (7.46) is a particular case of the generalized gauge-fixing function (5.33).
+The PT QCD full gluon propagator (6.38) cannot generate the gluon mass at some point, since it remains always massless.
 The full gluon propagator (5.34) provides a rather convenient framework in order to formulate and further develop the NP renormalization program for the massive gluon fields.
-This program will suppose to show that the full gluon propagator (7.5), and hence its free massive counterpart (8.3), can not be put on the mass-shell $q^{2}=M^{2}_{g}$.
+This program will suppose to show that the full gluon propagator (5.34), and hence its free massive counterpart (7.46), can not be put on the mass-shell $q^{2}=M^{2}_{g}$, where $M^{2}_{g}$ is the finite gluon pole mass.
 Acquiring mass dynamically, the gluon, nevertheless, will remain the off-mass-shell object.
-So that it will not appear in the physical spectrum (confinement of the massive gluon fields within our approach).
+So that, it will not appear in the physical spectrum.
+Some aspects of the massive gluon field’s renormalization are discussed in appendix C.
 In the presence of the mass squared scale parameter - the tadpole term - the role of the QCD coupling constant $g^{2}$ becomes unimportant.
-This is also evidence of the ’dimensional transmutation’, $g^{2}\rightarrow\Delta^{2}_{t}(D)$ 2 ; 35 ; 36 , which occurs whenever a massless theory acquires mass dynamically.
+This is also evidence of the ’dimensional transmutation’, $g^{2}\rightarrow\Delta^{2}_{t}(D)$ [2, 36, 37], which occurs whenever a massless theory acquires mass dynamically.
 It is a general feature of spontaneously symmetry breaking in field theories.
-We distinguish between the PT and full QCD by the explicit presence of the tadpole term in the latter one, and not by the magnitude of the coupling constant even at the regularized gluon fields level yet.
-In the both cases the gluon fields remain strongly interacted, apart from the asymptotic freedom (AF) regime 37 ; 38 .
-Any deviation of the full gluon propagator from the free one in QCD requires the presence of a mass squared scale parameter. Even in the AF regime there is a scale breaking,
-$D_{\mu\nu}(q)\sim\delta_{\mu\nu}\Bigl{[}{ig^{2}\over 1+g^{2}b_{0}\ln(q^{2}/\Lambda^{2}_{QCD})}\Bigr{]}(1/q^{2}),\quad q^{2}\rightarrow\infty,$ (8.4)
-which has been discovered many years ago 37 ; 38 .
+We distinguish between the PT/particular/conventional and general solutions to QCD by the explicit presence of the tadpole term in the latter one, and not by the magnitude of the coupling constant even at the regularized gluon fields level yet.
+In the both cases the gluon fields remain strongly interacted, apart from the AF regime [38, 39].
+So that, the general solution to QCD, investigated in section 5 is possible to call as the NP QCD as well.
+Any deviation of the full gluon propagator from the free one in QCD requires the presence of a mass squared scale parameter. Even in the AF regime $q^{2}\rightarrow\infty$ there is a scale breaking,
+$D_{\mu\nu}(q)\sim\delta_{\mu\nu}\Bigl{[}{ig^{2}\over 1+g^{2}b_{0}\ln(q^{2}/\Lambda^{2}_{QCD})}\Bigr{]}(1/q^{2}),$ (7.47)
+which has been discovered many years ago [37, 38].
 When $q^{2}\rightarrow\infty$ the dependence on the coupling constant $g^{2}$ vanishes, indeed.
-The expression (8.4) is written down in the t’ Hooft-Feynman gauge and in the approximation of the main PT logarithms with the colour group factor $b_{0}>0$.
+The well-known expression (7.47) is written down in the t’ Hooft-Feynman gauge and in the approximation of the main PT logarithms with the colour group factor $b_{0}>0$.
 $\Lambda^{2}_{QCD}$ determines the scale of the non-trivial PT dynamics in the QCD.
 However, in section 6, we have already shown that the PT cannot generate the scale parameter, having the dimension of a mass squared.
-From the renormalization group equations analysis, it follows that any mass to which can be assigned any physical meaning is exponentially suppressed in the weak coupling limit to every order of the PT 2 ; 35 , namely
+From the renormalization group equations analysis, it follows that any mass to which can be assigned any physical meaning is exponentially suppressed in the weak coupling limit to every order of the PT [2, 36], namely
 $M\sim\mu\exp(-1/b_{0}g^{2}),\quad\quad g^{2}\rightarrow 0,$ (7.48)
@@
 It is entirely of the NP origin!
-It has to come from the IR region which is controlled by the tadpole term, as discusses just above in this section.
+It has to come from the IR region which is controlled by the tadpole term, as discussed just above.
 Its finite version should be somehow related to $\Lambda^{2}_{QCD}$ (reminding that the ratio $(\Delta^{2}_{t}(D)/q^{2})$ may appear under the PT logarithms).
+The explanation of the AF phenomenon (i.e., a scale breaking in the PT regime) will be given when the NP renormalization program of the singular gluon field configurations will be performed.
 8 Conclusions
-Our formalism makes it possible to remove from the theory the QD constants of the gluon degrees of freedom in the same way as the QD constants of the quark degrees of freedom.
-This has lead to the discovery that in addition to the PT gluon field configurations in the QCD ground state (more precisely in its YM sector), preserving the exact gauge symmetry of its Lagrangian, there also exists such a type of the gluon field configurations which do not maintain this symmetry.
-But the PT renormalizable properties of the theory will not be affected.
-The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared.
+We have derived the novel exact constraint (3.15) on any solution to QCD.
+For this aim the ST identities for the gauge particles Green’s functions (propagators), complemented by the satisfied transverse relations for the quark and gluon contributions into the full gluon self-energy, have been used.
+The first of its solutions is called the general solution to QCD and given by the expression (3.16), which formulates the generalized gauge-fixing function approach to QCD.
+The existence of a new general solution to QCD makes it possible to prove in the most general and unique mathematical ways that the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is broken in its ground state.
+However, the PT renormalizability will not be affected within the framework of the full gluon propagator.
+In the PT $q^{2}\rightarrow\infty$ regime it will have the PT renormalizable behaviour $\sim 1/q^{2}$.
+The dynamical source of the Lagrangian’s gauge symmetry breakdown in the ground state has to be identified with the constant tadpole term, having the dimension of a mass squared.
 It is dynamically generated by the self-interaction of the multiple massless gluon modes involving only the point-like 4-gluon vertex, see eq.
@@
 1.
-The full 4-gluon vertex survives when all the gluon momenta involved go to zero, becoming thus its point-like counterpart, while the full 3-gluon vertex does not survive in this limit 10 ; 23 ; 39 .
+The full 4-gluon vertex survives when all the gluon momenta involved go to zero, becoming thus its point-like counterpart, while the full 3-gluon vertex does not survive in this limit.
+Just this shows the importance of the point-like 4-gluon vertex in the deep IR structure [10, 23, 40] of the theory.
 By the design the tadpole term is present in the QCD ground state from the very beginning, i.e., not introducing by hand.
@@
 1), while all the other tadpole-like terms don’t.
-Unlike them, being the QD but already regularized constant, the tadpole term is not connected to any external gluon momentum, and thus may generate a mass squared scale parameter 6 , indeed.
+Unlike them, being the QD but already regularized constant, the tadpole term is not connected to any external gluon momentum, and thus may generate a mass squared scale parameter [6], indeed.
 Thus, there is no need to extract such a scale parameter in the QCD vacuum in a some other way.
-In 10 the importance of the tadpole term has been recognized and exploited on an intuitive basis rather than a logical one.
+In [10] the importance of the tadpole term has been recognized and exploited on an intuitive basis rather than a logical one.
 However, from now on its existence cannot be undermined.
@@
 Its renormalized (i.e., finite) version will determine a scale of the NP dynamics in QCD and in its true ground state.
-For the both types of the gluon field configurations have been derived the corresponding expressions for the full gluon propagators describing their propagation in the vacuum of QCD.
-For the second type the corresponding NP QCD full gluon propagator (7.5) is written down in the generalized gauge (7.6).
-From these expressions it is clearly seen that in the PT $q^{2}\rightarrow\infty$ limit eq.
-(7.5) coincides with eq.
-(5.10).
-This is an evidence that the PT renormalizable properties of the theory will not be affected by the new type of the gluon field configurations.
-In other words, the renormalizability of the theory will be maintained within our approach to the full gluon propagator.
-It is worth emphasizing once more that the massive and severely singular gluon field configurations can be described by eq.
-(7.5) only, while eq.
-(5.10) describes the propagation of the massless gluon modes with the PT IR singularities $\sim 1/q^{2}$ only.
-Our ground-breaking findings are as follows:
-a. One of the transverse relations (3.5) is to be necessarily satisfied, in order to make the theory renormalizable in the PT sense.
-b. The subtraction scheme has been formulated, which makes it possible to remove from theory all the quadratically UV divergent constants on the general basis.
-c.
-The exact gauge symmetry of the YM Lagrangian is not a symmetry of its ground state.
-The dynamical source of this effect is the tadpole term.
-If it formally put zero then these gauge symmetries coincide.
-d. The system of the relations (6.6)-(6.7) are exact mathematical results, i.e., they are not prescriptions.
+Due to these impacts our advance findings are:
+a.
+The novel general constraint (3.15) on any solution to QCD has been exactly derived before any proper subtraction scheme has been introduced.
+The importance of the ST identities and the corresponding satisfied transverse relations has been explicitly shown.
+b. Its exact solution (3.16) determines the generalized gauge-fixing function approach to QCD.
+c. The proper subtraction scheme has been formulated to remove from the theory the QD constants, apart from the tadpole term, on the general mathematical basis.
+d. The system of the relations (5.36) are exact mathematical results, i.e., they are not prescriptions.
 e.
-The new NP approach to the YM theory and its ground state has been formulated, by keeping the corresponding ST identities always valid.
-It is based on the system of eqs.
-(7.5)-(7.6), which is defined up to the corresponding invariant function with the known asymptotic properties.
-Eqs.
-(7.5)-(7.6) are suitable for the NP renormalization program to be performed.
-f. The NL dependence $\xi$ on $\xi_{0}$ in the generalized gauge (7.4) in the PT $q^{2}\rightarrow\infty$ or, equivalently, formal $\Delta^{2}_{t}(D)=0$ limits becomes linear one $\xi=\xi_{0}$.
-g. The DRM 6 ; 25 ; 26 ; 27 prescription to simply ignore all the QD constants (including the tadpole term) in the PT QCD has been put on a firm mathematical ground from now on, see eqs. (5.4)-(5.6).
-In obtaining all these results no specific regularization schemes have been used.
-Neither special gauge choice nor any truncations, approximations and assumptions have been made either in dealing with the regularized skeleton loop integrals, contributing to the gluon self-energy.
-The exact analytic derivations have been done, based on a rigorous tensor algebra rules, such as the decomposition into the independent structures, the subtractions, etc.
-10 Summary
-It is shown in the most general and unique way that the gauge symmetry of the QCD Lagrangian is broken in its ground state, indeed.
-This has been achieved without solving the full gluon SD eq.
-(2.5), but recovering only its true dynamical and gauge structures in more detail in the framework of our formalism.
-So that, the significant peace of the information on these important issues was not missed within our general approach.
-Its breakthrough results can be summarized in terms of the following theorem:
-Tadpole term existence and the true QCD ground state: Prove that in the QCD ground state exist such gluon field configurations which break/violet the gauge symmetry of its Lagrangian, but the PT renormalizability of the theory is maintained.
-Existence of this effect includes establishing the true role of the tadpole term in the QCD ground state.
-Just this theorem has been proven above in this paper, which resolves the dilemma formulated at the end of section 2.
-In the proof of this theorem we have shown how to save the tadpole term in QCD and its ground state, without affecting the PT renormalizable properties of the theory, while removing all the other QD constants from it.
-The new derived solution with the gauge symmetry breaking will apparently satisfy the bounder conditions in QCD at the fundamental quark-gluon level: confinement at large distances ($q^{2}\rightarrow 0$) and AF at short distances ($q^{2}\rightarrow\infty$), as discussed in section 8.
-The solution with the gauge symmetry preservation can not satisfy them.
-All this underlines once more the importance of the theorem above.
-The perspectives for further work are as follows: to formulate the NP multiplicative (MP) renormalization programs for the full gluon propagator (7.5), which is only regularized one at this stage.
-Unfortunately, this is not a simple task, since such kind of a programs have to be different for the massive and severely singular gluon field configurations, briefly discussed in section 7.
+The new NP approach to the YM theory and its ground state has been formulated, based on the regularized system of eqs.
+(5.34)-(5.35).
+It is defined up to the invariant function with the known asymptotic properties.
+f.
+This system has a correct the PT limit, and is suitable for the NP renormalization program to be performed.
+It is called as the general solution to QCD, and just it resolves the dilemma formulated at the end of section 2.
+In obtaining all these results no any truncations, approximations and assumptions have been made in dealing with the regularized skeleton loop integrals, contributing to the gluon self-energy.
+The exact analytic derivations have been done, based on a tensor algebra rigorous rules.
+The second of the above-mentioned solutions is called the particular solution to QCD, and determined by the relations $\Delta^{2}_{t}(D)=0$ and $\xi=\xi_{0}$ in the exact constraint (3.15).
+This solution is nothing else but the well-known PT or, equivalently, conventional QCD, which has been studied last five decades.
+We confirm that the second solution preserves the gauge symmetry of the QCD Lagrangian in its ground state.
+Both equalities, shown above, are simply prescriptions, i.e., they have to be put valid by hand, as it has been discussed in section 6.
+In this connection, let us remind that for the tadpole term does not exist the independent mathematical relation according to which it has to be omitted from the theory, while such kind of the relations exist for the the quark and gluon QD regularized constants (see section 4).
+That is why the tadpole term has to be put zero by hand for the particular solution, which comes from the solution of the exact constraint (3.15), and see also discussion in appendix B.
+We can registered the existence of the two different types of the gluon field configurations in the QCD vacuum: the general one and its particular counterpart with the gauge symmetry broken and unbroken, respectively.
+The just derived general solution to QCD with a gauge symmetry breaking, but maintaining its PT renormalizability, will apparently satisfy the boundary conditions in QCD at the fundamental quark-gluon level: the confinement at large distances ($q^{2}\rightarrow 0$) and the AF at short distances ($q^{2}\rightarrow\infty$), as discussed in section 7.
+The particular or, equivalently, the PT/conventional QCD solution with a gauge symmetry preservation cannot satisfy the boundary condition at large distances.
+It shows the agreement with the AF regime at short distances, but cannot explain it, see discussion in section 7 as well.
+The perspectives for further work are as follows: to formulate the NP multiplicative (MP) renormalization programs for the regularized full gluon propagator (5.34).
+This is not a simple task, since such kind of a programs have to be different for the massive and severely singular gluon field configurations, briefly discussed in section 7.
 For this purpose it is necessary to take into account the transcendental and NL structure of eq.
@@
 To ensure a unitary of $S$-matrix elements by these NP MP renormalization programs should be requested as well.
+In general, our results open a new wide research areas in modern physics, leading to the multiple theoretical and phenomenological applications and comparing them with the results of the conventional QCD.
 To understand a mass dynamical generation at the quark-gluon level is the first necessary step to understand the existence of the physical mass spectrum at the hadronic level in QCD as a theory of strong interactions.
-The confinement first and only after the PCAC phase transitions in QCD have to be understood 2 ; 23 .
+The confinement first and only after the PCAC phase transitions in QCD have to be understood [2, 23].
 QCD is a self-consistent quantum gauge field theory.
-It does not need any extra degrees of freedom (such as Higgs fields 40 ) in order to generate a mass at the fundamental quark-gluon level.
+It does not need any extra degrees of freedom (such as Higgs fields [41]) to generate a mass at the fundamental quark-gluon level.
 Concluding, a few remarks are in order.
-The wide-spread opinion that the gauge symmetry cannot be broken/violated is misleading.
-The Elitzur’s theorem 41 has been proven in lattice field theory.
-It is supposed to be valid for the continuous theories as well, but if yes than only for the abelian systems, like QED.
-To the non-abelian systems, such as the YM fields, this theorem cannot be applied.
-In this paper we have just proven that the gauge symmetry of the YM theory is broken in its vacuum.
+The Elitzur’s theorem [42] has been proven in lattice field theory for the abelian system like QED (see appendix D as well).
+To the non-abelian systems like YM fields, it also can be applied, but only within its particular solution, as discussed above.
+At the same time, this theorem cannot be applied to the YM fields in the frame work of its exact solution for the generalized gauge-fixing function in the full gluon propagator.
+In this paper we have just shown that the gauge symmetry of the continuos YM theory can be broken in its ground state.
+Behind this effect is the mass squared parameter, which is nothing else but the renormalized tadpole term.
+It survives the NP renormalization but does not affect the PT renormalizability, as underlined just above.
 The relevant existing knowledge on a mass dynamical generation and spontaneously symmetry breaking subjects has been taken into account in the list of publications, presented in our paper (and see references in each of the cited items there).
+Also, let us note that contrary to the ghost term, the QCD Lagrangian’s gauge-fixing term is not explicitly present in the full gluon self-energy, that is why it is left out of our attention.
+It would be interesting to establish the character of its possible connection (if any) to the NL generalized gauge-fixing function approach.
+Apparently, this might be done only after the corresponding NP MP renormalization programs will be performed, not before, in our opinion.
+9 Summary
+Briefly, we can summarize our results as follows:
+The exact derivation of the general constraint (3.15) on any solution to QCD is present.
+This has been done in the mathematically understandable way for a broad readership of researches, since exploiting only the tensor algebra derivations rigorous rules, widely used in the theoretical physics, as pointed out above.
+It has the two possible exact solutions, in principle, independent from each other.
+The general solution to QCD is the solution with the gauge symmetry broken in its ground state.
+The dynamical source of this groundbreaking effect is the tadpole term.
+The general QCD remains the PT renormalizable, nevertheless.
+It will satisfy the boundary conditions at the fundamental quark-gluon level, as discussed in section 7.
+The particular or, equivalently, the PT/conventional solution to QCD is the solution with unbroken gauge symmetry in its ground state.
+It is the PT renormalizable as well, but not satisfying the boundary condition at large distances, as discussed in section 7.
+Its gauge symmetry hides the true role of the tadpole term in the QCD vacuum.
 Acknowledgments
@@
 Kluge, M.
-Vasúth for useful suggestions, remarks, discussions and help.
+Vasúth, P.
+Ván, G.
+Somogyi, S.
+Van Thurenhout for useful suggestions, remarks, discussions and help.
 The work was supported by the Hungarian National Research, Development and Innovation Office (NKFIH) under the contract numbers OTKA K135515 and NKFIH 2019-2.1.11-TÉT-2019-00078, 2019-2.1.11-TÉT-2019-00050, the Wigner Scientific Computing Laboratory.
 Appendix A The full gluon propagator up to one loop
-Figure 2: The SD equation for the full gluon propagator up to one loop as present in 6 . It is useful to compare it with Fig. 1 and its description in the text after it.
+Figure 2: The SD equation for the full gluon propagator up to one loop as present in [6]. It is useful to compare it with Fig. 1 and its description in the text after it.
 In order to explicitly compare the tadpole term (2.4) with the other tadpole-type terms let us consider the iteration of the gluon SD eq.
@@
 A.2.
-This iteration has been exactly calculated in 6 by using the DRM, while rightfully now ignoring the tadpole term itself and all the other tadpole-likes ones, in accordance with the relations (5.5)-(5.6).
-It is worth emphasizing that in 6 this iteration has not been called the PT expansion up to $O(g^{2})$, though looks like it, but expansion up to $O(\hbar)$.
+This iteration has been exactly calculated in [6] by using the DRM, while rightfully now ignoring the tadpole term itself and all the other tadpole-likes ones, in accordance with the relations (6.43).
+It is worth emphasizing that in [6] this iteration has not been called the PT expansion up to $O(g^{2})$, though looks like it, but expansion up to $O(\hbar)$.
 It is well-known that in QCD the expansion in powers of the coupling constant does not make any sense because it is strong.
-The PT expansion in QCD makes sense only in the AF 37 ; 38 regime when it becomes weak.
+The PT expansion in QCD makes sense only in the AF [37, 38] regime when it becomes weak.
 In order to perform analytical derivations below, we will take the interaction vertices as the point-like ones.
@@
 Also we will omit below all the colour group factors and other overall numerical numbers, since all of them are unimportant here.
-However, it is instructive first to show explicitly how works our subtraction scheme, using only the third diagram in the second line of Fig.
-A.2, for simplicity.
-From 6 by replacement $\mu\rightarrow\rho$ and $\nu\rightarrow\sigma$ and fixing $\omega=2$ and $\delta_{\rho\rho}=4$ in eq.
+However, it is instructive first to show explicitly how works our subtraction scheme, using only the third diagram in the second line of Fig.A.2, for simplicity.
+From [6] by replacement $\mu\rightarrow\rho$ and $\nu\rightarrow\sigma$ and fixing $\omega=2$ and $\delta_{\rho\rho}=4$ in eq.
 (6.22) there, one finally obtains an analytical expression for this term as follows:
-$\Pi^{(3)}_{\rho\sigma}(q)\sim\int d^{4}l\ {N_{\rho\sigma}(l,q)\over l^{2}(l+q)^{2}},$ (A.1)
+$\Pi^{(3)}_{\rho\sigma}(q)\sim\int d^{4}l\ {N_{\rho\sigma}(l,q)\over l^{2}(l+q)^{2}},$ (1.49)
 where, apart from below, we put the gluon loop propagators as free ones in the t’ Hooft-Feynman gauge as well, for the sake of transparency of our derivations for this diagram.
 The tensor function $N_{\rho\sigma}(l,q)$ is
-$N_{\rho\sigma}(l,q)=10l_{\rho}l_{\sigma}+5(l_{\rho}q_{\sigma}+q_{\rho}l_{\sigma})-2q_{\rho}q_{\sigma}+(5q^{2}+2(lq)+2l^{2})\delta_{\rho\sigma}.$ (A.2)
+$N_{\rho\sigma}(l,q)$ $=$ $10l_{\rho}l_{\sigma}+5(l_{\rho}q_{\sigma}+q_{\rho}l_{\sigma})-2q_{\rho}q_{\sigma}$ (1.50)
+$+$ $(5q^{2}+2(lq)+2l^{2})\delta_{\rho\sigma}.$
 Putting in eqs. (A.1) and (A.2) $q=0$, one gets
-$\Pi^{(3)}_{\rho\sigma}(0)\sim\int d^{4}l\ {N_{\rho\sigma}(l,0)\over(l^{2})^{2}}=\int d^{4}l\ {(10l_{\rho}l_{\sigma}+2l^{2}\delta_{\rho\sigma})\over(l^{2})^{2}},$ (A.3)
+$\Pi^{(3)}_{\rho\sigma}(0)\sim\int d^{4}l\ {N_{\rho\sigma}(l,0)\over(l^{2})^{2}}=\int d^{4}l\ {(10l_{\rho}l_{\sigma}+2l^{2}\delta_{\rho\sigma})\over(l^{2})^{2}},$ (1.51)
 and, obviously, this integral is quadratically divergent at the upper limit, but is assumed to be regularized from above and below from the very beginning, as emphasized throughout this work.
-Let us now compose the subtracted counterpart of the initial integral (A.1), in accordance with the subtractions (4.1), as follows:
-$\Pi^{(3s)}_{\rho\sigma}(q)=\Pi^{(3)}_{\rho\sigma}(q)-\Pi^{(3)}_{\rho\sigma}(0)\sim\int d^{4}l\ {N^{s}_{\rho\sigma}(l,q)\over(l^{2})^{2}(l+q)^{2}},$ (A.4)
+Let us now compose the subtracted counterpart of the initial integral (A.1), in accordance with the subtraction (4.18), as follows:
+$\Pi^{(3s)}_{\rho\sigma}(q)=\Pi^{(3)}_{\rho\sigma}(q)-\Pi^{(3)}_{\rho\sigma}(0)\sim\int d^{4}l\ {N^{s}_{\rho\sigma}(l,q)\over(l^{2})^{2}(l+q)^{2}},$ (1.52)
 where
-$N^{s}_{\rho\sigma}(l,q)$ $=$ $5l^{2}(l_{\rho}q_{\sigma}+q_{\rho}l_{\sigma})-20(ql)l_{\rho}l_{\sigma}+l^{2}(q^{2}-2(lq))\delta_{\rho\sigma}$ (A.5)
--- $2l^{2}q_{\rho}q_{\sigma}-10q^{2}l_{\rho}l_{\sigma}.$
-Because of the two last terms and term $l^{2}q^{2}\delta_{\rho\sigma}$ in this equation, the subtracted integral (A.4) is only logarithmically divergent at the upper limit.
-So that the integral (A.3) is one of the tadpole-like terms which should be disregarded on the general basis due to the relations (6.6)-(6.7) within our approach.
-The subtracted integral (A.4) finally will contribute to the invariant function which appear in the full gluon propagator (7.5).
+$N^{s}_{\rho\sigma}(l,q)$ $=$ $5l^{2}(l_{\rho}q_{\sigma}+q_{\rho}l_{\sigma})-20(ql)l_{\rho}l_{\sigma}-2l^{2}q_{\rho}q_{\sigma}$ (1.53)
+$+$ $l^{2}(q^{2}-2(lq))\delta_{\rho\sigma}-10q^{2}l_{\rho}l_{\sigma}.$
+[…]
 The tadpole term up to one skeleton loop, shown first in the third line of Fig. A.2, analytically can be written down as follows:
-$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lT^{0}_{\rho\sigma\alpha\beta}D_{\alpha\beta}(l)=\delta_{\rho\sigma}\Delta^{2}_{t}(D).$ (A.6)
+$\Pi_{\rho\sigma}^{t}(D)\sim\int d^{4}lT^{0}_{\rho\sigma\alpha\beta}D_{\alpha\beta}(l)=\delta_{\rho\sigma}\Delta^{2}_{t}(D),$ (1.54)
+see eq.
+(2.4) as well.
 Being the QD but already regularized constant, it is not connected to any external gluon momentum.
 In this respect it is different from all other tadpole-like terms which appear as a result of the subtraction scheme or in the NL iteration procedure.
-Also it directly generate the mass squared parameter, and thus explicitly violets the gauge symmetry of the QCD Lagrangian in its ground state 6 .
-We have shown that in the PT QCD it should be neglected on the general basis as well as all other tadpole-type terms.
-But in the QCD it should remain intact, while all other tadpole-types terms have to be removed from the theory, as it has been proven in section 6.
+Also it directly generate the mass squared parameter, and thus explicitly violets the gauge symmetry of the QCD Lagrangian in its ground state [6].
+We have shown that in the PT QCD or,equivalently, particular QCD (described in section 6) it was a prescription to neglect it, but in the full QCD or, equivalently, NP QCD (investigated in sections 4 and 5) it should remain intact.
+At the time, all other tadpole-type terms have to be disregarded on the exact mathematical basis in both theories.
 Let us now investigate the tadpole-like term with the gluon skeleton loop, shown second in the third line of Fig. A.2. Its analytical expression is
-$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l\ T^{0}_{\nu\zeta\mu}(l,p)D_{\nu\zeta}(l),$ (A.7)
+$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l\ T^{0}_{\nu\zeta\mu}(l,p)D_{\nu\zeta}(l),$ (1.55)
 where the gluon momentum $p$ is, in fact, zero, i.e., $p=0$ but it is convenient to go to this limit at the final step only.
@@
 Let us remind that such kind of the skeleton loop integrals are assumed to be regularized from above and below.
-Using now the Euclidean space Feynman rules for the corresponding vertices present in 6 , one arrives at
-$T^{0}_{\rho\sigma\mu^{\prime}}(q,p)$ $=$ $-(2q+p)_{\mu^{\prime}}\delta_{\rho\sigma}+(q-p)_{\sigma}\delta_{\rho\mu^{\prime}}+(2p+q)_{\rho}\delta_{\sigma\mu^{\prime}},$
-$T^{0}_{\nu\zeta\mu}(l,p)$ $=$ $2(l+p)_{\mu}\delta_{\nu\zeta}-l_{\zeta}\delta_{\nu\mu}-(2p+l)_{\nu}\delta_{\zeta\mu}.$ (A.8)
+Using now the Euclidean space Feynman rules for the corresponding vertices present in [6], one arrives at
+$T^{0}_{\rho\sigma\mu^{\prime}}(q,p)=-(2q+p)_{\mu^{\prime}}\delta_{\rho\sigma}+(q-p)_{\sigma}\delta_{\rho\mu^{\prime}}+(2p+q)_{\rho}\delta_{\sigma\mu^{\prime}}$ (1.56)
+and
+$T^{0}_{\nu\zeta\mu}(l,p)=2(l+p)_{\mu}\delta_{\nu\zeta}-l_{\zeta}\delta_{\nu\mu}-(2p+l)_{\nu}\delta_{\zeta\mu}.$ (1.57)
 Substituting these expressions into the eq. (A.7), one finally obtains
-$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p){2\over p^{2}}\int d^{4}l{d(l^{2})\over l^{2}}[3l_{\mu}+p_{\mu}].$ (A.9)
-By the symmetry integration ($d^{4}l=l^{3}dl=(1/2)l^{2}dl^{2}$ and all the overall numerical numbers due to the integration over angular variables in four-dimensional Euclidean space 6 omitted) the first loop integral is zero, while the second one leads to
-$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p)p_{\mu}{2\over p^{2}}\times\tilde{\Delta}^{2}_{g}(d),$ (A.10)
+$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p){2\over p^{2}}\int d^{4}l{d(l^{2})\over l^{2}}[3l_{\mu}+p_{\mu}].$ (1.58)
+By the symmetry integration ($d^{4}l=l^{3}dl=(1/2)l^{2}dl^{2}$ and all the overall numerical numbers due to the integration over angular variables in four-dimensional Euclidean space [6] omitted) the first loop integral is zero, while the second one leads to
+$\Pi^{(4)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu}(q,p)p_{\mu}{2\over p^{2}}\times\tilde{\Delta}^{2}_{g}(d),$ (1.59)
 where $\tilde{\Delta}^{2}_{g}(d)=\int d^{4}l(d(l^{2})/l^{2})$ is the QD but regularized skeleton loop integral, having the dimension of a mass squared.
-It is one of the constants present in the relations (6.7)-(6.8).
-It should be discarded on the general basis within our approach, i.e., put formally zero, so that $\tilde{\Delta}^{2}_{g}(d)=0$, even before the final $p^{2}=-M^{2}\rightarrow 0$ limit.
+It is one of the constants present in the relations (4.26).
+It should be discarded on the general mathematical basis within our approach, i.e., put zero, so that $\tilde{\Delta}^{2}_{g}(d)=0$, even before the final $p^{2}=-M^{2}\rightarrow 0$ limit.
 Let us note that the cancellation of the pole $1/p^{2}=-1/M^{2}$ in eq.
-(B.10) can be always achieved by going to the dimensionless variable $x=l/M$ in the skeleton loop integral (A.9), i.e., this pole is not a problem here.
+(A.11) can be always achieved by going to the dimensionless variable $x=l/M$ in the skeleton loop integral (A.10), i.e., this pole is not a problem here.
 Thus, $\Pi^{(4)}_{\rho\sigma}(q,p)=0$.
 Let us finally investigate the tadpole-like term with the ghost skeleton loop, shown third in the third line of Fig. A.2. In analogy with eq. (A.7) its analytical expression is
-$\Pi^{(5)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l{l_{\mu}\over l^{2}}d_{gh}(l^{2}),$ (A.11)
+$\Pi^{(5)}_{\rho\sigma}(q,p)\sim T^{0}_{\rho\sigma\mu^{\prime}}(q,p)D^{0}_{\mu\mu^{\prime}}(p)\int d^{4}l{l_{\mu}\over l^{2}}d_{gh}(l^{2}),$ (1.60)
 where $d_{gh}(l^{2})$ is the invariant function for the full ghost propagator $G(l)$.
 The vertex $T^{0}_{\rho\sigma\mu^{\prime}}(q,p)$ is explicitly shown in the relations (A.8), and the remaining free gluon propagator is also the same.
-The altered rule 6 for the gluon-ghost vertex gives $(1/2)[(l_{\mu}+l{\mu})=l_{\mu}$ and the gluon momentum $p$ is zero again at the final stage.
+The altered rule [6] for the gluon-ghost vertex gives $(1/2)[(l_{\mu}+l{\mu})=l_{\mu}$ and the gluon momentum $p$ is zero again at the final stage.
 Since the nominator of this skeleton loop integral contains the loop variable linearly, then by the symmetry integration it is simply zero from the very beginning, i.e., $\Pi^{(5)}_{\rho\sigma}(q,p)=0$.
@@
 This was not a prescription, but being an exact mathematical result within our formalism.
+Appendix B The origin of the conventional QCD
+In order to find the general solution to QCD, we present the full gluon self-energy (2.2) as a sum of the three independent terms, i.e., we distinguish between the gluon part and the tadpole term, though their sum forms the YM part of QCD, since the both are determined by the different self-interaction of multiple massless gluon modes.
+However, not doing so, one needs to present the full gluon self-energy in terms of the quark and YM compositions, i.e., one has to operate with
+$\Pi_{\rho\sigma}(q;D)=\Pi^{q}_{\rho\sigma}(q)+\Pi^{YM}_{\rho\sigma}(q;D),$ (2.61)
+where
+$\Pi^{YM}_{\rho\sigma}(q;D)=\Pi^{g}_{\rho\sigma}(q;D)+\delta_{\rho\sigma}\Delta^{2}_{t}(D).$ (2.62)
+The relation (3.13) in this case becomes
+$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{YM}_{\rho\sigma}(q;D)]={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (2.63)
+As we already know, the quark contribution is always satisfied, i.e., $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=0$, which leads to $\Delta^{2}_{q}=0$. Thus, the relation (B.3) is reduced to
+$q_{\rho}q_{\sigma}\Pi^{YM}_{\rho\sigma}(q;D)={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (2.64)
+In complete analogy with the subtraction (4.19), let us now introduce the subtraction for the YM part as follows:
+$\Pi^{s(YM)}_{\rho\sigma}(q;D)$ $=$ $\Pi^{YM}_{\rho\sigma}(q;D)-\Pi^{YM}_{\rho\sigma}(0;D)$ (2.65)
+$=$ $\Pi^{YM}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{YM}(D)$
+where $\Pi^{s(YM)}_{\rho\sigma}(0;D)=0$, by definition, and
+$\Delta^{2}_{YM}(D)=\Delta^{2}_{g}(D)+\Delta^{2}_{t}(D),$ (2.66)
+as it follows from eq. (B.2).
+Instead of the tensor decompositions (4.22), for the YM parts one gets
+$\Pi^{s(YM)}_{\rho\sigma}(q)=T_{\rho\sigma}(q)q^{2}\Pi^{s(YM)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(YM)}_{l}(q^{2}),$ (2.67)
+and
+$\Pi^{YM}_{\rho\sigma}(q)=T_{\rho\sigma}(q)q^{2}\Pi^{YM}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{YM}_{l}(q^{2}).$ (2.68)
+In all the quantities above and below the dependence on $D$ is omitted, for simplicity, and will be restored when necessary.
+Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary.
+However, all the invariant functions $\Pi^{s(YM)}_{t}(q^{2})$, $\Pi^{s(YM)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the subtraction (B.5).
+Substituting the relations (B.7)-(B.8) into the relations (B.4)-(B.5), by omitting the tedious tensor algebra derivations, one finally obtains
+$\Pi^{s(YM)}_{t}(q^{2})$ $=$ $\Pi^{YM}_{t}(q^{2})-{\Delta^{2}_{YM}(D)\over q^{2}},$
+$\Pi^{s(YM)}_{l}(q^{2})$ $=$ $-{(\xi_{0}-\xi)\over\xi\xi_{0}}+{\Delta^{2}_{YM}(D)\over q^{2}}.$ (2.69)
+However, the invariant function $\Pi^{s(YM)}_{l}(q^{2})$ cannot have the pole-type singularities, by definition, as underlined just above.
+So that one has to put the QD but regularized constant $\Delta^{2}_{YM}(D)=0$ on this mathematical ground.
+Taking into account that the quark constant $\Delta^{2}_{q}=0$ as well, one concludes that all the QD but regularized constants should be removed from theory, i.e., to be put zero as follows:
+$\Delta^{2}_{q}=\Delta^{2}_{g}(D)=\Delta^{2}_{t}(D)=0,$ (2.70)
+because of the relation (B.6) and all of them are independent from each other.
+If $\Delta^{2}_{t}(D)=0$ then from the relation (B.2) it follows that $\Pi^{YM}_{\rho\sigma}(q;D)=\Pi^{g}_{\rho\sigma}(q;D)$.
+But from the satisfied transverse relation $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ and previous equality it follows that $q_{\rho}q_{\sigma}\Pi^{YM}_{\rho\sigma}(q;D)=0$ as well.
+Then from the relation (B.4) one gets $\xi_{0}=\xi$.
+The relations (B.10) and this equality determines the particular, or, equivalently, conventional QCD, investigated in section 6.
+Here it has been obtained by treating all the QD constants on the same footing, completely ignoring their principle difference from the physical point of view, which has been emphasized throughout this work.
+This happened because in the previous derivations we deliberately have ignored that the subtraction (B.5), on account of the relation (B.2), in fact, looks like
+$\Pi^{s(YM)}_{\rho\sigma}(q;D)=\Pi^{g}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{g}(D),$ (2.71)
+i.e., $\Pi^{s(YM)}_{\rho\sigma}(q;D)=\Pi^{s(g)}_{\rho\sigma}(q;D)$, as it follows from eq.
+(4.18).
+Let us underline that the disappearance of the tadpole term is not a specific feature of our subtraction scheme.
+This always will happen in any subtraction scheme, since the constant minus the same constant is always zero, as pointed out in section 4.
+That is why explicitly present in the initial expressions it will remain in the final ones, compare eq.
+(2.5) with eq.
+(5.27), and thus with eq.
+(5.34).
+Substituting further the tensor decomposition (B.7) and second of the decompositions (4.22) into the subtraction (B.11), and again omitting some tedious tensor algebra derivations, one obtains
+$\Pi^{s(YM)}_{t}(q^{2})$ $=$ $\Pi^{g}_{t}(q^{2})-{\Delta^{2}_{g}(D)\over q^{2}},$
+$\Pi^{s(YM)}_{l}(q^{2})$ $=$ $\Pi^{g}_{l}(q^{2})+{\Delta^{2}_{g}(D)\over q^{2}}.$ (2.72)
+The general relation (B.4), on account of the first line of the subtraction (B.5), becomes
+$q_{\rho}q_{\sigma}[\Pi^{s(YM)}_{\rho\sigma}(q;D)+\Pi^{YM}_{\rho\sigma}(0;D)]={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (2.73)
+Substituting now the decomposition (B.7) and the relation $\Pi^{YM}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{YM}(D)$, which follows from the relations (B.5), into the previous equation, and omitting again tensor algebra derivations, one obtains
+$\Pi^{s(YM)}_{l}(q^{2})=-{(\xi_{0}-\xi)\over\xi\xi_{0}}+{(\Delta^{2}_{g}(D)+\Delta^{2}_{t}(D))\over q^{2}}.$ (2.74)
+Equating now this relation to the second of the relations (B.12), and doing simple algebra, one finally arrives at
+${(\xi_{0}-\xi)\over\xi\xi_{0}}q^{2}=\Delta^{2}_{t}(D),$ (2.75)
+which is nothing else but the general constraint (3.15).
+Note that in deriving this expression, we again use that from the satisfied transverse relation for the gluon part $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ it follows that $\Pi^{g}_{l}(q^{2})=0$.
+As in the case of the conventional QCD, the properties of the gluon degrees of freedom have been used at the last stage.
+Even introducing the YM contribution to the full gluon self-energy in eq.
+(B.1), which hides the tadpole term, but correctly treated its subtraction (B.5) by the subtraction (B.11), one again arrives at the exact constraint (B.15), nevertheless.
+Treated it literally (i.e., by eq.
+(B.5) only) leads to the conventional QCD.
+In this case the significant piece of the information on the true dynamical and gauge structures of the QCD vacuum is losing.
+In other words, the conventional QCD hides the true role of the tadpole term.
+Contrary, in general QCD, based on the exact solution of eq.
+(B.15), this information is not losing and the true role of the tadpole has been revealed.
+Concluding, let us emphasize that the exact constraint (3.15) has been obtained from the very beginning before introducing any proper subtraction scheme.
+This underlines its general and unique character.
+At the same time, the clarifying algebraic derivations performed here, leading to the relations (B.10), heavily depend on the not accurate (i.e., literally) treatment of the subtraction (B.5), as explained just above.
+Appendix C Renormalized ST identity
+Substituting eq. (3.16) into the ST identity (3.7), one obtains
+$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi=i{\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)}.$ (3.76)
+Let us introduce the non-perturbative (NP) renormalization constant for the tadpole term as follows: $\Delta^{2}_{t}(D)=Z_{\Delta}\Delta^{2}$, where $\Delta^{2}$ is finite and positive, by definition, while the NP renormalization constant $Z_{\Delta}$ depends on all the unphysical parameters, such as regulating parameters, mentioned above, etc.
+In principle, the string tension in the linear rising potential between heavy quarks has to be somehow related to the finite $\Delta^{2}$, since it is connected to the transverse projection operator structure, as underlined above.
+It is well-known that the ansatz $D_{\mu\nu}(q)\sim\delta_{\mu\nu}\Delta^{2}/(q^{2})^{2}$ just leads to a such potential.
+After doing some algebra, the ST identity for the renormalized full gluon propagator becomes
+$q_{\mu}q_{\nu}D^{R}_{\mu\nu}(q)=i\xi^{R}=i{\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}},$ (3.77)
+where the renormalized gauge-fixing parameters are defined as follows: $\tilde{\xi}_{0}=Z_{\Delta}\xi_{0},\ \xi^{R}=Z_{\Delta}\xi$ and $D^{R}_{\mu\nu}(q)=Z_{\Delta}D_{\mu\nu}(q)$, i.e., the ST identity (C.1) is expressed in terms of the finite quantities only.
+The correct PT limit is maintained, of course.
+Let us now demonstrate the interesting feature of the generalized gauge (C.2). It provides the self-consistency condition for the gauge choice in QCD as follows:
+${\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}}={aq^{2}\over q^{2}+a\Delta^{2}},$ (3.78)
+[…]
+${\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}}={q^{2}\over\Delta^{2}},$ (3.79)
+which is only satisfied at $q^{2}=0$, i.e., there is no any condition for the gauge-fixing parameter.
+The formulated self-consistency condition (C.3) points out on the inconsistency of the canonical gauge $\tilde{\xi}_{0}=\infty$ in QCD.
+Briefly, if everything is expressed in the terms of the finite quantities then the formal canonical gauge is forbidden to use in QCD, according to our approach to this theory.
+Let us now apply the above-developed formalism to the renormalized massive gluon fields.
+For this purpose, it is instructive to re-write eq.
+(7.46) for the free massive gluon propagator in the terms of the renormalized quantities.
+But before it is convenient to introduce the following notations: $\tilde{\xi}_{0}=\tilde{\lambda}^{-1}$ and $M^{2}=M^{2}_{g}$, and the corresponding renormalization constant $Z_{\Delta}=1$ in this case.
+Then eq.
+(7.46) will look like
+$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2}_{g})}\left[\delta_{\mu\nu}-{(1-\tilde{\lambda}^{-1})q_{\mu}q_{\nu}\over(q^{2}-\tilde{\lambda}^{-1}M^{2}_{g})}\right].$ (3.80)
+From this observation it follows that our approach makes it possible for gluon to acquire mass dynamically without introducing into the theory some extra degrees of freedom such as Higgs fields [41], mentioned above, and $M^{2}_{g}$ can be treated as exactly defined gluon pole mass.
+If one puts the expression (C.5) on the mass-shell $q^{2}=M^{2}_{g}$, then one arrives at
+$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2}_{g})}\left[\delta_{\mu\nu}-{q_{\mu}q_{\nu}\over M^{2}_{g}}\right],$ (3.81)
+which leads to the non-renormalizable behaviour in the PT $q^{2}\rightarrow\infty$ limit (i.e., it is constant and not $\sim 1/q^{2}$ as it is required for the renormalizable behaviour).
+However, the same expression (C.6) will be obtained if one expresses (C.5) in the canonical gauge ${\tilde{\lambda}}^{-1}=\infty$.
+As explained above, this gauge is not to be used in QCD.
+This means that within the generalized gauge-fixing function approach the free massive gluons cannot be put on the mass-shell.
+The finite values for the gauge-fixing parameter ${\tilde{\lambda}}^{-1}$ are only permitted in order to guarantee the theory to be renormalizable.The expression (C.5) has the correct PT limit, being off-mass-shell object at any finite ${\tilde{\lambda}}^{-1}$.
+So that, the canonical gauge is rather formal than real within the generalized gauge-fixing function approach to QCD.
+The canonical gauge and its connection to the structure of the full gluon propagator at the mass-shell will be investigated in our future work, where the NP renormalization theory of the massive gluon fields will be present in detail.
+Appendix D Preservation of the gauge symmetry in the QED ground state
+The propagation of the photon in the QED vacuum is formally described by eq.
+(2.1).
+Contrary to QCD (compare with eq.
+(2.2), the difference is that there is only one contribution to the full photon self-energy comes from the vacuum polarization tensor.
+Its skeleton term is similar present in Fig.
+1 by the first diagram.
+Analytically it looks like
+$\Pi^{ph}_{\rho\sigma}(q)\sim\int d^{4}l\gamma_{\rho}S(l-q)\Gamma_{\sigma}(l,q)S(l),$ (4.82)
+where all the overall numerical factors are omitted, for simplicity.
+Here $S$ and $G$ represent the full electron propagator and the full electron-photon vertex, respectively.
+The full photon and its free counterpart are analytically present similar to eq.
+(3.8) and eq.
+(3.9), respectively, as well as their ST identities by relations (3.7) and (3.10).
+Similar to the relation (3.11) for the transverse relation in this case, one obtains
+$q_{\rho}q_{\sigma}\Pi^{ph}_{\rho\sigma}(q;D)={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (4.83)
+It is well-known that because of the electric charge conversation, the transverse relation for the vacuum polarization tensor in QED is always satisfied, i.e., it is zero.
+Then from the previous equation immediately follows
+$q_{\rho}q_{\sigma}\Pi^{ph}_{\rho\sigma}(q;D)=0,\quad\rightarrow\quad\xi_{0}=\xi.$ (4.84)
+Let us now introduce the corresponding subtraction scheme similar to the relation (4.19) and their independent tensor decompositions (4.22), namely
+$\Pi^{s(ph)}_{\rho\sigma}(q)=\Pi^{ph}_{\rho\sigma}(q)-\Pi^{ph}_{\rho\sigma}(0)=\Pi^{ph}_{\rho\sigma}(q)-\delta_{\rho\sigma}\Delta^{2}_{ph},$ (4.85)
+where $\Pi^{s(ph)}_{\rho\sigma}(0)=0$, by definition, and
+$\Pi^{s(ph)}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(ph)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(ph)}_{l}(q^{2}),$
+$\Pi^{ph}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{ph}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{ph}_{l}(q^{2}),$ (4.86)
+respectively.
+Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary.
+However, the both invariant functions $\Pi^{s(ph)}_{t}(q^{2})$, $\Pi^{s(ph)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the initial subtraction (D.4).
+Substituting all these decompositions (D.5) into the subtraction (D.4), and doing some tensor algebra derivations, one finally obtains
+$\Pi^{s(ph)}_{t}(q^{2})$ $=$ $\Pi^{ph}_{t}(q^{2})-{\Delta^{2}_{ph}\over q^{2}},$
+$\Pi^{s(ph)}_{l}(q^{2})$ $=$ $\Pi^{ph}_{l}(q^{2})+{\Delta^{2}_{ph}\over q^{2}}.$ (4.87)
+From the first of the relations (D.3) and the second of the relations (D.5) it follows that $\Pi^{ph}_{l}(q^{2})=0$. Then the second of the relations (D.6) will be reduced to
+$\Pi^{s(ph)}_{l}(q^{2})={\Delta^{2}_{ph}\over q^{2}},$ (4.88)
+which is impossible since the corresponding invariant function can not have the pole-type singularities, by definition, as explained above.
+One has to put the QD but regularized constant to zero on this mathematical basis, i.e., $\Delta^{2}_{ph}=0$, and thus $\Pi^{s(ph)}_{l}(q^{2})=0$ as well.
+As we already know, the both relations
+$\xi=\xi_{0},\quad\Delta^{2}_{ph}=0$ (4.89)
+are necessary conditions for the coincidence of the gauge symmetry of the QED Lagrangian with the gauge symmetry of its ground state (like it occurs for the conventional QCD).
+This is our proof of the Elitzur theorem [42] for the continuos abelian fields within our approach, based on the formalism of the dynamical equations of motion for the gauge particle Euclidean Green’s functions.

### v4

@@
-The gauge symmetry violation in the QCD ground state
+Confining QCD theory and Mass Gap
 Abstract
-The purpose of this work is to report on convincing progress made in the investigation of the true dynamical and gauge structures of the QCD ground state (vacuum).
-It also sheds new significant insights into a mass squared dynamical generation in it.
-For this purpose, we use only the tensor algebra derivation rigorous rules in dealing with the corresponding skeleton loop integrals, contributing into the full gluon self-energy.
 The novel constraint on any solution to QCD has been exactly and uniquely derived in the framework of the the Slavnov-Taylor identities for the gauge particles propagators.
-Its general solution makes it possible to prove that the gauge symmetry of the QCD Lagrangian is broken in its ground state.
-The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared.
-The exact proof has been given that it cannot be disregarded from the theory and its ground state by any means.
-The gauge symmetry of the QCD Lagrangian forbids such kind of terms, but the perturbative renormalizability of the theory will not be affected.
-The particular solution of the novel constraint describes the gluon field configurations with unbroken gauge symmetry in the QCD vacuum.
-The solution with broken gauge symmetry dominates over the solution with unbroken one at large distances, while being in agreement with it at short distances.
+It has the two independent solutions.
+1).
+The first one leads to the general QCD, possessing the confinement phase transition.
+2).The second one leads to the particular QCD, or, equivalently, the conventional QCD, does not possessing it.
+The dynamical source of this effect has to be identified with the constant tadpole term, having the dimension of a mass squared, contributing to the full gluon self-energy.
+Its renormalized version is conventionally called a mass gap.
+By formulating the proper subtraction scheme, we prove that it cannot be disregarded from the theory and its ground state by any means, but the perturbative renormalizability of QCD will not be affected.The general QCD forbids the free gluons to appear in the physical spectrum at large distances, and, at the time, has an asymptotic freedom behaviour at high energies.
+Finally, we describe how our advance results are connected to the Jaffe-Witten’s theorem.
 keywords:
-Theoretical particle physics, Quantum Chromodynamics, ground state, full gluon self-energy, gauge
+Theoretical particle physics, Quantum Chromodynamics, vacuum, full gluon self-energy, gauge, mass gap
 PACS:
@@
 The second one at the (macroscopic) hadronic level is – the PCAC (Partially Conserved Axial Currents) phase transition – in order to explain the soft pion physics [2, 23].
-However, this purpose remains a formidable task yet because of the multiple dynamical and topological complexities of low-energy particles (hadrons) physics, originated from QCD and its ground state (vacuum).
+However, this purpose remains a formidable task yet because of the multiple dynamical and topological complexities of low-energy particles (hadrons) physics, originated from the QCD and its ground state (vacuum).
 This happens because QCD as a gauge theory of the gluon-gluon and quark-gluon strong interactions suffers from a few important conceptual problems.
-Here we focus on the two of them, being in the close connection with each other.
-A. Whether the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is a symmetry of its ground state or not?
-B.
-The dynamical generation of a mass squared in the vacuum of QCD, since the gauge invariance of its Lagrangian forbids such kind of terms, for example like the gluon mass term $M^{2}_{g}A_{\mu}A_{\mu}$.
-The solutions of these fundamental problems is a key to the explanation of the first confinement phase transition in QCD.
+The conventional QCD, which has been investigated the last five decades, cannot explain the colour confinement problem, i.e., why the coloured objects– gluons and quarks– cannot appear at low-energies/large distances, which is an experimental fact.
+Secondly, it cannot explain the asymptotic freedom (AF) phenomenon, i.e., a scale breaking at high-energies/short distances [38, 39], which is also an experimental fact.
+To explain the both experimental facts at low- and high-energies it is necessary to answer a long-standing question how does a mass come out of a massless theory (the mass gap) [2]?
+even at the fundamental quark gluon level.
+For example, the linear rasing potential between heavy quarks requires the mass scale parameter – the string tension.
+Also, a scale breaking in AF regime needs a mass squared scale parameter.
+So that, the goal arises how to solve the above-mentioned problems, and thus to formulate the confining QCD theory, and, at same time, explaining AF phenomenon as well.
+It may come from the investigation of the true gauge and dynamical structures of the QCD vacuum in more detail.
 The properties and symmetries of the QCD Lagrangian, and thus including its Yang – Mills (YM) part, are well-known [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
@@
 2 Gluon SD equation
-The system of the SD equations of motions in QCD, including the above-mentioned gluon SD equation, in general, contains much more informations on the properties of the theory than its Lagrangian can provide at all [2].
-So it might be the place where a mas squared scale parameter, forbidden by the QCD Lagrangian gauge symmetry, could be generated.
+The system of the SD equations of motions and the corresponding Slavnov-Taylor (ST) identities in QCD, in general, contains much more informations on the properties of the theory than its Lagrangian can provide at all [2].
+So it might be the place where a mas squared scale parameter could be generated.
 Among them, the special role belongs to the equations, describing the propagation of the fundamental constituents in the QCD vacuum – gluons and quarks.
-The gluon SD equation has a rather complicated tensor structure because of its above-mentioned NL character.
+The above-mentioned gluon SD equation has a rather complicated tensor structure because of its NL character.
 Let us emphasize that all the known interactions in nature (electroweak, gravitational and strong) are described by the gauge theories.
-The common mathematical language for them is the tensor algebra, which is only one used throughout this paper.
-That is why our works will be easily understood by a broad spectrum of researches from the different areas of the theoretical physics.
+The common mathematical language for them is the tensor algebra, which is only one used throughout this paper.That is why our work will be easily understood by a broad spectrum of researches from the different areas of the theoretical physics.
 The gluon SD equation analytically looks like
@@
 In what follows the constant $\Delta^{2}_{t}(D)$ will be called as the tadpole term itself, since it deviates from the definition (2.4) by the tensor $g_{\rho\sigma}$ only.
+So that, the tadpole term equally contributes to the transverse and longitudinal components of the full gluon self-energy (2.2).
 It is instructive for further purpose to present the initial gluon SD eq. (2.1), on account of the expressions (2.2) and (2.4) as follows:
@@
 They will be assumed but not shown in all regularized expressions, for simplicity.
-The tadpole term $\Delta^{2}_{t}(D)$, defined in eq.
-(2.4), is nothing else but the QD skeleton loop integral, which is already regularized from below and above as well as all the other such kind of constants which may appear in the theory by any possible ways (the subtraction schemes, the iteration solutions, etc., see below).
+The constant tadpole term $\Delta^{2}_{t}(D)$, defined in eq.
+(2.4), is nothing else but the QD, i.e, UV divergent skeleton loop integral, which is already regularized from below and above, as well as all the other such kind of constants which may appear in the theory by any possible ways (the subtraction schemes, the iteration solutions, etc., see below).
 Being a mass squared regularized quantity, it is explicitly present in the QCD ground state, see eq.
 (2.5).
-However, it violets the gauge symmetry of the QCD Lagrangian.
 Then a natural question arises why is it present in the vacuum of QCD at all if it makes the theory to look non-renormalizable from the very beginning?
-The standard solution to this dilemma was to remove this and all other QD constants in order to make theory renormalizable in accordance with the QCD Lagrangian’s exact gauge symmetry.
-However, here we are going to solve this dilemma by asking the question is it possible to retain the tadpole term in the QCD vacuum, but without affecting the PT renormalization properties of the theory?
-This is important to ask because, unlike to other QD constants, the tadpole term is only one which dynamically generates a mass squared scale parameter, dominated by the point-like four-gluon vertex only, and thus does not depending on the external gluon momentum [6].
-Also, one of our primary aims will be to clarify its role in the dynamical and gauge structures of the QCD vacuum in more detail.
+Let us remind that the theory possessing the QD quantities is not the PT renormalizable one since the logarithmic-type divergences only can be accounted for the PT.
+The standard solution was to remove this and all other QD constants in any possible gauge-invariant way from the theory in order to make it renormalizable in the PT sense.
+However, here we are going to ask the question is it possible to retain the tadpole term in the QCD vacuum, but without affecting the PT renormalizability of the theory?
+This is important to ask because, unlike to other QD constants, the tadpole term is only one which dynamically generates a mass squared scale parameter, dominated by the point-like four-gluon vertex only, and thus does not depending on the external gluon momentum at all [6].
+Also, we will clarify its role in the dynamical and gauge structures of the QCD ground state in more detail.
 Concluding, let us remind that the whole gluon momentum range is $q^{2}\in[0,\infty)$.
@@
 3 Exact constraint on any solution to QCD
-It is well known that all the Slavnov-Taylor (ST) identities which appear in QCD [1] are important for the renormalizability of the theory, and thus they ”are exact constraints on any solution to QCD” [2].
-Among them the ST identities [2, 24, 35] for the gauge particle propagators play a dominant role, since they determine the gauge structure of the QCD ground state (vacuum).
+It is well known that all the ST identities which appear in QCD are important for the renormalizability of the theory, and thus they ”are exact constraints on any solution to QCD” [2].
+Among them the ST identities [2, 24, 35] for the gauge particle propagators play a dominant role, since they determine the gauge structure of the QCD vacuum.
 The ST identity for the full gluon propagator looks like
@@
 (3.8), then one obtains the free gluon propagator
-$D^{0}_{\mu\nu}(q)=i\left[T_{\mu\nu}(q)+\xi_{0}L_{\mu\nu}(q)\right](1/q^{2}),$ (3.9)
+$D^{0}_{\mu\nu}(q)=i\left[T_{\mu\nu}(q)+\xi_{0}L_{\mu\nu}(q)\right]{1\over q^{2}},$ (3.9)
 where $\xi_{0}$ is the corresponding gauge-fixing parameter. The general ST identity (3.7) will look like
 $q_{\mu}q_{\nu}D^{0}_{\mu\nu}(q)=i\xi_{0}.$ (3.10)
-Let us note in advance that from the very beginning the gauge-fixing parameter for the full gluon propagator $\xi$ is not to be identified with its free counterpart $\xi_{0}$.
+It is important to note in advance that from the very beginning the gauge-fixing parameter for the full gluon propagator $\xi$ is not to be identified with its free counterpart $\xi_{0}$.
 Since the gauge freedom in QCD is determined by $\xi_{0}$, there should exist a relation between them.
@@
 (2.3).
-At the same time, it is well-known that in QCD that just the Faddeev – Popov (FP) ghost contribution [43], $\Pi^{gh}_{\rho\sigma}(q)$ makes the transverse relation for this term satisfied (independently from the tadpole term (2.4), which itself has the transverse projection operator structure).
+At the same time, it is well-known that in QCD just the Faddeev – Popov (FP) ghost contribution [42] $\Pi^{gh}_{\rho\sigma}(q)$ makes the transverse relation for this term satisfied (independently from the tadpole term (2.4), which itself has the transverse projection operator structure).
 It is instructive to emphasize that each term in the sum (2.3) cannot be made transverse, only the whole sum.
@@
 $\xi=f(q^{2};\xi_{0})={\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)}.$ (3.16)
-We will call it as the generalized gauge-fixing function approach to QCD as well.
-It has been exactly (i.e., without making any kind of the simplifications) and uniquely defined in a such new manner for the first time.This expression demonstrates the NL dependence $\xi=f(q^{2};\xi_{0})$ on $\xi_{0}$.
+This relation has been exactly (i.e., without making any kind of the simplifications) and uniquely defined in a such new manner for the first time.This expression demonstrates the NL dependence $\xi=f(q^{2};\xi_{0})$ on $\xi_{0}$.
 The linear relation $\xi\rightarrow\xi_{0}$ will be recovered in the PT $q^{2}\rightarrow\infty$ limit, when the ratio $(\Delta^{2}_{t}(D)/q^{2})$ is to be suppressed in this regime at finite $\xi_{0}$.
-We have established the general boundary condition in QCD for the behaviour of the gauge-fixing function in the PT limit for the regularized full gluon propagator.
-Since the general solution (3.16) has been obtained within the ST identities formalism, we will call it as the general solution to QCD.
-Let us underline that to know it means to know the true gauge structure of the QCD vacuum and its theory at the fundamental quark-gluon level (for more detailed description see section 5 as well).
+We have established the general boundary condition in QCD for the behaviour of this relation in the PT limit for the regularized full gluon propagator.
+Since the exact solution (3.16) has been obtained within the ST identities general formalism, we will call it as the general solution to QCD and its vacuum, i.e., at the fundamental quark-gluon level (see section 5 for detailed description).
 3.2 Particular solution
@@
 […]
-but these equalities should be put by hand in this case, i.e., they are the prescriptions rather than the exact mathematical results.
-We will call it as the particular solution to QCD.
-Let us note in advance that these conditions determine the so-called conventional or, equivalently, conventional PT QCD (though the coupling constant remains always strong apart from the asymptotic free (AF) regime [2, 38, 39]).
-Precisely, this solution to QCD has been studied the last five decades.
-In this solution the gauge symmetries of the QCD Lagrangian and its ground state coincide, like in QED (see section 6 as well).
+but these equalities should be put by hand in this case.
+They are the prescriptions rather than the exact mathematical results.
+In what follows we will call it as conventional or, equivalently, the PT QCD as well (though the coupling constant remains always strong apart from the asymptotic free (AF) regime [2, 38, 39]).
+For its description in some more details see section 6.
 4 Proper subtraction scheme
-QCD being much more complicated quantum field gauge theory than QED, has the two satisfied transverse relations (3.14), and so requires much more careful investigation in QCD.
+QCD being much more complicated quantum field gauge theory than QED, has the two independent satisfied transverse relations (3.14), and so requires much more careful investigation in QCD.
 They have to be included in a self-consistent way into the proper subtraction scheme.
 Finally this will lead to the removal of the corresponding QD constants from the theory.
-How they fixed the gauge-fixing function approach to QCD in the most general and unique ways has been just described here.
+How they fixed the relations (3.16) and (3.17) in the most general and unique ways has been just described above.
 The first step in the renormalization program of any gauge theory is the removal of the quadratic UV divergences in order to make the corresponding theory renormalizable in the PT sense.
@@
 Just this decreases the quadratic UV divergences of the corresponding skeleton loop integrals to a logarithmic ones, as it has been described in detail in this section.
-5 General solution to QCD
+5 General QCD
 Substituting the sum of the terms (4.27) into the eq.
@@
 (3.16).
-The generalized gauge-fixing function (5.33) is the known function of its arguments, while the relation (5.31) is the NL transcendental one.
+The function (5.33) is the known function of its arguments, while the relation (5.31) is the NL transcendental one.
 Behind the general inequality $\xi\neq\xi_{0}$ is the regularized constant $\Delta^{2}_{t}(D)$ as its dynamical source.
-Substituting equations (5.28) and (5.29)-(5.30) into the general decomposition (3.8) for the full gluon propagator, one finally obtains
+Substituting equations (5.31) and (5.33) into the general decomposition (3.8) for the full gluon propagator, one finally obtains
 $D_{\mu\nu}(q)$ $=$ $iT_{\mu\nu}(q){1\over q^{2}+q^{2}\Pi(q^{2};D)+\Delta^{2}_{t}(D)}$ (5.34)
@@
 $q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi(q^{2};\lambda^{-1})=i{\lambda^{-1}q^{2}\over q^{2}+\lambda^{-1}\Delta^{2}_{t}(D)}.$ (5.35)
-This is the generalized gauge-fixing function since it depends on the constant $\Delta^{2}_{t}(D)$, and when it is zero, one recovers the gauge-fixing parameter for the free gluon propagator.
-In this ST identity the gauge-fixing parameter $\lambda^{-1}$ is convenient to vary continuously from zero to infinity.
-The functional dependence of the generalized gauge-fixing function $\xi$ (5.33) is fixed up to an arbitrary gauge-fixing parameter $\xi_{0}=\lambda^{-1}$.
-Unless we fix it, and thus $\xi$ itself, we will call such situation as the general gauge dependence (GGD), see equations (3.7)-(3.8), (5.32)-(5.34).
+The ST identity (5.35) depends on the constant $\Delta^{2}_{t}(D)$, and when it is zero, one recovers the gauge-fixing parameter for the free gluon propagator.
+In this ST identity the gauge-fixing parameter $\lambda^{-1}$ for the free gluon propagator is convenient to vary continuously from zero to infinity.
+The functional dependence of $\xi$ (5.33) is fixed up to an arbitrary gauge-fixing parameter $\xi_{0}=\lambda^{-1}$.
+Unless we fix it, and thus $\xi$ itself, we will call such situation as the generalized gauge dependence (GGD), see equations (3.7)-(3.8), (5.32)-(5.35).
 Choosing $\xi_{0}=\lambda^{-1}$ explicitly, we will call such situation as the explicit gauge dependence (EGD).
@@
 The generalized gauge requires that there is no other functional expression for $\xi$, apart from given by the relation (5.33) at finite $\xi_{0}=\lambda^{-1}$, in the full gluon propagator (5.34) and the ST identity (5.35) for the regularized gluon fields.
-The system of the regularized equations (5.34)-(5.35) constitutes that the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is not a symmetry of its ground state, and it has been exactly and uniquely defined.
-They are present in the form suitable for the NP renormalization program to be performed.
+For the renormalized version of the ST identity (5.35) see section 6 below.
+The system of the regularized equations (5.34)-(5.35), explicitly depending on the tadpole term, are present in the form suitable for the NP renormalization program to be performed.
 The tadpole term enters the full gluon self-energy linearly, see the expression (3.12).
@@
 (5.28).
-The term $(\Delta^{2}_{t}(D)/q^{2})$ in the expressions (5.34)-(5.34) is to be suppressed in the PT $q^{2}\rightarrow\infty$ at the finite that $\xi_{0}=\lambda^{-1}$ (the above-mentioned canonical gauge will be investigated in detail in the forthcoming paper).
-The full gluon propagator will behave like the free gluon propagator $\sim 1/q^{2}$.
+The term $(\Delta^{2}_{t}(D)/q^{2})$ in the expressions (5.34)-(5.35) is to be suppressed in the PT $q^{2}\rightarrow\infty$ limit at the finite $\xi_{0}=\lambda^{-1}$ (the above-mentioned canonical gauge will be investigated in detail in the forthcoming paper).
+Then the full gluon propagator will behave like the free gluon propagator $\sim 1/q^{2}$ in this limit.
 This is the one of the necessary constraints that a theory is perturbative renormalizable.
 Other ones such as the corresponding behaviour of the spinor Green’s function, a unitary of $S$-matrix and analyticity (causality) [2], evidently are beyond the scope of the present work.
-Here we have proven that the true dynamical and gauge structures of the QCD/YM ground state are much more complicated than it is required by its Lagrangian’s gauge symmetry.
-But despite the gauge symmetry is broken in the QCD ground state, its PT renormalizability is not affected within the QCD full gluon propagator in the generalized gauge (5.33), i.e., it has the PT renormalizable behaviour at large $q^{2}$.
+Here one can conclude that the true dynamical and gauge structures of the QCD/YM ground state are much more complicated than it follows from its Lagrangian’s formulation.
+However, its PT renormalizability is not affected within the QCD full gluon propagator in the generalized gauge (5.34), i.e., it has the PT renormalizable behaviour at large $q^{2}$.
 Concluding, it is worth noting that in terms of the transverse relations and the corresponding QD regularized constants as well as the corresponding relations between $\xi$ and $\xi_{0}$, the general solution can be shown explicitly as the following system of the relations, namely
@@
 $\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0,\ \Delta^{2}_{t}(D)\neq 0,\ \xi\neq\xi_{0}=\lambda^{-1}.$ (5.37)
-It is necessary to underline that none of the prescriptions (i.e., putting some of them zero by hand) have been introduced and none of the truncations, approximations and assumptions have been done as well in the obtaining of these relations.
+It is necessary to underline that none of the prescriptions (i.e., putting some of them zero by hand) have been introduced and none of the truncations, approximations and assumptions have been done, as well as no gauge fixing by hand, in the obtaining of these relations.
 In other words, all of them are exact mathematical results in the general solution to QCD.
-6 Particular solution to QCD
-The widely used up to present days the linear relation $\xi=\xi_{0}$ is only a particular solution of the derived here the exact constraint (3.15).
+6 Renormalized ST identity and mass gap
+Since the novel constraint derived in section 3 has been mainly based on the ST identities formalism, it is possible to perform the NP renormalization program for the ST identity (5.35) itself.
+It is instructive to present (5.35) in terms of $\xi$ and $\xi_{0}$ again, then we have
+$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi=i{\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)}.$ (6.38)
+Let us now introduce the NP renormalization constant for the tadpole term as follows: $\Delta^{2}_{t}(D)=Z_{\Delta}\Delta^{2}$, where $\Delta^{2}$ is finite and positive, by definition, while the NP renormalization constant $Z_{\Delta}$ depends on all the unphysical parameters, such as regulating parameters, mentioned above, etc.
+In principle, the string tension in the linear rising potential between heavy quarks has to be somehow related to the finite $\Delta^{2}$, since it is connected to the transverse projection operator structure, as underlined above.
+It is well-known that the confining ansatz $D_{\mu\nu}(q)\sim T_{\mu\nu}(q)\Delta^{2}/(q^{2})^{2}$ just leads to a such potential.
+So that, the ST identity for the renormalized full gluon propagator becomes
+$q_{\mu}q_{\nu}D^{R}_{\mu\nu}(q)=i\xi^{R}=i{\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}},$ (6.39)
+where the renormalized gauge-fixing parameters are defined as follows: $\tilde{\xi}_{0}=Z_{\Delta}\xi_{0},\ \xi^{R}=Z_{\Delta}\xi$ and $D^{R}_{\mu\nu}(q)=Z_{\Delta}D_{\mu\nu}(q)$, i.e., the ST identity (6.39) is expressed in terms of the finite quantities only.
+The correct PT limit is maintained, of course.
+The renormalized version of the tadpole term $\Delta^{2}$ can be considered as a scale determining the NP structure of the general QCD.
+Conventionally, we call it as a mass gap and see discussion in section 10 as well.
+Such a finite mass gap $\Delta^{2}$ will appear in the full gluon propagator as well, i.e., not only in its longitudinal part (see sections 8 and 9).
+Let us now demonstrate the interesting feature of the generalized gauge (6.39). It provides the self-consistency condition for the gauge choice in QCD as follows:
+${\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}}={aq^{2}\over q^{2}+a\Delta^{2}},$ (6.40)
+[…]
+${\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}}={q^{2}\over\Delta^{2}},$ (6.41)
+which is only satisfied at $q^{2}=0$, i.e., there is no any condition for the gauge-fixing parameter.
+The formulated self-consistency condition (6.40) points out on the inconsistency of the canonical gauge $\tilde{\xi}_{0}=\infty$ in QCD.
+Briefly, if everything is expressed in the terms of the finite quantities then the formal canonical gauge is forbidden to use in QCD, according to our approach to this theory.
+7 Conventional QCD
+The widely used up to present days the linear relation $\xi=\xi_{0}$ is only a particular solution to the exact constraint (3.15).
 As mentioned above in section 3, it is defined when the tadpole term has to be put zero by hand, i.e., $\Delta^{2}_{t}(D)=0$, then $\xi=\xi_{0}=\lambda^{-1}$ as well and vice versa.
-The corresponding expressions for the full gluon propagator and the ST identity in the particular solution to QCD can be independently derived absolutely in the same way as has been derived the corresponding expressions for the general solution to QCD in the previous section.
-However, the simplest way to show the particular solution is to put formally $\Delta^{2}_{t}(D)=0$ in the system of eqs.
+The corresponding expressions for the full gluon propagator and the ST identity in this solution to QCD can be independently derived absolutely in the same way as has been derived the corresponding expressions for the general solution to QCD in the previous section.
+However, the simplest way to show the conventional solution is to put formally $\Delta^{2}_{t}(D)=0$ in the system of eqs.
 (5.34)-(5.35), then it looks like
@@
 Obviously, the expression for the PT full gluon propagator (7.42) describes the propagation of the PT massless gluons, since it has the PT singularity on the mass-shell $q^{2}=0$, i.e., the singularity of the free gluon propagator $\sim 1/q^{2}$.
+The gluon SD equation (7.42) does not provide any hint how to prevent the free and the PT massless gluon states to appear at large distance ($q^{2}\rightarrow 0$), i.e., this theory is not confining, as emphasized above.
 Let us note in advance that the equality $\xi=\xi_{0}=\lambda^{-1}$ takes place only for the regularized massless gluon fields.
 For their renormalized counterparts these gauge-fixing parameters differ by the corresponding renormalization constant.
-The system of the regularized eqs.(6.38)-(6.41), which is uniquely defined, is free of all the types of the scale parameters having the dimensions of a mass squared, forbidden by the exact gauge symmetry of the QCD Lagrangian.
-Therefore, it constitutes that the gauge symmetry of the QCD ground state, reflected by this system, coincides with the symmetry of its Lagrangian.
-Such coincidence is analogous to QED where the abelian $U(1)$ gauge symmetry of the Lagrangian is the same as of its ground state (omitting formally the YM degrees of freedom here, this coincidence has been proven just in this way in [10]).
-However, it is necessary to remind that the coupling constant remains strong, so to call it as $D^{PT}$ is only a mere convention, that is why the particular solution to QCD can be called as conventional QCD as well.
-Just this conventional solution to QCD has been investigated in full details the last five decades since its formulation in [1].
-The essential feature of the above-mentioned coincidence is the equality $\xi=\xi_{0}$.
-In this connection, let us note that one can start from the expression (3.8).
-It is the general decomposition of the full gluon propagator, depending on the one variable, into the independent tensor structures.
-Then it will automatically satisfy to the relation (3.7), called the ST identity, determining thus the gauge-fixing parameter.
-This has nothing to do with the statement that the ST identity is a consequence of the exact gauge symmetry.
-It becomes its consequence, indeed, when $\xi$ is fixed by the relation $\xi=\xi_{0}=\lambda^{-1}$.
-The well-known PT linear relation $\xi=\xi_{0}$ is the consequence of the exact gauge invariance both for the QCD Lagrangian and its ground state.
-In principle, the PT QCD or, equivalently, the conventional QCD is a well-elaborated quantum gauge field theory [2, 3, 4, 5, 6, 7, 8, 9], and for its framework of the corresponding dynamical equations of motion see recent review [34] (and references therein).
-However, this theory is not our concern (but for some clarifying discussion see appendix B).
+The conventional or, equivalently, the PT QCD is a well-elaborated quantum gauge field theory [2, 3, 4, 5, 6, 7, 8, 9], and for its framework of the corresponding dynamical equations of motion see recent review [34] (and references therein).
+Just it has been investigated in full details the last five decades since its formulation in [1].
+It is necessary to remind that the coupling constant remains strong, so to call it as $D^{PT}$ is only a mere convention.
+This theory is not our concern.
 We present it here within our formalism for the readers convenience to directly compare it with the general solution to QCD.
-For this aim it is useful to reproduce the system of the relations (5.36)-(5.37) for the PT QCD. It will look like
+For this aim it is useful to reproduce the system of the relations (5.36)-(5.37) for the PT QCD.
+It will look like
 $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)$ $=$ $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D^{PT})=0,$
@@
 $\Delta^{2}_{q}=\Delta^{2}_{g}(D^{PT})=0,\ \Delta^{2}_{t}(D^{PT})=0,\ \xi=\xi_{0}=\lambda^{-1}.$ (7.47)
-The first equalities in these relations coincide with the corresponding equalities in the relations (5.36)-(5.37), i.e., in both theories the general and the PT QCD or, equivalently, the particular QCD, they are exact mathematical results.
+The first equalities in these relations coincide with the corresponding equalities in the relations (5.36)-(5.37), i.e., in both theories the general and the PT QCD they are exact mathematical results.
 At the same time, the second equalities in these relations remain prescriptions, since they have to be put zero by hand, as emphasized above.
@@
 However, from now on this prescription has been put on a firm mathematical ground.
-How to understand such kind of equalities has been explained in section 5.
-7 Discussion
-Let us now discuss some interesting physical consequences of the general solution to QCD.
-It requires the explicit presence of the tadpole term in the QCD ground state at the fundamental quark-gluon level, i.e., it cannot be discarded in the QCD/YM theory by any means.
-This was exactly proven in section 4 and used in section 5.
-Then it will drastically change the structure of the full gluon propagator in the deep IR region $q^{2}\rightarrow 0$.
-For the sake of transparency and clarity let us use the free gluon propagator in the t’ Hooft-Feynman gauge $\lambda^{-1}=1$, i.e., $D^{0}_{\mu\nu}(q)=i{\delta_{\mu\nu}/q^{2}}$.
+How to understand such kind of equalities has been already explained in section 5.
+8 Confinement of the free gluon states
+For the sake of transparency and simplicity, we use the free gluon propagator in the t’ Hooft-Feynman gauge $\lambda^{-1}=1$, i.e., $D^{0}_{\mu\nu}(q)=i{\delta_{\mu\nu}/q^{2}}$.
 Then the full gluon propagator (5.28), on account of the definitions (5.29) and (5.30), up to the first iteration term looks like
@@
 Evidently, such a summation is only possible due to the existence of the tadpole term.
-The dependence of the coefficients $\Phi_{k}(g^{2},...)$ on the coupling constant is not known, i.e., such kind of series are the cluster expansions, mentioned above [23].
-These coefficients, being the sum of the infinite number of terms by themselves, are resulted from the skeleton loop integrals, contributing into full gluon self-energy through the invariant function $\Pi(q^{2};D)$ because of the NL character of the iteration procedure.
+The dependence of the coefficients functions $\Phi_{k}(g^{2},...)$ on the coupling constant is not known, i.e., such kind of series are the cluster expansions, mentioned above [23].
+These coefficients, being the sum of the infinite number of terms by themselves, are resulted from the skeleton loop integrals, contributing into the full gluon self-energy through the invariant function $\Pi(q^{2};D)$.
+Let us note here that eq.
+(8.49) can be expressed in terms of the mass gap through the relation $\Delta^{2}_{t}(D)=Z_{\Delta}\Delta^{2}$ and further including $Z_{\Delta}$ into the above-mentioned coefficient functions.
 Evidently, the iteration ab  infinity leads to the essential singularity at $q^{2}\rightarrow 0$!
-in the full gluon propagator (7.44), which will remain massless, but depending on the tadpole term.
+in the full gluon propagator (8.49), which will remain massless, but depending on the tadpole term (scale breaking in the IR region).
 Let us note that the same series can be obtained by expanding the initial NL transcendental relation (5.31) in powers of the ratio $(\Delta^{2}_{t}(D)/q^{2})$.
-From the NL iteration expressions (7.44)-(7.45) it follows that the full gluon propagator in the IR region ($q^{2}\rightarrow 0$) is dominated by the tadpole contribution, while in the UV region ($q^{2}\rightarrow\infty$) it will be dominated by the free gluon one.
+From the NL iteration expressions (8.48)-(8.49) it follows that the full gluon propagator in the IR region ($q^{2}\rightarrow 0$) is dominated by the tadpole term contribution, while in the UV region ($q^{2}\rightarrow\infty$) it will be dominated by the free gluon one.
 This is completely different to QED, where the dressed photon propagator behaves at small photon momentum as its free counterpart ($\sim 1/q^{2}$).
-This is also true for the full gluon propagator in the PT QCD with unbroken gauge symmetry, investigated in section 6.
+This is also true for the full gluon propagator in the conventional QCD, investigated in section 6.
 The principal difference between these theories is due to the existence of the tadpole term in the QCD ground state, which precisely controls the structure of a single full gluon propagator at small $q^{2}$.
@@
 This is important since there is no any other dynamical mechanism to confine these gluon states to the vacuum of QCD or to keep them inside hadrons.
-To prove the confinement of the singular full gluon propagator itself due to the above-mentioned essential singularity, though possible, but requires completely different investigation within the theory of distributions [26], and thus is the subject of the future work.
-Let us note that by the NP program we mean that the full gluon propagator will become free from all the non-physical parameters of the theory in any of its possible solutions (not only in the iteration one).
+To prove the confinement of the singular full gluon propagator itself due to the above-mentioned essential singularity, though possible, but requires completely different investigation within the theory of distributions [26].
+It is the subject of the future work.
+Let us note that by the NP program we mean that the full gluon propagator will become free from all the non-physical parameters (like the coefficient functions mentioned above) of the theory in any of its possible solutions (not only in the iteration one).
+9 Asymptotic freedom
+let us now investigate the PT $q^{2}\rightarrow\infty$ limit of the full gluon propagator (5.34).
+For this aim it is enough to neglect the ratio $(\Delta^{2}_{t}(D)/q^{2})$ in the invariant function eq.
+(5.31).
+Also, to replace the the invariant function $\Pi(q^{2})$ by its to leading order logarithm, and to introduce the following relation $d(q^{2})=\alpha_{s}(q^{2};\Lambda^{2})/\alpha_{s}(\lambda)$, where $\Lambda^{2}$ and $\lambda$ are the UV regulating parameter and its dimensionless counterpart, respectively.
+Then eq.
+(5.34) will look like
+$D_{\mu\nu}(q)\sim iT_{\mu\nu}(q)\Bigl{[}{\alpha_{s}(\lambda)\over 1+\alpha_{s}(\lambda)b_{0}\ln(q^{2}/\Lambda^{2})}\Bigr{]}{1\over q^{2}}+...,$ (9.50)
+where $b_{0}>0$ is the colour group factor and $\alpha_{s}(\lambda)$ is the coupling constant squared.
+Here and below the trivial longitudinal part is not shown, for simplicity.
+Since we already know that the finite mass gap exists, then without loosing generality, one can put $\Lambda^{2}=f(\lambda)\Delta^{2}=f(\lambda)a^{-1}a\Delta^{2}=f^{\prime}(\lambda)\Lambda^{2}_{QCD}$ with $\Lambda^{2}_{QCD}=a\Delta^{2}$, where $a$ is a some finite number because one cannot identify directly the mass gap $\Delta^{2}$ with $\Lambda^{2}_{QCD}$.
+The dimensionless functions $f^{\prime}(\lambda)=a^{-1}f(\lambda)$ behave like $\sim\lambda^{2}$ in the $\lambda\rightarrow\infty$ limit, otherwise remaining arbitrary.
+Doing some algebra, one obtains
+$D_{\mu\nu}(q)\sim iT_{\mu\nu}(q)\Bigl{[}{\alpha_{s}\over 1+\alpha_{s}b_{0}\ln(q^{2}/\Lambda^{2}_{QCD})}\Bigr{]}{1\over q^{2}}+...,$ (9.51)
+where finite $\alpha_{s}=\alpha_{s}(M_{Z})=0.1184$ is determined by the relation connecting $\alpha_{s}(\lambda)$ to $\ln f^{\prime}(\lambda)$.
 Eq.
-(5.34) shows that gluon may acquire mass dynamically if the denominator of this equation becomes zero at some point $q^{2}=M^{2}$.
-The expression for the free massive vector particle propagator follows in a rather simple way, indeed.
-For this goal it is necessary to omit the corresponding invariant function and identify the tadpole term with the gluon mass $M^{2}$.
-In Minkowski metric (by replacing $q^{2}\rightarrow-q^{2}$ and $\delta_{\mu\nu}\rightarrow g_{\mu\nu}$), one gets
-$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2})}\left[g_{\mu\nu}-{(1-\lambda^{-1})q_{\mu}q_{\nu}\over(q^{2}-\lambda^{-1}M^{2})}\right],$ (7.46)
-i.e., it will be expressed in the Stueckelberg gauge [4].
-Contracting this expression with respect to $q_{\mu}$ and $q_{\nu}$, one obtains the ST identity (5.35) by replacing there $q^{2}\rightarrow-q^{2}$.
-In other words, the Stueckelberg gauge (7.46) is a particular case of the generalized gauge-fixing function (5.33).
-The PT QCD full gluon propagator (6.38) cannot generate the gluon mass at some point, since it remains always massless.
-The full gluon propagator (5.34) provides a rather convenient framework in order to formulate and further develop the NP renormalization program for the massive gluon fields.
-This program will suppose to show that the full gluon propagator (5.34), and hence its free massive counterpart (7.46), can not be put on the mass-shell $q^{2}=M^{2}_{g}$, where $M^{2}_{g}$ is the finite gluon pole mass.
-Acquiring mass dynamically, the gluon, nevertheless, will remain the off-mass-shell object.
-So that, it will not appear in the physical spectrum.
-Some aspects of the massive gluon field’s renormalization are discussed in appendix C.
-In the presence of the mass squared scale parameter - the tadpole term - the role of the QCD coupling constant $g^{2}$ becomes unimportant.
-This is also evidence of the ’dimensional transmutation’, $g^{2}\rightarrow\Delta^{2}_{t}(D)$ [2, 36, 37], which occurs whenever a massless theory acquires mass dynamically.
-It is a general feature of spontaneously symmetry breaking in field theories.
-We distinguish between the PT/particular/conventional and general solutions to QCD by the explicit presence of the tadpole term in the latter one, and not by the magnitude of the coupling constant even at the regularized gluon fields level yet.
-In the both cases the gluon fields remain strongly interacted, apart from the AF regime [38, 39].
-So that, the general solution to QCD, investigated in section 5 is possible to call as the NP QCD as well.
-Any deviation of the full gluon propagator from the free one in QCD requires the presence of a mass squared scale parameter. Even in the AF regime $q^{2}\rightarrow\infty$ there is a scale breaking,
-$D_{\mu\nu}(q)\sim\delta_{\mu\nu}\Bigl{[}{ig^{2}\over 1+g^{2}b_{0}\ln(q^{2}/\Lambda^{2}_{QCD})}\Bigr{]}(1/q^{2}),$ (7.47)
-which has been discovered many years ago [37, 38].
-When $q^{2}\rightarrow\infty$ the dependence on the coupling constant $g^{2}$ vanishes, indeed.
-The well-known expression (7.47) is written down in the t’ Hooft-Feynman gauge and in the approximation of the main PT logarithms with the colour group factor $b_{0}>0$.
-$\Lambda^{2}_{QCD}$ determines the scale of the non-trivial PT dynamics in the QCD.
-However, in section 6, we have already shown that the PT cannot generate the scale parameter, having the dimension of a mass squared.
+(9.51) has been discovered many years ago [38, 39].
+In the PT $q^{2}\rightarrow\infty$ limit the dependence on the coupling constant $\alpha_{s}$ vanishes, indeed.
+The well-known expression (9.51) is written down in the so-called approximation of the main PT logarithms and $\Lambda^{2}_{QCD}$ determines the scale of the non-trivial PT dynamics in the QCD.
+However, in section 6, we have already shown that the PT QCD cannot generate the scale parameter, having the dimension of a mass squared.
 From the renormalization group equations analysis, it follows that any mass to which can be assigned any physical meaning is exponentially suppressed in the weak coupling limit to every order of the PT [2, 36], namely
@@
 It is entirely of the NP origin!
-It has to come from the IR region which is controlled by the tadpole term, as discussed just above.
-Its finite version should be somehow related to $\Lambda^{2}_{QCD}$ (reminding that the ratio $(\Delta^{2}_{t}(D)/q^{2})$ may appear under the PT logarithms).
-The explanation of the AF phenomenon (i.e., a scale breaking in the PT regime) will be given when the NP renormalization program of the singular gluon field configurations will be performed.
-8 Conclusions
-We have derived the novel exact constraint (3.15) on any solution to QCD.
-For this aim the ST identities for the gauge particles Green’s functions (propagators), complemented by the satisfied transverse relations for the quark and gluon contributions into the full gluon self-energy, have been used.
-The first of its solutions is called the general solution to QCD and given by the expression (3.16), which formulates the generalized gauge-fixing function approach to QCD.
-The existence of a new general solution to QCD makes it possible to prove in the most general and unique mathematical ways that the $SU(3)$ colour gauge symmetry of the QCD Lagrangian is broken in its ground state.
-However, the PT renormalizability will not be affected within the framework of the full gluon propagator.
-In the PT $q^{2}\rightarrow\infty$ regime it will have the PT renormalizable behaviour $\sim 1/q^{2}$.
-The dynamical source of the Lagrangian’s gauge symmetry breakdown in the ground state has to be identified with the constant tadpole term, having the dimension of a mass squared.
-It is dynamically generated by the self-interaction of the multiple massless gluon modes involving only the point-like 4-gluon vertex, see eq.
-(2.4) and Fig.
-1.
-The full 4-gluon vertex survives when all the gluon momenta involved go to zero, becoming thus its point-like counterpart, while the full 3-gluon vertex does not survive in this limit.
-Just this shows the importance of the point-like 4-gluon vertex in the deep IR structure [10, 23, 40] of the theory.
-By the design the tadpole term is present in the QCD ground state from the very beginning, i.e., not introducing by hand.
-It is only one which explicitly contributes into the skeleton loops decomposition of the full gluon propagator (shown in Fig.
-1), while all the other tadpole-like terms don’t.
-Unlike them, being the QD but already regularized constant, the tadpole term is not connected to any external gluon momentum, and thus may generate a mass squared scale parameter [6], indeed.
-Thus, there is no need to extract such a scale parameter in the QCD vacuum in a some other way.
-In [10] the importance of the tadpole term has been recognized and exploited on an intuitive basis rather than a logical one.
-However, from now on its existence cannot be undermined.
-Besides its combinatorial (mathematical) meaning it has a deep dynamical one as a source of the gauge symmetry violation in the QCD vacuum.
-Its renormalized (i.e., finite) version will determine a scale of the NP dynamics in QCD and in its true ground state.
-Due to these impacts our advance findings are:
-a.
-The novel general constraint (3.15) on any solution to QCD has been exactly derived before any proper subtraction scheme has been introduced.
-The importance of the ST identities and the corresponding satisfied transverse relations has been explicitly shown.
-b. Its exact solution (3.16) determines the generalized gauge-fixing function approach to QCD.
-c. The proper subtraction scheme has been formulated to remove from the theory the QD constants, apart from the tadpole term, on the general mathematical basis.
-d. The system of the relations (5.36) are exact mathematical results, i.e., they are not prescriptions.
-e.
-The new NP approach to the YM theory and its ground state has been formulated, based on the regularized system of eqs.
-(5.34)-(5.35).
-It is defined up to the invariant function with the known asymptotic properties.
-f.
-This system has a correct the PT limit, and is suitable for the NP renormalization program to be performed.
-It is called as the general solution to QCD, and just it resolves the dilemma formulated at the end of section 2.
-In obtaining all these results no any truncations, approximations and assumptions have been made in dealing with the regularized skeleton loop integrals, contributing to the gluon self-energy.
-The exact analytic derivations have been done, based on a tensor algebra rigorous rules.
-The second of the above-mentioned solutions is called the particular solution to QCD, and determined by the relations $\Delta^{2}_{t}(D)=0$ and $\xi=\xi_{0}$ in the exact constraint (3.15).
-This solution is nothing else but the well-known PT or, equivalently, conventional QCD, which has been studied last five decades.
-We confirm that the second solution preserves the gauge symmetry of the QCD Lagrangian in its ground state.
-Both equalities, shown above, are simply prescriptions, i.e., they have to be put valid by hand, as it has been discussed in section 6.
-In this connection, let us remind that for the tadpole term does not exist the independent mathematical relation according to which it has to be omitted from the theory, while such kind of the relations exist for the the quark and gluon QD regularized constants (see section 4).
-That is why the tadpole term has to be put zero by hand for the particular solution, which comes from the solution of the exact constraint (3.15), and see also discussion in appendix B.
-We can registered the existence of the two different types of the gluon field configurations in the QCD vacuum: the general one and its particular counterpart with the gauge symmetry broken and unbroken, respectively.
-The just derived general solution to QCD with a gauge symmetry breaking, but maintaining its PT renormalizability, will apparently satisfy the boundary conditions in QCD at the fundamental quark-gluon level: the confinement at large distances ($q^{2}\rightarrow 0$) and the AF at short distances ($q^{2}\rightarrow\infty$), as discussed in section 7.
-The particular or, equivalently, the PT/conventional QCD solution with a gauge symmetry preservation cannot satisfy the boundary condition at large distances.
-It shows the agreement with the AF regime at short distances, but cannot explain it, see discussion in section 7 as well.
-The perspectives for further work are as follows: to formulate the NP multiplicative (MP) renormalization programs for the regularized full gluon propagator (5.34).
-This is not a simple task, since such kind of a programs have to be different for the massive and severely singular gluon field configurations, briefly discussed in section 7.
-For this purpose it is necessary to take into account the transcendental and NL structure of eq.
+It has to come from the IR region which is controlled by the tadpole term, as discussed above.
+So that, the scale breaking appears under the PT logarithm, while the gluon itself remains massless like in eq.
+(8.49), where the scale breaking appears in the NL way.
+In the both cases the scale breaking masses cannot be directly identified with the mass gap, defined in section 6.
+It should be multiplied by the finite numbers different in each case, in order to get scale breaking in the IR and UV regions.
+Its exact derivation, and thus the explanation of the AF phenomenon (i.e., a scale breaking in the PT regime) will be given when the corresponding NP renormalization program for the full gluon propagator (5.34) will be performed.
+Concluding, all this can be manifestation that ”the problems encountered in perturbation theory are not mere mathematical artefacts but rather signify deep properties of the full theory” [43].
+The message we are trying to convey is that the AF phenomenon is the first direct evidence of the validity of the general QCD.
+In our opinion, the gluon confinement and AF are the two boundary conditions at low- and high-energies in the QCD and its vacuum, to which the general QCD satisfies, while the conventional QCD does not.
+For these features of the general QCD, one can call it as the confining QCD theory as well.
+Finally, it is worth emphasizing once more that in the both asymptotics of the iteration/singular solution to the confining QCD, the gluons always remain massless.
+So that, the QCD Lagrangian’s gauge symmetry is not violated by it, but the mass gap is needed in order to provide its scaling properties, as described above.
+10 Discussion and outlook
+The general QCD requires the explicit presence of the mass squared scale parameter - the tadpole term - in the QCD ground state, i.e., at the fundamental quark-gluon level, as this was exactly proven in section 4 and used in section 5.
+In its presence the role of the QCD coupling constant $g^{2}$ becomes unimportant.
+This is also evidence of the ’dimensional transmutation’, $g^{2}\rightarrow\Delta^{2}_{t}(D)$ [2, 36, 37], which occurs whenever a massless theory acquires mass dynamically.
+It is a general feature of spontaneously symmetry breaking in field theories.
+We distinguish between the PT/particular/conventional and general QCD by the explicit presence of the tadpole term in the latter one, and not by the magnitude of the coupling constant even at the regularized gluon fields level yet.
+In the both cases the gluon fields remain strongly interacted, apart from the AF regime [38, 39].
+The general character of our results is convenient to summarize in connection with the Jaffe-Witten’s (JW) [23] theorem, formulated as follows:
+JW’s theorem: Yang-Mills existence and Mass Gap.
+Prove that for any compact simple gauge group $G$, a non-trivial quantum Yang-Mills theory exists on $R^{4}$ and has a mass gap $\Delta>0$.
+Existence includes establishing some strong axiomatic properties for the Euclidean Green’s functions.
+In their description of the theorem it is explained why a correct QCD theory must have the following three properties: 1).
+The nuclear force is strong but short-ranged due to the existence of the ”mass gap”, a scale parameter $\Delta>0$.
+2).
+The physical particle states are $SU(3)$ invariant,i.e., confinement of colour degrees of freedom at the fundamental quark-gluon level.
+3).
+The chiral symmetry breaking to account for the ”current algebra” theory of soft pions at the hadronic level.
+The definition of the mass gap given in [23] as the solution of the Hamiltonian problem that every excitation of the vacuum has energy at least $\Delta>0$ is the mass gap responsible for the second PCAC phase transition at the hadronic level (in our opinion).
+For example, the glueball state with the most lowest mass, though none of such state has been discovered yet (but this is not a problem, remanding that the Higgs particle [41] has been experimentally found after many years of its prediction).
+The literary meaning/definition of the mass gap, mentioned above, may appear at the fundamental quark-gluon level as well in the general QCD only in its massive solution for the full gluon propagator (5.34).
+The gluon may acquire mass dynamically, indeed, if the denominator of this equation becomes zero at some finite point $q^{2}=M^{2}=\Delta^{2}_{t}(D)$.
+This solution is completely different case from the singular solution, and will be subject for the independent consideration.
+The request for a mass gap in the general QCD, can be also summarized like the JW theorem:
+Theorem I: Mass Gap existence and Yang-Mills theory.
+If a non-trivial quantum Yang-Mills theory with gauge group $SU(3)$ exists on $R^{4}$ then it has a mass gap $\Delta^{2}>0$, responsible for the first confinement phase transition in QCD.
+Existence includes establishing the confining QCD theory, which remaining the PT renormalizable.
+In connection with our theorem, let us note that the tadpole term cannot be removed from the QCD theory and its vacuum by any means, apart from putting it zero by hand in order to get to the conventional QCD.
+Its renormalized (finite) version survives the NP renormalization and thus can be assigned a physical meaning as a scale parameter responsible for the NP effects in the QCD, leading to the formulation of the confining QCD theory.
+Just the existence of a such scale, conventionally called as a mass gap, provides the suppression of the PT and free gluon states to appear at large distances, i.e., in the physical spectrum, as explained above.
+It is dynamically generated by the self-interaction of the multiplied massless gluon modes, dominated by the point-like four-gluon vertex [10, 23, 40], i.e., it is a fundamental quantity (”mass without mass” [44]), so that it cannot be a bound state of anything.
+In [10] the importance of the mass gap has been recognized and exploited on an intuitive basis rather than a logical one, but from now on its existence cannot be undermined.
+A few additional clarifying remarks are in order.
+Let us remind that the present QCD theory, i.e., conventional one in our terminology, possesses none of the above-mentioned three necessary properties required by JW for QCD to be successful theory of the strong interactions at the both microscopic and macroscopic levels.
+Unlike this theory, the confining QCD theory, discovered in this work, possesses at least the two of those properties.
+It has a characteristic mass scale parameter even at the fundamental quark-gluon level.
+Precisely, the existence of this scale in the theory confines the colour gluons to the QCD ground state.
+We strongly believe that quark confinement and ”chiral symmetry breaking” will be also confirmed after solution of the quark full SD equation by the substitution of the full gluon propagator, determined within the confining QCD theory.
+It is worth emphasize once more that we have explicitly shown that the first confinement phase transition in QCD should have its own finite scale parameter $\Delta^{2}$, reflected in our Theorem I.
+In order to describe the scales of the different NP effects, such as the scales breaking, the existence of the string tension, etc., it should undergo the finite transformations, i.e., to be multiplied by the different finite constants.
+Taking into account that the JW theorem has been formulated in the most general terms, absorbing the different interpretations/meanings/understanding of a mass gap itself, it was quite natural and suitable for us to adopt the JW mass gap name for our characteristic mass scale parameter.
+In other words, we understand the mass gap conception in more wider context than its definition given by JW.
+Nevertheless, to summarize all our advance results in terms of the theorem similar to the JW theorem, makes sense and is very useful for further development of the correct (confining) QCD theory and understanding of its true vacuum dynamical and gauge structures.
+In obtaining them no any truncations, approximations and assumptions have been made in dealing with the regularized skeleton loop integrals, contributing into the full gluon self-energy.
+The exact analytical derivations have been done, based on a tensor algebra rigorous rules only.
+The perspectives for further work are as follows: the NP renormalization program of the ST identity performed in section 6 was enough to prove the existence of a mass gap, as stated in Theorem I.
+However, to formulate the NP multiplicative (MP) renormalization program for the regularized full gluon propagator (5.34) is not a simple task.
+Such kind of a program has to take into account of all the peculiarities of the severely singular gluon field configurations, briefly discussed in section 8: first of all, the transcendental character and the NL structure of eq.
 (5.31).
-They are also needed to be formulated in the most general and unique ways.
-To ensure a unitary of $S$-matrix elements by these NP MP renormalization programs should be requested as well.
-In general, our results open a new wide research areas in modern physics, leading to the multiple theoretical and phenomenological applications and comparing them with the results of the conventional QCD.
+It has to be replaced by the corresponding cluster expansion, suitable for the NP renormalization programs to be performed.
+This is also needed to be formulated in the most general and unique ways.
+To ensure a unitary of $S$-matrix elements by these NP MP renormalization program should be requested as well.
+In general, our results open a new wide research areas in the general QCD, leading to the multiple theoretical and phenomenological applications and comparing them with the results of the conventional QCD.
+In this connection, let us note that the general QCD may have another type of the solution, different from the singular one, due to the above-underlined transcendental character and the NL structure of eq.
+(5.31), namely the massive gluon solution, mentioned above.
 To understand a mass dynamical generation at the quark-gluon level is the first necessary step to understand the existence of the physical mass spectrum at the hadronic level in QCD as a theory of strong interactions.
@@
 QCD is a self-consistent quantum gauge field theory.
-It does not need any extra degrees of freedom (such as Higgs fields [41]) to generate a mass at the fundamental quark-gluon level.
-Concluding, a few remarks are in order.
-The Elitzur’s theorem [42] has been proven in lattice field theory for the abelian system like QED (see appendix D as well).
-To the non-abelian systems like YM fields, it also can be applied, but only within its particular solution, as discussed above.
-At the same time, this theorem cannot be applied to the YM fields in the frame work of its exact solution for the generalized gauge-fixing function in the full gluon propagator.
-In this paper we have just shown that the gauge symmetry of the continuos YM theory can be broken in its ground state.
-Behind this effect is the mass squared parameter, which is nothing else but the renormalized tadpole term.
-It survives the NP renormalization but does not affect the PT renormalizability, as underlined just above.
+It does not need any extra degrees of freedom (such as Higgs fields [41]) or to extract the mass by some other ways to generate a mass at the fundamental quark-gluon level.
 The relevant existing knowledge on a mass dynamical generation and spontaneously symmetry breaking subjects has been taken into account in the list of publications, presented in our paper (and see references in each of the cited items there).
-Also, let us note that contrary to the ghost term, the QCD Lagrangian’s gauge-fixing term is not explicitly present in the full gluon self-energy, that is why it is left out of our attention.
-It would be interesting to establish the character of its possible connection (if any) to the NL generalized gauge-fixing function approach.
-Apparently, this might be done only after the corresponding NP MP renormalization programs will be performed, not before, in our opinion.
-9 Summary
-Briefly, we can summarize our results as follows:
-The exact derivation of the general constraint (3.15) on any solution to QCD is present.
-This has been done in the mathematically understandable way for a broad readership of researches, since exploiting only the tensor algebra derivations rigorous rules, widely used in the theoretical physics, as pointed out above.
-It has the two possible exact solutions, in principle, independent from each other.
-The general solution to QCD is the solution with the gauge symmetry broken in its ground state.
-The dynamical source of this groundbreaking effect is the tadpole term.
-The general QCD remains the PT renormalizable, nevertheless.
-It will satisfy the boundary conditions at the fundamental quark-gluon level, as discussed in section 7.
-The particular or, equivalently, the PT/conventional solution to QCD is the solution with unbroken gauge symmetry in its ground state.
-It is the PT renormalizable as well, but not satisfying the boundary condition at large distances, as discussed in section 7.
-Its gauge symmetry hides the true role of the tadpole term in the QCD vacuum.
+Let us note that contrary to the ghost term, the QCD Lagrangian’s gauge-fixing term is not explicitly present in the full gluon self-energy, that is why it is left out of our consideration.
 Acknowledgments
@@
 In this respect it is different from all other tadpole-like terms which appear as a result of the subtraction scheme or in the NL iteration procedure.
-Also it directly generate the mass squared parameter, and thus explicitly violets the gauge symmetry of the QCD Lagrangian in its ground state [6].
-We have shown that in the PT QCD or,equivalently, particular QCD (described in section 6) it was a prescription to neglect it, but in the full QCD or, equivalently, NP QCD (investigated in sections 4 and 5) it should remain intact.
-At the time, all other tadpole-type terms have to be disregarded on the exact mathematical basis in both theories.
+We have shown that in the conventional QCD (described in section 6) it was a prescription to neglect it, but in the confining QCD (investigated in sections 4 and 5) it should remain intact.
+At the same time, all the other tadpole-like terms, mentioned above, have to be disregarded on the exact mathematical basis in the both theories.
 Let us now investigate the tadpole-like term with the gluon skeleton loop, shown second in the third line of Fig. A.2. Its analytical expression is
@@
 This was not a prescription, but being an exact mathematical result within our formalism.
-Appendix B The origin of the conventional QCD
-In order to find the general solution to QCD, we present the full gluon self-energy (2.2) as a sum of the three independent terms, i.e., we distinguish between the gluon part and the tadpole term, though their sum forms the YM part of QCD, since the both are determined by the different self-interaction of multiple massless gluon modes.
-However, not doing so, one needs to present the full gluon self-energy in terms of the quark and YM compositions, i.e., one has to operate with
-$\Pi_{\rho\sigma}(q;D)=\Pi^{q}_{\rho\sigma}(q)+\Pi^{YM}_{\rho\sigma}(q;D),$ (2.61)
-where
-$\Pi^{YM}_{\rho\sigma}(q;D)=\Pi^{g}_{\rho\sigma}(q;D)+\delta_{\rho\sigma}\Delta^{2}_{t}(D).$ (2.62)
-The relation (3.13) in this case becomes
-$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{YM}_{\rho\sigma}(q;D)]={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (2.63)
-As we already know, the quark contribution is always satisfied, i.e., $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=0$, which leads to $\Delta^{2}_{q}=0$. Thus, the relation (B.3) is reduced to
-$q_{\rho}q_{\sigma}\Pi^{YM}_{\rho\sigma}(q;D)={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (2.64)
-In complete analogy with the subtraction (4.19), let us now introduce the subtraction for the YM part as follows:
-$\Pi^{s(YM)}_{\rho\sigma}(q;D)$ $=$ $\Pi^{YM}_{\rho\sigma}(q;D)-\Pi^{YM}_{\rho\sigma}(0;D)$ (2.65)
-$=$ $\Pi^{YM}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{YM}(D)$
-where $\Pi^{s(YM)}_{\rho\sigma}(0;D)=0$, by definition, and
-$\Delta^{2}_{YM}(D)=\Delta^{2}_{g}(D)+\Delta^{2}_{t}(D),$ (2.66)
-as it follows from eq. (B.2).
-Instead of the tensor decompositions (4.22), for the YM parts one gets
-$\Pi^{s(YM)}_{\rho\sigma}(q)=T_{\rho\sigma}(q)q^{2}\Pi^{s(YM)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(YM)}_{l}(q^{2}),$ (2.67)
-and
-$\Pi^{YM}_{\rho\sigma}(q)=T_{\rho\sigma}(q)q^{2}\Pi^{YM}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{YM}_{l}(q^{2}).$ (2.68)
-In all the quantities above and below the dependence on $D$ is omitted, for simplicity, and will be restored when necessary.
-Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary.
-However, all the invariant functions $\Pi^{s(YM)}_{t}(q^{2})$, $\Pi^{s(YM)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the subtraction (B.5).
-Substituting the relations (B.7)-(B.8) into the relations (B.4)-(B.5), by omitting the tedious tensor algebra derivations, one finally obtains
-$\Pi^{s(YM)}_{t}(q^{2})$ $=$ $\Pi^{YM}_{t}(q^{2})-{\Delta^{2}_{YM}(D)\over q^{2}},$
-$\Pi^{s(YM)}_{l}(q^{2})$ $=$ $-{(\xi_{0}-\xi)\over\xi\xi_{0}}+{\Delta^{2}_{YM}(D)\over q^{2}}.$ (2.69)
-However, the invariant function $\Pi^{s(YM)}_{l}(q^{2})$ cannot have the pole-type singularities, by definition, as underlined just above.
-So that one has to put the QD but regularized constant $\Delta^{2}_{YM}(D)=0$ on this mathematical ground.
-Taking into account that the quark constant $\Delta^{2}_{q}=0$ as well, one concludes that all the QD but regularized constants should be removed from theory, i.e., to be put zero as follows:
-$\Delta^{2}_{q}=\Delta^{2}_{g}(D)=\Delta^{2}_{t}(D)=0,$ (2.70)
-because of the relation (B.6) and all of them are independent from each other.
-If $\Delta^{2}_{t}(D)=0$ then from the relation (B.2) it follows that $\Pi^{YM}_{\rho\sigma}(q;D)=\Pi^{g}_{\rho\sigma}(q;D)$.
-But from the satisfied transverse relation $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ and previous equality it follows that $q_{\rho}q_{\sigma}\Pi^{YM}_{\rho\sigma}(q;D)=0$ as well.
-Then from the relation (B.4) one gets $\xi_{0}=\xi$.
-The relations (B.10) and this equality determines the particular, or, equivalently, conventional QCD, investigated in section 6.
-Here it has been obtained by treating all the QD constants on the same footing, completely ignoring their principle difference from the physical point of view, which has been emphasized throughout this work.
-This happened because in the previous derivations we deliberately have ignored that the subtraction (B.5), on account of the relation (B.2), in fact, looks like
-$\Pi^{s(YM)}_{\rho\sigma}(q;D)=\Pi^{g}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{g}(D),$ (2.71)
-i.e., $\Pi^{s(YM)}_{\rho\sigma}(q;D)=\Pi^{s(g)}_{\rho\sigma}(q;D)$, as it follows from eq.
-(4.18).
-Let us underline that the disappearance of the tadpole term is not a specific feature of our subtraction scheme.
-This always will happen in any subtraction scheme, since the constant minus the same constant is always zero, as pointed out in section 4.
-That is why explicitly present in the initial expressions it will remain in the final ones, compare eq.
-(2.5) with eq.
-(5.27), and thus with eq.
-(5.34).
-Substituting further the tensor decomposition (B.7) and second of the decompositions (4.22) into the subtraction (B.11), and again omitting some tedious tensor algebra derivations, one obtains
-$\Pi^{s(YM)}_{t}(q^{2})$ $=$ $\Pi^{g}_{t}(q^{2})-{\Delta^{2}_{g}(D)\over q^{2}},$
-$\Pi^{s(YM)}_{l}(q^{2})$ $=$ $\Pi^{g}_{l}(q^{2})+{\Delta^{2}_{g}(D)\over q^{2}}.$ (2.72)
-The general relation (B.4), on account of the first line of the subtraction (B.5), becomes
-$q_{\rho}q_{\sigma}[\Pi^{s(YM)}_{\rho\sigma}(q;D)+\Pi^{YM}_{\rho\sigma}(0;D)]={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (2.73)
-Substituting now the decomposition (B.7) and the relation $\Pi^{YM}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{YM}(D)$, which follows from the relations (B.5), into the previous equation, and omitting again tensor algebra derivations, one obtains
-$\Pi^{s(YM)}_{l}(q^{2})=-{(\xi_{0}-\xi)\over\xi\xi_{0}}+{(\Delta^{2}_{g}(D)+\Delta^{2}_{t}(D))\over q^{2}}.$ (2.74)
-Equating now this relation to the second of the relations (B.12), and doing simple algebra, one finally arrives at
-${(\xi_{0}-\xi)\over\xi\xi_{0}}q^{2}=\Delta^{2}_{t}(D),$ (2.75)
-which is nothing else but the general constraint (3.15).
-Note that in deriving this expression, we again use that from the satisfied transverse relation for the gluon part $q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$ it follows that $\Pi^{g}_{l}(q^{2})=0$.
-As in the case of the conventional QCD, the properties of the gluon degrees of freedom have been used at the last stage.
-Even introducing the YM contribution to the full gluon self-energy in eq.
-(B.1), which hides the tadpole term, but correctly treated its subtraction (B.5) by the subtraction (B.11), one again arrives at the exact constraint (B.15), nevertheless.
-Treated it literally (i.e., by eq.
-(B.5) only) leads to the conventional QCD.
-In this case the significant piece of the information on the true dynamical and gauge structures of the QCD vacuum is losing.
-In other words, the conventional QCD hides the true role of the tadpole term.
-Contrary, in general QCD, based on the exact solution of eq.
-(B.15), this information is not losing and the true role of the tadpole has been revealed.
-Concluding, let us emphasize that the exact constraint (3.15) has been obtained from the very beginning before introducing any proper subtraction scheme.
-This underlines its general and unique character.
-At the same time, the clarifying algebraic derivations performed here, leading to the relations (B.10), heavily depend on the not accurate (i.e., literally) treatment of the subtraction (B.5), as explained just above.
-Appendix C Renormalized ST identity
-Substituting eq. (3.16) into the ST identity (3.7), one obtains
-$q_{\mu}q_{\nu}D_{\mu\nu}(q)=i\xi=i{\xi_{0}q^{2}\over q^{2}+\xi_{0}\Delta^{2}_{t}(D)}.$ (3.76)
-Let us introduce the non-perturbative (NP) renormalization constant for the tadpole term as follows: $\Delta^{2}_{t}(D)=Z_{\Delta}\Delta^{2}$, where $\Delta^{2}$ is finite and positive, by definition, while the NP renormalization constant $Z_{\Delta}$ depends on all the unphysical parameters, such as regulating parameters, mentioned above, etc.
-In principle, the string tension in the linear rising potential between heavy quarks has to be somehow related to the finite $\Delta^{2}$, since it is connected to the transverse projection operator structure, as underlined above.
-It is well-known that the ansatz $D_{\mu\nu}(q)\sim\delta_{\mu\nu}\Delta^{2}/(q^{2})^{2}$ just leads to a such potential.
-After doing some algebra, the ST identity for the renormalized full gluon propagator becomes
-$q_{\mu}q_{\nu}D^{R}_{\mu\nu}(q)=i\xi^{R}=i{\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}},$ (3.77)
-where the renormalized gauge-fixing parameters are defined as follows: $\tilde{\xi}_{0}=Z_{\Delta}\xi_{0},\ \xi^{R}=Z_{\Delta}\xi$ and $D^{R}_{\mu\nu}(q)=Z_{\Delta}D_{\mu\nu}(q)$, i.e., the ST identity (C.1) is expressed in terms of the finite quantities only.
-The correct PT limit is maintained, of course.
-Let us now demonstrate the interesting feature of the generalized gauge (C.2). It provides the self-consistency condition for the gauge choice in QCD as follows:
-${\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}}={aq^{2}\over q^{2}+a\Delta^{2}},$ (3.78)
-[…]
-${\tilde{\xi}_{0}q^{2}\over q^{2}+\tilde{\xi}_{0}\Delta^{2}}={q^{2}\over\Delta^{2}},$ (3.79)
-which is only satisfied at $q^{2}=0$, i.e., there is no any condition for the gauge-fixing parameter.
-The formulated self-consistency condition (C.3) points out on the inconsistency of the canonical gauge $\tilde{\xi}_{0}=\infty$ in QCD.
-Briefly, if everything is expressed in the terms of the finite quantities then the formal canonical gauge is forbidden to use in QCD, according to our approach to this theory.
-Let us now apply the above-developed formalism to the renormalized massive gluon fields.
-For this purpose, it is instructive to re-write eq.
-(7.46) for the free massive gluon propagator in the terms of the renormalized quantities.
-But before it is convenient to introduce the following notations: $\tilde{\xi}_{0}=\tilde{\lambda}^{-1}$ and $M^{2}=M^{2}_{g}$, and the corresponding renormalization constant $Z_{\Delta}=1$ in this case.
-Then eq.
-(7.46) will look like
-$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2}_{g})}\left[\delta_{\mu\nu}-{(1-\tilde{\lambda}^{-1})q_{\mu}q_{\nu}\over(q^{2}-\tilde{\lambda}^{-1}M^{2}_{g})}\right].$ (3.80)
-From this observation it follows that our approach makes it possible for gluon to acquire mass dynamically without introducing into the theory some extra degrees of freedom such as Higgs fields [41], mentioned above, and $M^{2}_{g}$ can be treated as exactly defined gluon pole mass.
-If one puts the expression (C.5) on the mass-shell $q^{2}=M^{2}_{g}$, then one arrives at
-$D^{0}_{\mu\nu}(q)={-i\over(q^{2}-M^{2}_{g})}\left[\delta_{\mu\nu}-{q_{\mu}q_{\nu}\over M^{2}_{g}}\right],$ (3.81)
-which leads to the non-renormalizable behaviour in the PT $q^{2}\rightarrow\infty$ limit (i.e., it is constant and not $\sim 1/q^{2}$ as it is required for the renormalizable behaviour).
-However, the same expression (C.6) will be obtained if one expresses (C.5) in the canonical gauge ${\tilde{\lambda}}^{-1}=\infty$.
-As explained above, this gauge is not to be used in QCD.
-This means that within the generalized gauge-fixing function approach the free massive gluons cannot be put on the mass-shell.
-The finite values for the gauge-fixing parameter ${\tilde{\lambda}}^{-1}$ are only permitted in order to guarantee the theory to be renormalizable.The expression (C.5) has the correct PT limit, being off-mass-shell object at any finite ${\tilde{\lambda}}^{-1}$.
-So that, the canonical gauge is rather formal than real within the generalized gauge-fixing function approach to QCD.
-The canonical gauge and its connection to the structure of the full gluon propagator at the mass-shell will be investigated in our future work, where the NP renormalization theory of the massive gluon fields will be present in detail.
-Appendix D Preservation of the gauge symmetry in the QED ground state
-The propagation of the photon in the QED vacuum is formally described by eq.
-(2.1).
-Contrary to QCD (compare with eq.
-(2.2), the difference is that there is only one contribution to the full photon self-energy comes from the vacuum polarization tensor.
-Its skeleton term is similar present in Fig.
-1 by the first diagram.
-Analytically it looks like
-$\Pi^{ph}_{\rho\sigma}(q)\sim\int d^{4}l\gamma_{\rho}S(l-q)\Gamma_{\sigma}(l,q)S(l),$ (4.82)
-where all the overall numerical factors are omitted, for simplicity.
-Here $S$ and $G$ represent the full electron propagator and the full electron-photon vertex, respectively.
-The full photon and its free counterpart are analytically present similar to eq.
-(3.8) and eq.
-(3.9), respectively, as well as their ST identities by relations (3.7) and (3.10).
-Similar to the relation (3.11) for the transverse relation in this case, one obtains
-$q_{\rho}q_{\sigma}\Pi^{ph}_{\rho\sigma}(q;D)={(\xi_{0}-\xi)\over\xi\xi_{0}}(q^{2})^{2}.$ (4.83)
-It is well-known that because of the electric charge conversation, the transverse relation for the vacuum polarization tensor in QED is always satisfied, i.e., it is zero.
-Then from the previous equation immediately follows
-$q_{\rho}q_{\sigma}\Pi^{ph}_{\rho\sigma}(q;D)=0,\quad\rightarrow\quad\xi_{0}=\xi.$ (4.84)
-Let us now introduce the corresponding subtraction scheme similar to the relation (4.19) and their independent tensor decompositions (4.22), namely
-$\Pi^{s(ph)}_{\rho\sigma}(q)=\Pi^{ph}_{\rho\sigma}(q)-\Pi^{ph}_{\rho\sigma}(0)=\Pi^{ph}_{\rho\sigma}(q)-\delta_{\rho\sigma}\Delta^{2}_{ph},$ (4.85)
-where $\Pi^{s(ph)}_{\rho\sigma}(0)=0$, by definition, and
-$\Pi^{s(ph)}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(ph)}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{s(ph)}_{l}(q^{2}),$
-$\Pi^{ph}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{ph}_{t}(q^{2})-q_{\rho}q_{\sigma}\Pi^{ph}_{l}(q^{2}),$ (4.86)
-respectively.
-Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary.
-However, the both invariant functions $\Pi^{s(ph)}_{t}(q^{2})$, $\Pi^{s(ph)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the initial subtraction (D.4).
-Substituting all these decompositions (D.5) into the subtraction (D.4), and doing some tensor algebra derivations, one finally obtains
-$\Pi^{s(ph)}_{t}(q^{2})$ $=$ $\Pi^{ph}_{t}(q^{2})-{\Delta^{2}_{ph}\over q^{2}},$
-$\Pi^{s(ph)}_{l}(q^{2})$ $=$ $\Pi^{ph}_{l}(q^{2})+{\Delta^{2}_{ph}\over q^{2}}.$ (4.87)
-From the first of the relations (D.3) and the second of the relations (D.5) it follows that $\Pi^{ph}_{l}(q^{2})=0$. Then the second of the relations (D.6) will be reduced to
-$\Pi^{s(ph)}_{l}(q^{2})={\Delta^{2}_{ph}\over q^{2}},$ (4.88)
-which is impossible since the corresponding invariant function can not have the pole-type singularities, by definition, as explained above.
-One has to put the QD but regularized constant to zero on this mathematical basis, i.e., $\Delta^{2}_{ph}=0$, and thus $\Pi^{s(ph)}_{l}(q^{2})=0$ as well.
-As we already know, the both relations
-$\xi=\xi_{0},\quad\Delta^{2}_{ph}=0$ (4.89)
-are necessary conditions for the coincidence of the gauge symmetry of the QED Lagrangian with the gauge symmetry of its ground state (like it occurs for the conventional QCD).
-This is our proof of the Elitzur theorem [42] for the continuos abelian fields within our approach, based on the formalism of the dynamical equations of motion for the gauge particle Euclidean Green’s functions.